### v1

# Archilles’ Heel in Semi-open LLMs: Hiding Bottom against Recovery Attacks

Abstract
Closed-source large language models deliver strong performance but have limited downstream customizability. Semi-open models, combining both closed-source and public layers, were introduced to improve customizability. However, parameters in the closed-source layers are found vulnerable to recovery attacks. In this paper, we explore the design of semi-open models with fewer closed-source layers, aiming to increase customizability while ensuring resilience to recovery attacks. We analyze the contribution of closed-source layer to the overall resilience and theoretically prove that in a deep transformer-based model, there exists a transition layer such that even small recovery errors in layers before this layer can lead to recovery failure. Building on this, we propose SCARA [^1] Code is available at: https://github.com/OTTO-OTO/SCARA-Semi-Open, a novel approach that keeps only a few bottom layers as closed-source. SCARA employs a fine-tuning-free metric to estimate the maximum number of layers that can be publicly accessible for customization. We apply it to five models (1.3B to 70B parameters) to construct semi-open models, validating their customizability on six downstream tasks and assessing their resilience against various recovery attacks on sixteen benchmarks. We compare SCARA to baselines and observe that it generally improves downstream customization performance and offers similar resilience with over 10 times fewer closed-source parameters. We empirically investigate the existence of transition layers, analyze the effectiveness of our scheme and finally discuss its limitations.
1 Introduction
Figure 1: Semi-open vs. fully-closed model.
Vendors of Large Language Models (LLMs) have recently launched several closed-source models with impressive capabilities, serving diverse user needs across different scenarios (Minaee et al., 2024; Zhao et al., 2023). Access to closed-source models is typically through black-box APIs, such as GPT-4o and Claude 3.5. These APIs hide model weights and internal structures, restricting downstream customizability. To address this, semi-open LLMs have emerged, offering more flexibility while still hiding certain infrastructure details (Eiras et al., 2024; White et al., 2024). As shown in Figure 1, these semi-open models keep some modules closed-source but allow access to others. This enables users to interact with closed-source modules via embedding APIs and fine-tune the open modules for downstream tasks.
Open-sourcing more parameters and structure details apparently enhances downstream customizability. However, Zanella-Beguelin et al. (2021) showed that semi-open LLMs with only a few closed-source parameters are vulnerable to model recovery attacks. Recovery attackers query the closed-source module and then train a new module that imitates its functionality. This can lead to the full replication and theft of closed-source modules (Solaiman, 2023). Although the community has extensively studied recovery attacks against closed-source models (Chen et al., 2023; Jiang et al., 2023b), defending against recovery attacks in the semi-open setting is an uncharted area. Recovery attackers targeting fully closed-source models seek to fine-tune a new model that precisely replicates the closed-source model (Tamber et al., 2024; Dubiński et al., 2024). In contrast, attackers in semi-open settings are not required to exactly replicate the closed-source module. Instead, they can fine-tune the closed-source module alongside the public module to reconstruct the overall functionality. While open-sourcing more layers enhances downstream flexibility, it also facilitates easier replication.
Beyond the closed-source amount, the specific sections concealed are vital for defending against recovery attacks. Shen et al. (2023) suggested concealing several top layers and keeping the bottom public. However, the benefits of hiding bottom layers (near the input) versus top layers (near the output) are still unclear. Therefore, we examine the impact of each layer on resilience and theoretically identify a transition layer. Any recovery error in bottom layers before this transition layer leads to a high probability of recovery failure. In contrast, errors in later layers have limited impact. This finding suggests that keeping the bottom layers closed-sourced offers better protection against recovery attacks than the top layers, even when the same number of layers is hidden.
In this paper, we introduce SCARA, a selective closed-sourcing method for designing semi-open models that balances customizability with resilience against recovery attacks. Building on our theoretical findings, SCARA selectively hides a few bottom layers. It determines the minimal number of closed-source layers using a recovery difficulty score, a metric that estimates recovery performance without requiring fine-tuning. This score is based on the initial average recovery loss during the attack. SCARA identifies the optimal closed-source strategy by selecting the layers with metric corresponding to the worst recovery performance. Consequently, models designed by SCARA retain most layers as publicly accessible, achieving customizability comparable to fully open-source models while remaining resilient to recovery attacks. Our main contributions are as follows:
  • We theoretically demonstrate the existence of a transition layer in LLMs. We prove that small recovery errors in bottom decoder layers before this layer can lead to recovery failure with high probability, whereas errors in later layers have a limited impact. (see Section 4.1)
  • We introduce SCARA, a selective closed-source approach that conceals a few bottom decoder layers to enhance customizability while maintaining resilience. Specifically, we propose a metric that does not require fine-tuning, but correlates with the recovered performance under attacks, enabling us to approximately find the minimal number of hidden layers. (see Section 4.2)
  • We compare our approach with two baselines across five models (1.3B to 70B parameters), assessing customizability on six tasks and resilience against three recovery strategies across sixteen benchmarks in six domains. Experiments show that our method significantly improves downstream performance while maintaining comparable resilience against recovery attacks, with over 10 times fewer closed-source parameters than the baselines. For example, the semi-open Llama2-70B produced by our method hides only 2.5% of the parameters but achieves a 30% higher downstream performance score than the baselines in the Financial domain. We also observe a performance improvement of over 40% on Mistral-7B. Additionally, our method maintains similar resilience against recovery attacks compared to both baselines. (see Section 5.2)
  • We empirically investigate the presence of transition layers and the correlation between our metric and the recovered performance of each closesd-source combination. We conclude by analyzing the hyper-parameter sensitivity and discuss the limitations of our approach. (see Section 5.4)
2 Related Works
Model Customization. Vendors have introduced three main strategies for model customization, each with distinct trade-offs. First, fine-tuning APIs allow users to customize fully closed-source models (e.g., La Plateforme, Azure AI Services), while restricting access (Finlayson et al., 2024). Second, embedding models provide richer customization by allowing users to select and modify subsequent structures (Sarıtaş et al., 2024; Lee et al., 2024). However, the absence of joint pre-training can degrade performance (Nussbaum et al., 2024) and expose the closed-sourced models to recovery attacks (Caspari et al., 2024; Tamber et al., 2024). Third, open-source models enable full customization flexibility through free fine-tuning, yet pose challenges to model control and usage supervision (Bommasani et al., 2022; Roumeliotis et al., 2023).
Model Recovery Attacks. Prior attacks (Tramèr et al., 2016; Krishna et al., 2020; Dziedzic et al., 2023a) attempt to recover the functionality of fully closed-source models through API queries.  Carlini et al. (2024) advanced these by entirely extracting the embedding projection layer and hidden dimension size. Recently, various defenses against fully closed-source model recovery attacks have been proposed (Jiang et al., 2024), including malicious queries detection (Shang et al., 2024), watermarking (Zhang et al., 2021), fingerprinting (Guan et al., 2022), etc. These methods do not directly apply to the semi-open settings with only partial model information.
Semi-Open Model. Previous studies (Lin et al., 2024; Chen et al., 2024; Qiao & Zhou, 2023) explore opening bottom layers of models for user customization, while keeping later layers closed-sourced to maintain vendor control. For example, Shen et al. (2023) introduced SAP, which open-sources the first six transformer layers but limits customization options. Meanwhile, Dubiński et al. (2024); Dziedzic et al. (2023b) proposed a semi-open approach where encoder models are offered as APIs, allowing users to customize task-specific subsequent modules. However, (Liu et al., 2022) and Sha et al. (2023) showed that these encoder models are still vulnerable to recovery attacks.
3 Preliminaries
3.1 Security Threat: Semi-open Model Recovery
Semi-open LLMs. Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ denote the input data matrix, where each row corresponds to a $d$-dimensional feature vector representing a single token. Let $f:\mathbb{R}^{n\times d}\rightarrow\mathcal{Y}$ denote a victim model, capable of processing the feature matrix $\mathbf{X}$ and producing an element in the set $\mathcal{Y}$ as output. Modern LLMs typically adopt a multi-layer architecture to capture complex patterns in the input data. Specifically, $f$ is a composition of multiple decoder layers, i.e., $f(\mathbf{X};\bm{\theta})=\varphi_{L}\circ...\circ\varphi_{1}(\mathbf{X})$. All decoder layers $\varphi_{1},...,\varphi_{L}$ share the same architecture but each layer is equipped with distinct parameters. The parameters of all layers are denoted by the vector $\bm{\theta}$. We consider a semi-open setting, in line with  Zanella-Beguelin et al. (2021) and Xu et al. (2021), where certain layers of the LLM are closed-sourced while others remain public. Let the closed-sourced set $I\subseteq\{1,\ldots,L\}\triangleq[L]$ denote the index set of hidden layers, while its complement $I^{c}$ contains the public layer indices.
Figure 2: Workflow of semi-open model recovery attack
Semi-open Model Recovery Attack. The semi-open model recovery attack is one of the high-fidelity attacks (Carlini et al., 2024) that aims to recover the general functionality of the victim model. As illustrated in Figure 2, the adversary is able to interact with the victim LLM and possesses access to the architecture and parameters of its public layers. The adversary aims to emulate the behavior of the victim LLM by learning the architecture and parameters associated with the closed-source layers. This task is accomplished through a structured procedure comprising three main steps: (1) Attack Dataset Construction: The adversary begins by querying the victim model, thereby gathering an attack dataset $\mathcal{D}$ containing samples representing the capabilities of the victim LLM. (2) Parameter Initialization: Next, the adversary randomly initializes the parameters associated with the closed-sourced layers, setting the stage for subsequent model fine-tuning. (3) Model Fine-Tuning: Leveraging the dataset $\mathcal{D}$ obtained in the first step, the adversary fine-tunes the parameters of the entire model, iteratively adjusting the parameters to better align with the observed behavior of the victim LLM. Let $\bm{\theta}_{\text{FT}}(I,\mathcal{D})$ denote the recovered parameters under the attack dataset $\mathcal{D}$ and closed-sourced set $I$. Prior work has shown that this approach enables the adversary to replicate the general functionality of the semi-open LLMs, despite limited access.
3.2 Problem Formulation
[…]
$R(I)=\frac{\mathbb{E}[s(f(\mathbf{X};\bm{\theta}_{\text{FT}}(I,\mathcal{D})),Y)]}{\mathbb{E}[s(f(\mathbf{X};\bm{\theta}),Y)]}.$ (1)
Here, $\mathbb{E}$ in the numerator reflects the expectation computed over random samples $(\mathbf{X},Y)$ drawn from $\mathbb{P}_{\mathbf{X}\times Y}$, the random attack set $\mathcal{D}$, and the random initialization of parameters within the closed-source layers during fine-tuning. Conversely, the term $\mathbb{E}$ in the denominator solely considers the expectation over random samples. With this definition, the term $R([L])$ denotes the recovery ratio of the recovered model under a fully-closed approach, where $[L]=\{1,...,L\}$. Hence, the following question arises.
Given $\varepsilon>0$, what is the smallest closed-source set $I$ for which $R(I)\leq(1+\varepsilon)R([L])$?
This question essentially asks whether it is feasible to identify a minimal closed-source index set $I$, such that, under this closed-source strategy, the resulting recovered model exhibits similarity to the model recovered under fully-closed approach. In other words, the recovery score does not surpass that of fully-closed approach by more than a factor of $(1+\varepsilon)$.
4 Methodology
Figure 3: Customizability and resilience comparison in Llama2-70B. Higher scores indicate better customizability in Fig. (a) and weaker resilience in Fig. (b). Details can be found in Appendix C.1
In this section, we investigate how each layer affects customizability and resilience against recovery attacks. We begin with an experiment involving two semi-open Llama2-70B models, each with either the first two (Semi-Open-1) or the last two (Semi-Open-2) decoder layers closed-sourced. We compare their customization performance and recovered performance under the recovery attack. Figure 3 (a) and (b) show that although two semi-open models perform similarly on six downstream tasks, closed-sourcing the first two layers offers significantly greater resilience than the last two. Moreover, we compare the Semi-Open-1 model to the fully-closed model and observe that this semi-open model can achieve better customizability and comparable resilience at the same time. Therefore, we conjecture that, with a sufficient number of closed-sourced layers before a certain transition layer, a semi-open model can simultaneously achieve great customizability on downstream tasks and strong resilience against recovery attacks. In this section, we first present a theoretical result showing the existence of transition layers and then introduce our selective closed-sourcing approach.
4.1 Resilience Transition Layer in Infinitely Deep Transformers
Model Overview. Let us revisit our large language model composed of $L$ layers, denoted as $f(\mathbf{X};\bm{\theta})=\varphi_{L}\circ...\circ\varphi_{1}(\mathbf{X})$. Recall that each row of the feature matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ represents a $d$-dimensional vector for an input token. We treat each layer $\varphi_{i}$ as a transformer layer, where each layer processes an $n\times d$ dimensional matrix as input and outputs another $n\times d$ matrix. Thus, the model $f$ outputs a matrix of $n$ rows and $d$ columns, indicating that the large language model outputs a feature vector for each token. Moreover, we assume that each layer contains a normalized residual self-attention function, defined as
$\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)=\mathbf{X}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\mathbf{X},$ (2)
where $Q_{i}\in\mathbb{R}^{d\times d_{Q}}$ and $K_{i}\in\mathbb{R}^{d\times d_{Q}}$ are projection parameter matrices for the $Q$ and $K$ matrices in the transformer, respectively. Additionally, $\sqrt{d_{Q}}$ and the matrix norm $\|\mathbf{X}\|$ denote normalization factors provided by the normalization layer. We consider the strategy of concealing the $\alpha L$-th layer with $\alpha\in[0,1]$ and $\alpha L\in\mathbb{N}$ while keeping other layers public. After the semi-open model recovery, we assume the parameters of the recovered model in the public layers are identical to the victim model, while those in the proprietary layer deviate. Let $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ denote the recovered weight matrix of the proprietary layer, i.e., ${\bm{\theta}}_{\text{FT}}(\{\alpha L\})=\{(K_{1},Q_{1}),...,(\hat{K}_{\alpha L},\hat{Q}_{\alpha L}),...,({K}_{L},{Q}_{L})\}$. Let $\hat{\varphi}_{\alpha L}$ denote the function of the recovered proprietary layer, i.e., the $\alpha L-$th layer, in the recovered model. In this subsection, we consider the normalized output of an infinitely deep model whose $\alpha L$-th layer is closed-sourced and subjected to the attack. The output of the recovered model is
$\hat{f}_{\infty}(\mathbf{X})=\lim_{L\rightarrow\infty}\frac{\varphi_{L}\circ...\varphi_{\alpha L+1}\circ\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ...\circ\varphi_{1}(\mathbf{X})}{\|\varphi_{L}\circ...\varphi_{\alpha L+1}\circ\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ...\circ\varphi_{1}(\mathbf{X})\|_{F}},$
where $\|\cdot\|_{F}$ denotes the Frobenius norm of a given matrix. We consider this infinitely deep network as our ideal model because, in real-world settings, most large-scale models are sufficiently deep. Next, we present the following theorem to illustrate the existence of a critical value $\alpha^{*}$ such that if $\alpha<\alpha^{*}$, the recovered LLM outputs identical feature vectors for all tokens. Conversely, if $\alpha>\alpha^{*}$, the output feature vectors may vary across tokens.
Theorem 1.
Assume that $\mathbb{P}_{\mathbf{X}\times Y}$ is defined on a countable domain $\mathcal{X}\times\mathcal{Y}$ with $\mathbf{0}_{n\times d}\notin\mathcal{X}$. Assume that parameter matrices $\{K_{i},Q_{i}\}_{i\geq 1}$ in the victim model $f$ have uniform bounded norms, i.e., $\|K_{i}\|\leq D$ and $\|Q_{i}\|\leq D$ for some $D>0$. There exists an $\alpha^{*}\in(0,1)$ depending on $D$ such that the following two statements are true.
(1) Let $\alpha<\alpha^{\star}$ and $\{K_{i},Q_{i}\}_{i\geq 1}$ be any parameter matrix sequence in the victim model. Let $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ be the recovered parameter matrices drawn from a continuous distribution supported on $\mathbb{R}^{n\times d}$. With probability one, for any input $\mathbf{X}\in\mathcal{X}$, the row vectors in the matrix $\hat{f}_{\infty}(\mathbf{X})$ are identical.
(2) Let $\alpha>\alpha^{\star}$. There exists a victim model with parameter matrix sequence $\{K_{i},Q_{i}\}_{i\geq 1}$ such that for any recovered parameter matrices $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$, the row vectors in the matrix $\hat{f}_{\infty}(\mathbf{X})$ are not entirely the same for some input feature matrix $\mathbf{X}\in\mathcal{X}$.
Remark 1: The proof is provided in Appendix A. This theorem demonstrates that if the recovered parameters of the bottom layers (i.e., $\alpha<\alpha^{*}$) are obtained through a randomized algorithm, such as stochastic gradient descent, with a continuous distribution supported on $\mathbb{R}^{n\times d}$, the recovery will certainly fail, as it will produce the same feature vector for every token. In contrast, keeping the later layers closed-sourced (i.e., $\alpha>\alpha^{*}$) does not maintain this property, indicating that it is more effective to closed-source the bottom layers before the transition layer, rather than the later ones.
Remark 2: The theorem relies on the assumption that the distribution is defined over a countable domain, $\mathcal{X}\times\mathcal{Y}$, typically satisfied by inputs such as sentences or images. We show in the proof that for each input matrix $\mathbf{X}\in\mathcal{X}$, there are two zero-measure sets $\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}(\mathbf{X})$ such that the recovered matrices must avoid to satisfy the theorem. Hence, the countable unions $\mathcal{K}=\bigcup_{\mathbf{X}\in\mathcal{X}}\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}=\bigcup_{\mathbf{X}\in\mathcal{X}}\mathcal{Q}(\mathbf{X})$ are also zero-measure sets, ensuring that when recovered matrices do not belong to these sets, the conditions in the theorem are met for any input matrix $\mathbf{X}$ in the input space.
Theorem 1 shows that hiding bottom layers offers greater resilience against model recovery attacks, indicating that closed-sourcing layers starting from the first layer may be an effective approach. In the next subsection, we present a closed-sourcing approach to identify the minimal set of hidden layers.
4.2 SCARA: Selective Closed-sourcing Approach Against Recovery Attack
We propose a method to approximately find the smallest bottom layer index set $I$ that satisfies $R(I)\leq(1+\varepsilon)R([L])$. A simple approach is to start with $I_{l}=\{1,\ldots,l\}$ for each $l$ beginning from 1, then evaluate the recovery ratio $R(I_{l})$ after the attack, and identify the smallest $l$ that meets the inequality. This extensive fine-tuning process is time-consuming, prompting the critical question: Can we create a fine-tuning-free metric that predicts LLM performance under semi-open model recovery attacks? Hence, our goal is to establish a metric directly correlated with the recovery ratio.
In the recovery ratio $R(I)$, each $I$ has the same denominator, so our focus is on a metric related to the numerator, specifically $\mathbb{E}[s(f(\mathbf{X};\bm{\theta}_{\text{FT}}(I,\mathcal{D})),Y)]$, which measures the average performance score of the recovered model. This average performance score generally inversely correlates with the average testing loss $\mathbb{E}[\ell(f(\mathbf{X};\bm{\theta}_{\text{FT}}(I,\mathcal{D})),Y)]$. Therefore, our goal becomes finding the $l$ such that
$\mathbb{E}[\ell(f(\mathbf{X};\bm{\theta}_{\text{FT}}(\{1,...,l\},\mathcal{D})),Y)]\geq(1-\varepsilon)\mathbb{E}[\ell(f(\mathbf{X};\bm{\theta}_{\text{FT}}([L],\mathcal{D})),Y)].$
However, calculating both sides of this inequality requires knowing the recovered parameters from the fine-tuning process. To bypass this, we aim for an approximate solution. The recovered parameters are generated through gradient descent, starting from the initial parameters $\bm{\theta}_{0}(I)$, with the hidden layers being randomly initialized. Using the Taylor Expansion, we find
$\mathbb{E}\left[\ell\left(f(\mathbf{X};\bm{\theta}_{\text{FT}}(I,\mathcal{D})),Y\right)\right]=\mathbb{E}\left[\ell\left(f(\mathbf{X};\bm{\theta}_{0}(I)),Y\right)\right]+\mathcal{O}(\|\bm{\theta}_{\text{FT}}(I,\mathcal{D})-\bm{\theta}_{0}(I)\|_{2}).$
Previous research (Choi et al., 2024; Bailly et al., 2022) suggests the difference $\|\bm{\theta}_{\text{FT}}(I,\mathcal{D})-\bm{\theta}_{0}(I)\|_{2}$ is minor for large networks compared to the dataset size $|\mathcal{D}|$. In models like a single-layer ReLU network (Anthony et al., 1999; Zou et al., 2020), the difference $\|\bm{\theta}_{\text{FT}}(I,\mathcal{D})-\bm{\theta}_{0}(I)\|_{2}$ is of order $\mathcal{O}\left(\frac{|\mathcal{D}|}{\sqrt{N}}\right)$ (Jacot et al., 2018; Wei et al., 2019), where $N$, the number of model parameters, which is much larger than the dataset size in LLMs (Dubey et al., 2024; Liu et al., 2024). Hence, the first term that does not require fine-tuning dominates, suggesting it as a viable metric for predicting the recovery ratio. Thus, we define the first term as “Recovery Difficulty” (RD$(I)$) with the expression:
$\text{RD}(I)=\mathbb{E}_{\mathbf{X},Y,\bm{\theta_{0}}(I)}\left[\ell\left(f(\mathbf{X};\bm{\theta}_{0}(I),Y\right)\right].$
This score, which can be estimated using a sample average, represents the recovered performance of the model when specific layers $I$ are closed-sourced. A higher RD$(I)$ suggests worse recovery performance, indicating a lower recovery ratio $R(I)$. Therefore, our SCARA operates in the following way. SCARA begins by sampling evaluation data from the underlying distribution and then computes RD$(I_{l})$ for each set of closed-sourced layers $I_{l}=\{1,...,l\}$ for $l=1,...,L$. SCARA stops at the smallest $l^{*}$ that satisfies RD$(I_{l^{*}})\geq(1-\varepsilon)\text{RD}([L])$.
5 Experiments
5.1 Experimental Settings
Models. We consider five open-source, decoder-only structured LLMs with various architectures. Specifically, we select Llama2-70B-chat, Llama2-7B-chat (Touvron et al., 2023), Mistral-7B-v0.1 (Jiang et al., 2023a), Phi-2 (Abdin et al., 2024), and Phi-1.5 (Li et al., 2023). We designate these pre-trained models as the base models for customization and victims in semi-open model recovery attacks. Details of each model is provided in Appendix B.1.
Attack Methods. We recover models produced by different closed-source approaches using three attack methods: FT-all, FT-closed, and SEM (Stealing Embedding Model) (Tamber et al., 2024). FT-all finetunes the entire model, whereas FT-closed fine-tunes the closed-sourced set only. We also consider SEM, a typical attack that trains a substitute embedding module with data from the closed-sourced set. Following (He et al., 2021), a diverse attack set is required for full recovery. Therefore, we merge data evenly for two general datasets, MMLU benchmark (Hendrycks et al., 2021) and Alpaca 52k (Wang et al., 2022), resulting in a 51k combined set. Details are in Appendix B.2
Baselines. We compare SCARA with the other two baselines: SAP-DP and the fully-closed (Eiras et al., 2024) approach. The SAP (Shen et al., 2023) framework keeps the first six decoder layers open and the rest closed-source. SAP-DP extends SAP by adding Laplace noise to the model outputs, a common strategy for model protection (Lee et al., 2018). The fully-closed approach represents the extreme, where all layers are closed-sourced. Further details can be found in Appendix B.3
Implementation Details of SCARA. We apply the SCARA algorithm to identify the smallest closed-sourced set $I$ such that $R(I)\leq(1+\varepsilon)R([L])$. To calculate the recovery difficulty (RD), we use a 1.5k evaluation set sourced from the same datasets as the attack set. We explore $\varepsilon$ values from 0.05 to 1 and discover that setting $\varepsilon$ to 0.05 yields optimal performance on the evaluation set. We randomly initialize the hidden layers and average the RD across three seeds to select the smallest layer set. For more details on the dataset and the SCARA sensitivity on $\varepsilon$, see Appendix B.4 and Section 5.3.
Evaluation Benchmarks We assess customizability on six downstream tasks: Code (Zheng et al., 2024b), Math (Yue et al., 2023), Medical (Zhang et al., 2023), Finance (Wang et al., 2023b), Law (Guha et al., 2024), and Alignment (Meng et al., 2024). To fully evaluate recovered functionalities, we focus on six capabilities domains following Llama2 report (Touvron et al., 2023). Specifically, we assess the recovered model across sixteen benchmarks grouped into (1) Commonsense Reasoning (Rsn.); (2) Reading Comprehension (Read.); (3) World Knowledge (Knl.); (4) Code; (5) Math; and (6) General Ability (Gen.). More details can be found in Appendix B.5.
Metrics. We measure model customizability through its improvements on benchmarks. For resilience, we calculate the “Average Recovery Ratio” (ARR) by averaging the recovery ratios across benchmarks. A lower ARR indicates higher resilience offered by the closed-sourced set. Additionally, we define $\Delta\textbf{ARR}(I)=\text{ARR}(I)-\text{ARR}([L])$ to compare the resilience between closed-sourcing set $I$ and the fully-closed approach. A smaller $\Delta\textbf{ARR}$ suggests similar resilience to the fully-closed model.
Figure 4: Customization performance of models closed-sourced by SCARA on six downstream tasks.
Table 1: Recovery ratios on 6 functionalities under FT-all (SCARA ||||SAP-DP|||| Fully-closed). “H.E.” in Code domain presents the benchmark “HumanEval”. More details are available in Appendix C.2.
Benchmark Llama2-70B Llama2-7B Mistral-7B Phi-2 Phi-1.5
Rsn. PIQA 62.6||||59.8||||63.0 64.7||||64.7||||64.6 63.0||||61.2||||60.2 68.3||||65.6||||65.7 70.6||||69.5||||66.7
Winogrande 68.5||||67.7||||68.3 76.8||||74.8||||76.6 67.2||||69.0||||68.3 68.3||||64.9||||64.8 70.3||||67.8||||67.6
ARC-easy 31.9||||32.8||||31.3 36.3||||35.5||||34.9 32.3||||34.7||||32.0 43.2||||35.3||||33.9 40.5||||37.8||||36.1
ARC-challenge 38.5||||38.1||||44.2 47.8||||46.6||||50.9 39.7||||42.6||||44.5 36.8||||36.6||||35.3 46.1||||44.4||||47.5
Hellaswag 31.4||||31.4||||32.4 33.9||||34.0||||35.0 32.2||||32.0||||31.3 37.4||||37.3||||34.3 42.0||||41.0||||40.0
Read. LAMBADA 0.01||||0.00||||0.00 0.02||||0.00||||0.01 0.16||||0.00||||0.01 1.34||||0.04||||0.00 1.37||||0.00||||0.00
BoolQ 47.2||||47.1||||53.9 59.5||||56.0||||65.0 48.3||||46.8||||56.7 56.7||||50.3||||55.8 61.7||||54.9||||60.8
SQuADv2-EM 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00
SQuADv2-F1 1.50||||1.68||||0.34 0.68||||0.88||||0.82 1.69||||0.36||||0.93 3.65||||0.39||||0.90 1.28||||1.07||||2.64
OBQA 54.5||||54.5||||57.1 57.4||||52.5||||59.2 57.7||||56.8||||56.3 0.00||||0.00||||0.02 0.04||||0.00||||0.00
Knl. NaturalQuestions 0.00||||0.02||||0.00 0.01||||0.01||||0.08 0.00||||0.00||||0.02 0.01||||0.00||||0.06 0.21||||0.00||||0.00
TriviaQA 0.00||||0.02||||0.00 0.00||||0.00||||0.03 0.00||||0.00||||0.01 0.01||||0.00||||0.01 0.01||||0.00||||0.00
Code MBPP&H.E. 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00
Math GSM8K 0.02||||0.00||||0.06 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00
Gen. MMLU 36.8||||38.3||||36.5 52.9||||50.0||||53.3 40.4||||36.9||||37.2 42.6||||40.3||||40.5 56.7||||54.1||||54.1
BBH 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.01||||0.00||||0.00 0.00||||0.00||||0.00
Average Recovery Ratio($\downarrow$) 21.9||||21.8||||22.8 25.3||||24.4||||25.9 22.5||||22.4||||22.8 23.9||||22.3||||22.4 26.2||||25.3||||25.4
Closed-source Ratio($\downarrow$) 2.50||||92.5||||100. 3.16||||81.3||||100. 3.16||||81.6||||100. 6.25||||81.3||||100. 8.33||||75.0||||100.
Table 2: Recovery ratios on Llama2-70B.
Strat. Method Rsn. Read. Knl. C.&M. Gen. ARR
FT-c. SCARA 47.1 21.6 0.00 0.03 18.7 22.6
SAP-DP 46.2 19.5 0.00 0.00 19.0 21.8
F-Closed 47.8 21.2 0.00 0.08 18.5 22.8
SEM SCARA 48.2 21.9 0.00 0.00 18.5 22.4
SAP-DP 47.1 21.1 0.00 0.00 18.3 22.3
F-Closed 47.8 21.2 0.00 0.08 18.5 22.8
Table 3: ARR of SCARA vs. attacks.
Model FT-all FT-closed SEM
Llama2-70B 21.9 22.6 22.4
Llama2-7B 25.3 24.8 25.0
Mistral-7B 22.5 22.5 22.0
Phi-2 23.9 23.7 22.1
Phi-1.5 26.2 26.9 24.7
5.2 Main Results
In this subsection, we compare SCARA with three baselines, demonstrating its superior customizability on downstream domains while preserving similar resilience against model recovery attacks.
Customizability: SCARA vs. Baselines. We compare the customization performance of SCARA with closed-source baselines. Results are shown in Figure 4 and detailed in Appendix B.6.
On 70B and 7B models, SCARA consistently surpasses SAP-DP and fully-closed approaches across six domains and aligns closely with the performance of the fully-open approach, where all parameters are accessible. For instance, in the Law domain, SCARA improves scores by 10% over SAP-DP and fully-closed approaches on Llama2-70B, with this improvement rising to 35% on 7B models. Similar patterns of enhanced customizability are also evident in Phi-2 model, though the improvement on the Law domain narrows to only 1%. Furthermore, SCARA maintains performance comparable to the fully-open approach, demonstrating that hiding a small portion of parameters has a small impact on downstream customizability. More results are reported in Appendix C.3.
Resilience: SCARA vs. Baselines. We compare the resilience of SCARA with other closed-source baselines under three recovery strategies: FT-all, FT-closed, and SEM attack. As shown in Table 1, under FT-all attack, SCARA, SAP-DP, and fully-closed approach show similar resilience across various architectures and domains, with ARR differences within 1.4%. For instance, on Llama2-70B, SCARA keeps only 1.25% of parameters hidden, yet achieves an ARR of 21.9%, comparable to SAP-DP (21.8%) and the fully-closed approach (22.8%), which keep 92.5% and 100% of parameters closed-sourced, respectively. This pattern also extends across five architectures, indicating that our approach effectively preserves resilience with limited parameter closed-sourced. As shown in Table 3, SCARA achieves comparable resilience under FT-closed and SEM attacks. Under FT-closed attack, the recovery ratio differences among the three approaches remain under 2.1% across six domains. Similarly, under SEM attack, recovery ratios of SCARA align closely with the other approaches. These results show that SCARA offers resilience against model recovery on par with fully-closed approach, despite requiring fewer private parameters. More details are in Appendix C.4.
Figure 5: Amplification of error in Llama2-7B.
Table 4: SCARA vs. dataset scales.
Scale Rsn. Read. Knl. C.&M. Gen. ARR
51k 51.7 21.6 0.01 0.00 28.3 25.3
100k 51.3 21.5 0.13 0.00 29.6 25.3
200k 51.4 21.7 0.11 0.00 29.7 25.2
300k 51.6 21.7 0.11 0.00 30.5 25.5
500k 51.8 22.0 0.09 0.00 30.8 25.8
Resilience: SCARA vs. Recovery Strategies. Table 3 shows that SCARA effectively defends three recovery attack strategies on all models. We observe that SEM, a typical and effective attack for recovering embedding models, does not show a significant boost in recovery performance. This can be because SEM attackers focus on recovering only the proprietary embedding module, while the semi-open model recovery attackers aim at recovering the full functionality of the entire model, including both proprietary and public modules. The targets of these two attackers are different since even small errors in bottom layers can lead to significant output deviations. To see this, we add small perturbations on parameters in the first layer of Llama2-7B model and evaluate the embedding deviation at the output of each decoder layer. Figure 5 shows that the norm of deviation increases as the layer index increases, indicating that small errors are amplified by subsequent layers, leading to large deviations in the final output. Therefore, SCARA closed-sources the first several layers, effectively leveraging this amplification, making the functionality recovery more difficult and ensuring strong resilience against recovery attack. We report details in Appendix B.7 and C.5.
Resilience: SCARA vs. Recovery Dataset Scales. We further evaluate the resilience of SCARA against FT-all by increasing the recovery dataset scale on Llama2-7B to determine if larger datasets would compromise its effectiveness. More details on the attack dataset are in Appendix 7. Table 4 shows the recovery ratio achieved by SCARA under each attack dataset. We observe that increasing the scale of the attack dataset leads to only a mild increase in recovery ratios, indicating a limited impact on SCARA. For instance, recovering with the 500k samples results in only a 0.5% ARR improvement over 51k samples. This suggests that the resilience provided by SCARA cannot be easily compromised by simply increasing the dataset scale. Details are reported in Appendix C.6.
5.3 Analysis on SCARA
In this subsection, we assess the effectiveness of SCARA on four LLMs, showing the resilience transition in LLMs and the strong correlation between recovery difficulty and ARR.
Figure 6: (a) shows customizability and resilience changes in Llama2-7B and Phi-2 when different closed-sourced layer sets. Each point for customizability represents hidden layers from the start up to the indicated set. (b) presents $\Delta$ARR for different closed-sourced parameter quantities and proportions in Llama2-7B. (c) depicts recovery ratio changes ($\Delta$RR) in specific functions of Llama2-7B with varying closed-sourced parameter ratios.
Transition Layer. We investigate the existence of transition layers in resilience against model recovery and customizability in the math domain. Results are in Figure 6(a) and detailed in Appendix C.7
For resilience, we hide same-sized layer sets with different start points, and attack them using FT-all. Specifically, the sets consist of one layer for Llama2-7B and two layers for Phi-2. We observe clear resilience transitions in both models. For example, in Llama2-7B, the transition occurs at the eighth layer set. For sets preceding the eighth, $\Delta$ARR remains close to zero, indicating that concealing any set up to the eighth layer set yields resilience comparable to the fully-closed approach. However, concealing sets beyond the eighth leads to decreased resilience, as indicated by increasing $\Delta$ARR values. In the smaller Phi-2 model, the transition happens earlier at the first layer set, with hiding later sets resulting in reduced resilience. These patterns suggest resilience transitions occur deeper in larger models. Similar trends of transitions in Mistral-7B and Phi-1.5 can be found in Appendix B.8.
For customizability, we conceal varying numbers of layers from the start and fine-tune the open set. We observe the customizability transition in the 7B model. Specifically in Llama2-7B, concealing only the first layer set improves performance by 43%, but this gain drops to 7% when concealing up to the third layer set and eventually fluctuates around 0%. However, we barely observe transitions in the smaller model, with improvements limited to 3%. This suggests that fewer closed-sourced layers enhance customizability, and larger models possess greater customizability than smaller ones.
Resilience vs. Closed-source Size. We explore how closed-sourced set size impacts the model resilience in Llama2-7B. We start from the smallest self-attention module (k_projection) in a decoder layer and extend to the entire model. Results are shown in Figure 6 and detailed in Appendix C.8.
In the general domain (Figure 6(b)), we observe that resilience emerges when a sufficient number of parameters are closed-sourced. Specifically in Llama2-7B, we observe that when the proportion of closed-sourced increases to 3%, resilience emerges. Furthermore, we observe similar patterns in specific domains, yet resilience transitions vary across capabilities as shown in Figure 6(c). For instance, technical skills such as Math show earlier transitions, with resilience emerging at 1% parameters closed-sourced, whereas domains such as Commonsense Reasoning require hiding 3%. In summary, closed-sourcing a small portion of parameters can provide sufficient resilience against model recovery, meanwhile, technical capabilities tend to be more challenging to recover than other domains. Other models are analyzed in Appendix B.9.
Figure 7: (a) presents the Pearson coefficient between recovery ratio (RR) and recovery difficulty (RD) across four models and six domains. (b) and (c) depict the link between $\Delta$ARR and RD for Llama2-7B and OPT-350M.
5.4 discussions
Effectiveness of RD on large models. We assess the efficacy of the recovery difficulty (RD) in estimating the performance of the recovered model. Specifically, we calculate the Pearson and Spearman correlation coefficients between RD and ARR across different capability groups. As shown in Figure 7(a), we observe a negative correlation between the recovery difficulty and average recovery ratio. For example, in Llama2-7B, the Pearson coefficient is consistently below -0.80, reaching as low as -0.98. We observe similar phenomena in other models with varying architectures and sizes, confirming RD as a reliable predictor of recovered model performance and the effectiveness of SCARA. Further analysis and results of Spearman coefficients can be found in Appendix B.10.
[…]
Figure 8: Sensitivity to $\varepsilon$.
Table 5: Performance of SCARA defending adversarial attacks. $\downarrow$ indicates the smaller the better.
[…]
Gold Std. 58.0 43.9 0.00
SCARA 72.3 85.0 26.5
SAP-DP 72.2 83.9 24.9
[…]
Limited Defense against Adversarial Attack. We compare SCARA and SAP-DP in defending against three black-box adversarial attacks on Llama2-7B. Specifically, we apply the membership inference (Fu et al., 2023) (MIA), attribute inference (Staab et al., 2023) (AIA), and prompt injection (Liu et al., 2023) (PIA) attacks to the semi-open models produced by SAP-DP and SCARA. As shown in Table 5, we observe that SAP-DP outperforms SCARA across all three attacks, but still performs worse than the gold standard. This is because SCARA does not introduce additional output perturbation and thus provide limited defense against black-box adversarial attacks. Details can be found in Appendix B.11.
6 Conclusion
In this paper, we explored finding minimal closed-sourced sets to enhance LLM customizability while preserving their resilience against semi-open model recovery attacks. We theoretically prove that minor errors in bottom decoder layers prior to a transition layer greatly reduce recovery attack success. We introduced SCARA, which selectively closed-sources a small set of layers, achieving superior customizability and comparable resilience to SAP-DP and fully-closed. We empirically investigated the existence of customization and resilience transitions, showed the impact of closed-source size on model resilience, analyzed the effectiveness of our approach, and finally discussed its limitations.
Appendix A Proof of Theorem 1
In this section, we prove Theorem 1. We first revisit the our model, present several important lemmas and finally present the proof.
A.1 Model Overview
The recovered model $f(\mathbf{X};\bm{\theta})$ is structured as a sequence of $L$ transformer layers,
$f(\mathbf{X})=\varphi_{L}\circ\varphi_{L-1}\circ...\circ\varphi_{\alpha L+1}\circ\hat{\varphi}_{\alpha L}\circ_{\alpha L-1}\circ...\circ\varphi_{1}(\mathbf{X}),$ (3)
where $\mathbf{X}\in\mathbb{R}^{n\times d}$ represents the input, interpreted as an assembly of $n$ tokens, each possessing $d$ hidden dimensions. Each transformer layer, indexed by $1\leq i\leq L$, is represented by $\varphi_{i}$, which maps $\mathbb{R}^{n\times d}$ to $\mathbb{R}^{n\times d}$ and can be defined as follows,
$\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)=\left[\mathbf{I}_{n}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\right]\mathbf{X},$ (4)
where $Q_{i}\in\mathbb{R}^{d\times d_{Q}}$, $K_{i}\in\mathbb{R}^{d\times d_{Q}}$ represent projection parameter matrices. Here, the $\alpha L$-th layer is the recovered layer and the others are the public layers. For simplicity, we use the function $\hat{\varphi}_{\alpha L}$ to denote mapping of the recovered layer, i.e., $\hat{\varphi}_{\alpha L}(\mathbf{X})=\varphi_{\alpha L}(\mathbf{X};\hat{K}_{\alpha L},\hat{Q}_{\alpha L})$.
A.2 Bounds on Different Orthogonal Components
Lemma 1.
For any $1\leq l\leq L$, $1\leq p\leq d$, any $\mathbf{X}\in\mathbb{R}^{n\times d}$, we have
$\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\varphi_{l}\left(\mathbf{X};K_{l},Q_{l}\right)[p]\right|\leq(1+\beta_{D})\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\mathbf{X}[p]\right|,$ (5)
where $\mathbb{I}_{n}$ is a column vector with dimensions $n\times 1$ and each element is 1, $\mathbf{X}[p]$ is the $p$-th column of the input $\mathbf{X}$, $\varphi_{l}\left(\mathbf{X};K_{l},Q_{l}\right)[p]$ is the $p$-th column of the $l$-th self-attention output, the coefficient $\beta_{D}$ satisfies $10<\beta_{D}<$ 1 and it is related to the upper bound of the L2-norm of matrices $K_{l},Q_{l}$.
Proof.
Let $\bm{u}=\left\{\bm{u}_{l,1}=\frac{\mathbb{I}_{n}}{\sqrt{n}},\bm{u}_{l,2},\ldots,\bm{u}_{l,n}\right\}$ denote the eigenvectors of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. Assume $\sigma_{l,1},\sigma_{l,2},\ldots,\sigma_{l,n}$ denote the eigenvalues of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$ and $-1<\sigma_{l,n}<\beta_{D}$ for any $l,n$. Thus we have
$\bm{v}^{\top}\varphi_{l}\left(\mathbf{X};K_{l},Q_{l}\right)[p]$ $=\bm{v}^{\top}\left[\mathbf{I}_{n}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\right]\mathbf{X}[p]$ (6a)
$=\bm{v}^{\top}\left[\mathbf{I}_{n}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\right]\sum_{k=1}^{n}\alpha_{pk}\bm{u}_{l,k}$ (6b)
$=\bm{v}^{\top}\sum_{k=1}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{u}_{l,k}$ (6c)
$\leq\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{v}^{\top}{u}_{l,k}\right|$ (6d)
$=\left\|\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{u}_{l,k}\right\|_{2}$ (6e)
$=\left[\sum_{k=2}^{n}\alpha^{2}_{pk}(1+\sigma_{l,k})^{2}\right]^{1/2}$ (6f)
$\leq(1+\beta_{D})\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\mathbf{X}[p]\right|,$ (6g)
where
$\beta_{D}=\max_{\left\|K_{l}\right\|_{2}\leq D,\left\|Q_{l}\right\|_{2}\leq D}\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}{\left\|\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\bm{v}\right\|_{2}}<1.$
The equation equation 6c is due to $\bm{u}_{l,k}$ are the eigenvectors of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. The inequality equation 6e is because when $\bm{v}=\frac{\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{u}_{l,k}}{\left\|\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k}){u}_{l,k}\right\|_{2}}$, we have the maximum value.
Lemma 2.
For any $K_{l},Q_{l}\in\mathbb{R}^{d\times s}$ and any $\mathbf{X}\in\mathbb{R}^{n\times d}$, the following equation always holds:
$\left|\mathbb{I}_{n}^{\top}\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]\right|=2\left|\mathbb{I}_{n}^{\top}\mathbf{X}[p]\right|,$ (7)
where $\mathbf{X}[p]$ is the $p$-th column of the input $\mathbf{X}$, $\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]$ is the $p$-th column of the $l$-th self-attention output.
Proof.
Assume that a set of orthogonal basis for $\mathbb{R}^{n}$ is $\left\{{\bm{u_{1}},\bm{u_{2}},\dots,\bm{u_{n}}}\right\}$, where $\bm{u_{1}}=\frac{\mathbb{I}_{n}}{\sqrt{n}}$. Then we can rewrite $\mathbf{X}[p]$ as $\mathbf{X}[p]=\sum_{j=1}^{n}\alpha_{pj}\bm{u_{j}}$, where $\alpha_{pj}$$(1\leq p\leq d)$ are the corresponding coefficients for the $p$-th column of $\mathbf{X}$ under the orthogonal basis. Next, we calculate $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|$ and $\left|\mathbb{I}_{n}^{\top}\mathbf{X}[p]\right|$, respectively. Note that $\mathbb{I}_{n}^{\top}\bm{u_{j}}=0$ for all $j\neq 1$. Therefore, we can obtain that,
$\mathbb{I}_{n}^{\top}\mathbf{X}[p]=\sqrt{n}\alpha_{p1}.$ (8)
Then we can get
$\left|\mathbb{I}_{n}^{\top}\mathbf{X}[p]\right|=|\sqrt{n}\alpha_{p1}|.$ (9)
Let $\sigma_{i1},\sigma_{i2},\ldots,\sigma_{in}$ denote the eigenvalues of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. Applying the Perron–Frobenius theorem for Markov matrices Lemmens & Nussbaum (2012), we deduce that for the matrix $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$, there exists only one eigenvalue equal to 1, while all other eigenvalues in absolute value are strictly less than 1. Without loss of generality, we assume $\sigma_{i1}=1$, implying $\left|\sigma_{ij}\right|<1$ for $j\neq 1$. Recalling the definition of $\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)$ and considering the linear operation, we can rewrite it as follows:
$\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]=\sum_{j=1}^{n}\alpha_{pj}\left(1+\sigma_{ij}\right)\bm{u_{j}}.$ (10)
Then we calculate the term $\left|\mathbb{I}_{n}^{\top}\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]\right|$ as follows,
$\left|\mathbb{I}_{n}^{\top}\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]\right|$ $=\left|\mathbb{I}_{n}^{\top}(\sum_{j=1}^{n}\alpha_{pj}\left(1+\sigma_{ij}\right)\bm{u_{j}}\right|$ (11a)
$=\left|\sqrt{n}\left(\alpha_{p1}(1+\sigma_{i1})\right)\right|$ (11b)
$=2|\sqrt{n}\alpha_{p1}|,$ (11c)
where equation 11a is induced by substituting the equation equation 10 into $\left|\mathbb{I}_{n}^{\top}\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]\right|$, equation 11b is due to $\mathbb{I}_{n}^{\top}\bm{u_{j}}=0$ for all $j\neq 1$, equation 11c follows the fact that $\sigma_{i1}=1$
A.3 Proof of Theorem 1
We first prove the following result. For simplicity of notations, we use $f(\mathbf{X})\left[p\right]$ to denote the $p$-th ($1\leq p\leq d$) column of the the recovered model $f(\mathbf{X})$, where the parameters in the $\alpha L$-th layer is replaced with the matrices $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$. We use the function $\hat{\varphi}_{\alpha L}(\mathbf{X})=\varphi_{\alpha L}(\mathbf{X};\hat{K}_{\alpha L},\hat{Q}_{\alpha L})$ to denote the mapping of the $(\alpha L)$-th layer. Then we are going to show that there exists $\alpha^{\star}=\log_{2}{\frac{2}{1+\beta_{D}}}$ and $10<\beta_{D}<$ 1 makes the following equations hold.
(1) Assume $\alpha<\alpha^{\star}$. For any $\mathbf{X}$, $\left\|K_{i}\right\|_{2}\leq D,\left\|Q_{i}\right\|_{2}\leq D$, there exists a zero measure set $\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}(\mathbf{X})$ such that
$\lim_{L\to\infty}\left\|\frac{f(\mathbf{X})\left[p\right]}{\left\|f(\mathbf{X})\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}=0.$ (12)
(2) For any $\alpha>\alpha^{\star}$, there exists a sequence of matrix $\{K_{i},Q_{i}\}_{i\geq 1}$ such that for any recovered matrix $K_{\alpha L}$ and $Q_{\alpha L}$, we have $\left\|K_{i}\right\|_{2}\leq D,\left\|Q_{i}\right\|_{2}\leq D$, we have,
$\lim_{L\to\infty}\left\|\frac{f(\mathbf{X})\left[p\right]}{\left\|f(\mathbf{X})\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}=\sqrt{2}.$ (13)
Proof.
Based on Lemma equation 1, we obtain that
$\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}f\left(\mathbf{X}\right)[p]\right|\leq(1+\beta)^{L}\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\mathbf{X}[p]\right|.$ (14)
Based on Lemma equation 2, we know that
$\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=2^{(1-\alpha)L-1}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|.$ (15)
We firstly prove the equation equation 12. When
$\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=2^{(1-\alpha)L-1}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|\neq 0,$ (16)
then we have
$\left\|\frac{f(\mathbf{X})\left[p\right]}{\left\|f(\mathbf{X})\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}$ $=\left[2-\frac{2\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]}{\sqrt{n}\sqrt{\frac{(\mathbb{I}_{n}^{\top}f(\mathbf{X})[p])^{2}}{n}+(\bm{v}^{\top}f(\mathbf{X})[p])^{2})}}\right]^{1/2}$ (17a)
$=\sqrt{2}\left[1-\frac{1}{\sqrt{1+\frac{n(\bm{v}^{\top}f(\mathbf{X})[p])^{2}}{(\mathbb{I}_{n}^{\top}f(\mathbf{X})[p])^{2}}}}\right]^{1/2}$ (17b)
$\leq\sqrt{2}\left[1-\frac{1}{\sqrt{1+\frac{n(1+\beta)^{2L}|\bm{v}^{\top}\mathbf{X}[p]|^{2}}{2^{2[(1-\alpha)L-1]}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|^{2}}}}\right]^{1/2}$ (17c)
$\leq 2\sqrt{2n}\left(\frac{1+\beta}{2^{1-\alpha}}\right)^{L}\frac{\left|\bm{v}^{\top}\mathbf{X}[p]\right|}{\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|},$ (17d)
where the inequality equation 17c is based on the inequality equation 14 and equation 15. The inequality equation 17d is based on Lemma equation 3. Therefore, if $\alpha<\log_{2}{\frac{2}{1+\beta_{D}}}$ and $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|\neq 0$, then we have $\lim_{L\to\infty}\left(\frac{1+\beta_{D}}{2^{1-\alpha}}\right)^{L}=0$. Now we can consider when $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=0.$ In fact, it is easy to show that this can only happens when $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ belong to certain sets making $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=0,$ which corresponds to zero measure set $\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}(\mathbf{X})$ depending on the input $\mathbf{X}$. Since the input space is countable, therefore, the union $\cup_{\mathbf{X}\in\mathcal{X}}\mathcal{K}(\mathbf{X})$ and $\cup_{\mathbf{X}\in\mathcal{X}}\mathcal{Q}(\mathbf{X})$ are also zero-measure sets.
To prove equation equation 13, let $K^{\star}$, $Q^{\star}$ with $\left\|K^{\star}\right\|_{2}\leq D,\left\|Q^{\star}\right\|_{2}\leq D$ satisfy the following condition,
$\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}{\left\|\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\bm{v}\right\|_{2}}=\beta_{D}.$ (18)
Let $\bm{v}^{\star}$ be the solver of the above optimization problem equation 18 and consider the $K_{l}=K^{\star}$, $Q_{l}=Q^{\star}$ and $\mathbf{X}^{\star}=[\bm{v}^{\star},\bm{v}^{\star},\cdots,\bm{v}^{\star}]$. Clearly, $\bm{v}^{\star}\perp\mathbb{I}_{n}.$ Assume there exists $\bm{u}:\left\|\bm{u}^{\star}\right\|_{2}=1$ satisfying $\bm{u}^{\star}\perp\mathbb{I}_{n}$, $\bm{u}^{\star}\perp\bm{v}^{\star}$, therefore we can rewrite $f(\mathbf{X}^{\star})\left[p\right]$ as follows,
$f(\mathbf{X}^{\star})\left[p\right]=\frac{\mathbb{I}_{n}^{\top}}{\sqrt{n}}f(\mathbf{X}^{\star})\frac{\mathbb{I}_{n}}{\sqrt{n}}+\bm{v}^{\star\top}f(\mathbf{X}^{\star})\bm{v}^{\star}+\bm{u}^{\star\top}f(\mathbf{X}^{\star})\bm{u}^{\star}.$ (19)
For any $1\leq l\leq L$, based on Lemma equation 1, we know that
$\left|\bm{v}^{*\top}f\left(\mathbf{X}^{\star}\right)[p]\right|=(1+\beta_{D})^{L}\left|\bm{v}^{*\top}\mathbf{X}^{\star}[p]\right|.$ (20)
Since
$\left|\mathbb{I}_{n}^{\top}f\left(\mathbf{X}^{\star}\right)[p]\right|=2^{L}\left|\mathbb{I}_{n}^{\top}\mathbf{X}^{\star}[p]\right|=|\mathbb{I}_{n}^{\top}\bm{v}^{\star}|=0$ (21)
and
$\left|\bm{v}^{*\top}f\left(\mathbf{X}^{\star}\right)[p]\right|=(1+\beta_{D})^{L}\left|\bm{v}^{*\top}\mathbf{X}^{\star}[p]\right|\neq 0,$ (22)
then we have
$\left\|\frac{f(\mathbf{X}^{\star})\left[p\right]}{\left\|f(\mathbf{X}^{\star})\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}$ $=\left[2-\frac{2\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p]}{\sqrt{n}\left\|f(\mathbf{X}^{\star})\left[p\right]\right\|_{2}}\right]^{1/2}$ (23a)
$=\left[2-\frac{2\mathbb{I}_{n}^{\top}}{\sqrt{n}}\frac{f(\mathbf{X}^{\star})[p]}{\sqrt{\frac{1}{n}(\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p])^{2}+(\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p])^{2}+(\bm{u}^{\star\top}f(\mathbf{X}^{\star})[p])^{2}}}\right]^{1/2}$ (23b)
$\geq\left[2-\frac{2\mathbb{I}_{n}^{\top}}{\sqrt{n}}\frac{f(\mathbf{X}^{\star})[p]}{\sqrt{\frac{1}{n}(\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p])^{2}+(\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p])^{2}}}\right]^{1/2}$ (23c)
$=\left[2-2\frac{\frac{\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p]}{\sqrt{n}\left|\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p]\right|}}{\sqrt{1+\frac{\left|\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p]\right|^{2}}{n\left|\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p])\right|^{2}}}}\right]^{1/2}$ (23d)
$=\left[2-2\frac{\frac{2^{(1-\alpha)L-1}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X}^{\star})[p]\right|}{\sqrt{n}(1+\beta_{D})^{L}\left|\bm{v}^{\star\top}\mathbf{X}^{\star}[p]\right|}}{\sqrt{1+\frac{2^{2[(1-\alpha)L-1]}}{n(1+\beta_{D})^{2L}}\frac{\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X}^{\star})[p]\right|^{2}}{\left|\bm{v}^{\star\top}\mathbf{X}^{\star}[p]\right|^{2}}}}\right]^{1/2},$ (23e)
where equation equation 23b is based on equation 19, equation equation 23e is based on equation 22 and equation 15. When $\alpha>\log_{2}{\frac{2}{1+\beta_{D}}}$, we have $\lim_{L\to\infty}\left(\frac{2^{1-\alpha}}{1+\beta_{D}}\right)^{L}=0$. Thus we have $\lim_{L\to\infty}\left\|\frac{f(\mathbf{X}^{\star})\left[p\right]}{\left\|f(\mathbf{X}^{\star}\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}=\sqrt{2}$. This indicates that the $p$-th column of the output matrix $f(\mathbf{X}^{\star})$ is not parallel to $\mathbf{I}_{n}$ for any $p$. This further indicates that the output matrix does not have the identical vector in each row. ∎
A.4 Technical Lemma
Lemma 3.
For any $x\in(0,1)$, it always holds $\left[1-\frac{1}{\sqrt{1+x^{2}}}\right]^{1/2}\leq x$.
Proof.
To establish the inequality $\left[1-\frac{1}{\sqrt{1+x^{2}}}\right]^{1/2}\leq x$, we begin by proving,
$1-\frac{1}{\sqrt{1+x^{2}}}\leq x^{2}.$ (24)
To demonstrate equation 24, we equivalently show
$1-x^{2}\leq\frac{1}{\sqrt{1+x^{2}}}.$ (25)
Subsequently, it suffices to verify
$(1-x^{2})(\sqrt{1+x^{2}})\leq 1.$ (26)
This is equivalent to proving
$(1-x^{2})^{2}(1+x^{2})\leq 1.$ (27)
Thus, our focus shifts to demonstrating
$(1-x^{2})(1-x^{4})\leq 1.$ (28)
Clearly, equation 28 holds true for any $x\in(0,1)$. ∎
Appendix B Experiment Details
B.1 Model Details.
The foundation models we use in our experiments are selected from open-source repositories, and Table 6 shows the basic information of the models and their sources. Specifically, we employ Llama2-70B-chat [^2] https://huggingface.co/meta-llama/Llama-2-70b-chat-hf, Llama2-7B-chat [^3] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, and Mistral-7B-v0.1444https://huggingface.co/mistralai/Mistral-7B-v0.1. For smaller models, we select Phi-2555https://huggingface.co/microsoft/phi-2 and Phi-1.5666https://huggingface.co/microsoft/phi-1_5. We also consider OPT model [^7] https://huggingface.co/facebook/opt-350m, which has only 350 million parameters and 24 decoder layers.
Table 6: Model Info
Model Size Decoder Layers
Llama2-70B-chat (Touvron et al., 2023) 70B 80
Llama2-7B-chat (Touvron et al., 2023) 7B 32
Mistral-7B-v0.1 (Jiang et al., 2023a) 7B 32
Phi-2 (Abdin et al., 2024) 2.7B 32
Phi-1.5 (Li et al., 2023) 1.3B 24
OPT (Zhang et al., 2022) 350M 24
B.2 Recovery Attacks.
Attack implementation details. In performing FT-all and FT-closed model recovery attacks, we adhere to the training hyper-parameters outlined in the Llama2 report (Touvron et al., 2023), employing the AdamW optimizer with a cosine learning rate scheduler. The initial learning rate is set to $2\times 10^{-5}$, with a weight decay of 0.1, a batch size of 128, and bfloat16 precision for input sequences of 512 tokens. The LLaMA2-70B model is trained for 3 epochs with a random seed of 42, while other models are trained for 5 epochs across three seeds: 42, 1234, and 20. Despite limiting training to 3 epochs for the 70B model, the training loss stabilized effectively. Our implementation builds upon the llama-recipes repository provided by META.
For SEM attacks, distinct configurations were employed for SCARA and SAP-DP. In the case of SCARA, embeddings from the closed-source components were collected and paired with the input data to train a substitute model. In contrast, for SAP-DP, embeddings from the sixth decoder layer and the model’s final logits were utilized to construct the training dataset. In accordance with (Tamber et al., 2024), we applied a learning rate of 1.5e-4, a weight decay of 0.01, and a linear learning rate scheduler with 500 warmup steps. Both training and validation batch sizes were set to 32, with MSE as the loss function. SCARA was trained for 30 epochs due to its smaller model size, whereas SAP-DP was trained for 5 epochs.
All recovery experiments were conducted on Nvidia 4090 24G, 6000 Ada 48G, and A100 80G GPUs, utilizing PyTorch 2.2.0 and CUDA 11.8 on Ubuntu 20.04.6 LTS.
Base 51k Recovery Dataset. We ensure dataset coverage and reliability by using a 1:1 ratio of the MMLU auxiliary training set  [^8] https://github.com/hendrycks/test and Alpaca dataset  [^9] https://github.com/tatsu-lab/stanford_alpaca/blob/main/alpaca_data.json, extracting 25.5k samples from each. From the MMLU auxiliary training data (Hendrycks et al., 2021), we sample 50%, and from Alpaca (Taori et al., 2023), we use a step size of 2 to enhance diversity. The datasets are then formatted for model training, applying Alpaca and MMLU prompts from Table 7.
Table 7: Prompts for Alpaca and MMLU auxiliary training data
Dataset Prompt Type Description
Alpaca with input Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request.
w/o input Below is an instruction that describes a task. Write a response that appropriately completes the request.
MMLU Question Answering Below is a question with no choices. Write the correct answer that appropriately solves the question.
Multiple Choice The following is a multiple choice question, paired with choices. Answer the question in the format: “Choice:content”.
Extra Recovery Datasets. To enhance dataset diversity, the 100K, 200K, 300K, and 500K datasets integrate additional specialized sources. As detailed in Table 8, these sources include Baize (Xu et al., 2023) (158K English multi-turn conversations via ChatGPT’s self-chat), MathInstruct (Yue et al., 2023) (260K curated math instruction instances focusing on hybrid reasoning), and OpenOrca (Mukherjee et al., 2023) (augmented FLAN collection with 1M GPT-4 completions and 3.2M GPT-3.5 completions). These enrichments are intended to support complex computational and theoretical tasks, offering broader topic coverage.
Table 8: Composition of variously sized datasets
Raw Data Set 51k 100k 200k 300k 500k
Alpaca 25.5 50 40 50 50
MMLU auxiliary training set 25.5 50 40 100 100
Baize-MedQuAD 0 0 40 50 50
Baize-Quora 0 0 40 50 50
Baize-Stackoverflow 0 0 40 50 50
MathInstruct 0 0 4 6 20
OpenOrca 0 0 0 0 180
Validation Datasets. Table 9 outlines the composition of the validation datasets. For Validation Dataset 1, we extracted 50% from each of the 57 MMLU validation sub-datasets, totaling 1.5K instances, paired with Alpaca data selected using a step size of 751. This dataset is used with the 51K and 100K training sets. For larger training sets (200K, 300K, and 500K), Validation Dataset 2 was created by adding 400 instances from three Baize subsets, expanding the validation set to 4.0K.
Table 9: Composition of validation datasets of different sizes
Raw Data Set Validation Set Evaluation Set
Alpaca 765 765
MMLU auxiliary training set 751 751
Baize-MedQuAD 0 850
Baize-Quora 0 850
Baize-Stackoverflow 0 850
Total Length 1516 4066
B.3 Baselines.
In this section, we provide further details on the baselines used in our comparisons: SAP-DP and fully-closed. These schemes represent different strategies, each with distinct trade-offs in terms of customizability and resilience against model recovery attacks.
SAP. The Split-and-Privatize (SAP) framework (Shen et al., 2023) offers an approach to balance between protecting model privacy and data privacy while maintaining competitive performance. Specifically, the SAP framework keeps the bottom six encoder layers open, allowing user access and fine-tuning while closing the deeper layers on the vendor.
SAP-DP. To further strengthen protection while maintaining competitive performance, we extend SAP by incorporating differential privacy techniques by adding Laplace noise to perturb the logits during the fine-tuning process (Lee et al., 2018). The Laplace Distribution with mean $\mu$ and scale $b$ is the distribution with probability density function:
$\text{Laplace}(x|\mu,b)=\frac{1}{2b}\exp\left(-\frac{|x-\mu|}{b}\right)$
Specifically, in SAP-DP, the noise $n$ is sampled: $n\sim\text{Laplace}(0,0.5)$ and added to the output logits of the model to balance privacy protection and model performance.
Fully-closed. Following (Eiras et al., 2024), we use the fully-closed approach as a baseline. This assumes the adversary has no access to internal model parameters, treating the model as a black-box, where only output data can be collected. We slightly broaden this setup by assuming the adversary knows the model’s architecture but no other details. Thus, recovering the fully-closed model involves using the collected data to retrain a model with the same architecture to restore its general functionality.
B.4 Implementation Details of SCARA.
Evaluation Datasets. We created a 1.5K Evaluation Set to assess model resilience under various closed-sourcing strategies. This set includes 50% of entries from each of the 57 MMLU validation sub-datasets (Hendrycks et al., 2021), distinct from Validation Set outlined in Table 9. Additionally, we selected an equal number of Alpaca dataset (Taori et al., 2023), using a step size of 751, ensuring no overlap with the Validation Set.
Hyper-parameter Sensitivity. As shown in Figure 9, we evaluate SCARA’s sensitivity to tolerance magnitude $\varepsilon$, adjusting it from 0.05 to 1 in 0.05 increments while calculating the $\Delta\text{ARR}$ for six recovered models. The results indicate that SCARA is minimally sensitive to changes in $\varepsilon$, with $\Delta\text{ARR}$ values stabilizing as $\varepsilon$ increases. This stability arises from the need for a smaller closed-sourced layer at higher $\varepsilon$, allowing the condition $R(I)\leq(1+\varepsilon)R([L])$ to be met with fewer layers. Additionally, the increase in $\Delta\text{ARR}$ is smaller for larger models, suggesting that privatizing more parameters beyond a certain point offers diminishing returns in resilience.
Figure 9: Sensitivity on $\varepsilon$.
B.5 Evaluation Benchmarks
Most of our evaluations are conducted using the lm-evaluation suite (Gao et al., 2023), the bigcode-evaluation-harness platform (Ben Allal et al., 2022), and MT-Bench (Zheng et al., 2023). For specific domains, such as finance and law, we utilize the official benchmark testing codes provided by their respective communities, as detailed below.
Evaluation on Customizabilities. We assess the customizability of models across six domains, as detailed in Table 10. Each domain includes specific benchmarks and metrics designed to evaluate different aspects of the model’s performance in relation to customizability. In particular, for evaluating medical capabilities, we select two subcategories from the MMLU benchmark that are related to the medical domain: mmlu_anatomy and mmlu_professional_medicine. For assessing legal reasoning, we select 10 multiple-choice and judgment-based subcategories from Legalbench. The performance of the model in these legal tasks is measured using perplexity, following the prompt structure provided by Legalbench. Specifically, the selected subcategories include: cuad_audit_rights, canada_tax_court_outcomes, definition_classification, cuad_affiliate_license-licensee, learned_hands_business, contract_nli_survival_of_obligations, contract_nli_explicit_identification, contract_nli_confidentiality_of_agreement, hearsay, and contract_qa.
Table 10: Details of the Six Customizability Benchmarks
Domain Benchmark Metric n-shot Reference
Code HumanEval Pass@1 0 Chen et al. (2021)
MBPP Pass@1 1 Austin et al. (2021)
Math GSM8K Exact Match 8 Cobbe et al. (2021)
Medical MMLU_Medical Accuracy 5 Hendrycks et al. (2021)
Finance FPB F1 0 Wang et al. (2023a)
Law LegalBench Accuracy 0 Guha et al. (2023)
Alignment MT-Bench Score (GPT-4) Zheng et al. (2023)
Evaluation on Resilience. We follow the Llama-2 report Touvron et al. (2023) to evaluate the recovered model, including 16 benchmarks, which are categorized into 6 groups. Table 11 summarizes the functionality benchmarks used in our experiments, along with their test methods and performance metrics. Our model ranks choices in multiple-choice tasks and generates answers for open-ended generation tasks.
Table 11: Details of the Sixteen Functionality Benchmarks
Domain Benchmark Metric n-shot Reference
Commonsense Reasoning PIQA Accuracy 0 Bisk et al. (2020)
Hellaswag Accuracy 0 Zellers et al. (2019)
Winogrande Accuracy 0 Sakaguchi et al. (2019)
ARC_easy Accuracy 0 Clark et al. (2018)
ARC_challenge Accuracy 0 Clark et al. (2018)
Reading Comprehension OpenBookQ Accuracy 0 Mihaylov et al. (2018)
LAMBADA Accuracy 0 Paperno et al. (2016)
BoolQ Accuracy 0 Clark et al. (2019)
SQuADv2 HasAns_EM 2 Rajpurkar et al. (2018)
SQuADv2 HasAns_F1 2 Rajpurkar et al. (2018)
World Knowledge NaturalQuestions Exact Match 5 Kwiatkowski et al. (2019)
TriviaQA Exact Match 5 Joshi et al. (2017)
Code HumanEval Pass@1 0 Chen et al. (2021)
MBPP Pass@1 1 Austin et al. (2021)
Math GSM8K Exact Match 8 Cobbe et al. (2021)
General Ability MMLU Accuracy 5 Hendrycks et al. (2021)
BBH Accuracy 3 Suzgun et al. (2022)
B.6 Model Customization
Datasets. To fine-tune the models for domain-specific tasks, we utilized several datasets tailored to different sectors, including Code (Zheng et al., 2024b), Math (Yue et al., 2023), Medical (Zhang et al., 2023), Finance (Wang et al., 2023b), Law (Guha et al., 2024), and Alignment (Meng et al., 2024). Table 12 lists the customization training datasets used in the experiments. For the code domain, we combine the datasets from CodeFeedback and CodeAlpaca. For law and finance, we merge all training datasets from Legalbench and FinGPT respectively. These datasets are then prepared for model training using the Alpaca prompts outlined in Table 7. Additionally, we randomly select 3,000 samples to serve as the validation dataset.
Table 12: Customization Training Datasets Composition
Domain Dataset Name Size Reference
Code CodeFeedback 156k  Zheng et al. (2024a)
CodeAlpaca 20k Chaudhary (2023)
Math MathInstruction 262K Yue et al. (2023)
Medical MedMCQA 183k Zhang et al. (2023)
Law Legalbench 90k Guha et al. (2023)
Finance FinGPT 204k Wang et al. (2023a)
Alignment Ultrafeedback 62k Cui et al. (2024)
Customization Training Hyperparameters. In model customization, we use different hyperparameters depending on the model size. For LLaMA2-70B, we apply QLoRA with the settings outlined in Table 13, while for 7B models, we use LoRA. For smaller models like Phi2 and Phi-1.5, we fine-tune all model parameters. For LLaMA2-70B, we fine-tune it as a quantized 4-bit model over 1 epoch, starting with a learning rate of $1.5\times 10^{-6}$. For the 7B models, we train for 3 epochs, with a seed value of 42. The training setup includes a weight decay of 0.1, a batch size of 128, a warmup ratio of 0.03, and input sequences of 512 tokens, following standard experimental practices (Hu et al., 2021). For Phi2 and Phi-1.5, we use the training hyperparameters from the LLaMA2 report. We employ the AdamW optimizer with a cosine learning rate scheduler, starting with a learning rate of $2\times 10^{-5}$, a weight decay of 0.1, a batch size of 128, and use bfloat16 precision for 512-token input sequences. Specifically, for alignment, we follow SimPO Meng et al. (2024) and set the preference parameters $\beta=2$ and $\gamma=1$. The learning rate is $1\times 10^{-6}$ for LLaMA2-70B and $5\times 10^{-7}$ for the 7B and smaller models. All experiments are conducted using the LLaMA-Factory on Nvidia 4090 24G, 6000 Ada 48G, and A100 80G GPUs, with PyTorch 2.2.0 and CUDA 11.8 on Ubuntu 20.04.6 LTS.
Table 13: The Hyperparameters for Customization Training.
Model Method Rank $r$ Lora $\alpha$ Dropout Learning Rate Epochs Warmup R.
Llama2-70B QLoRA 96 16 0.05 1.50E-04 1 0.03
Llama2-7B LoRA 32 64 0.05 2.00E-05 3 0.03
Mistral-7B LoRA 32 64 0.05 1.00E-06 3 0.03
B.7 Amplification of small recovery error
To investigate the amplification of minor recovery errors in the pre-transition layers, we conducted experiments on the LLaMA2-7B model. Specifically, we added Gaussian noise, $x\sim N(0,0.01^{2})$, to the parameters of the first decoder layer of the model. Then, we compared the embedding outputs generated by the noisy model with those from the original model. The difference between the two sets of outputs was measured using the Frobenius norm. We collected embedding outputs from layers 0, 3, 6, 9, 12, 15, 18, 21, 24, 27, and 30, totaling 45,000 output samples. The results showed a significant amplification of these small errors.
B.8 Resilience and Customization Transitions
For the LLaMA2-7B model, the smallest closed-source layer set identified by SCARA consists of a single decoder layer, whereas for Phi-2, it includes two decoder layers. Consequently, for LLaMA2-7B, we opted to closed-source each even-indexed layer, while for Phi-2, we chose to closed-source non-overlapping pairs of layers (e.g., layers 0-1, 2-3). For each selected layer set, we first closed-source them, then subjected the semi-open model to FT-all attacks, and subsequently calculated the $\Delta\text{ARR}$ of the layer set to assess its resilience.
When verifying the customization transition, due to computational constraints, we validated only every other layer set for both models (e.g., closed-source layers 0, 0-4, 0-8 …). Specifically, we applied LoRA-based customization on LLaMA2-7B in the math domain, while for Phi-2, we utilized the full finetuning approach. The experimental hyperparameters remain consistent with those outlined in the Appendix B.6.
We further computed the $\Delta\text{ARR}$ for each closed-source set within Mistral-7B-v0.1 and Phi-1.5. In these models, the smallest closed-source set identified by SCARA consists of one decoder layer and two decoder layers, respectively. Following the same experimental configuration as LLaMA2-7B and Phi-2, we closed-sourced each even-indexed layer for Mistral-7B, and non-overlapping pairs of layers for Phi-1.5. The complete results demonstrating the transition layers within the Mistral-7B and Phi-1.5 model that closed two non-overlapping consecutive layers are depicted in Figure 10. Once again, we observed a distinct presence of transition layers. Specifically, in Mistral-7B, the transition layer appears at the 24th layer, while in Phi-1.5, it is located within the first layer set. Further results for can be found in Appendix C.7.
Figure 10: Resilience changes in Miatral-7B and Phi-1.5.
B.9 Resilience Across Closed Sizes
To examine the influence of Closed layer size on model resilience, we conduct experiments on Closed-sourcing different amounts and proportions of parameters in the model’s decoder layer. We give instructions on the detailed setting of closed-sourced models in Table 14. The module names are all derived from the overall implementation functions of each model in the Transformers open-source repositories in Table 6. We utilize abbreviated module names to denote specific settings.
Table 14: Closed-sourced Sizes Setting. “*” indicates an entire decoder layer.
Llama-7B Mistral-7B Phi2-2.7B Phi1.5-1.3B
0.25% $\mathrm{W}_{k}$ $\mathrm{W}_{q},\mathrm{W}_{k}$ $\mathrm{W}_{k}$ $\mathrm{W}_{k}$
0.50% $\mathrm{W}_{q},\mathrm{W}_{k}$ $\mathrm{W}_{o},MLP_{\text{up}}$ $\mathrm{W}_{q},\mathrm{W}_{k}$ $\mathrm{W}_{q},\mathrm{W}_{k}$
1% $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{o}$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{o}$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{d}$ $\mathrm{W}_{q},\mathrm{W}_{k},,\mathrm{W}_{v}$
3% 0 0 0 0
7% 0-1 0-1 0-1 0-1
15% 0-4 0-4 0-3 0-3
30% 0-9 0-9 0-9 0-6
50% 0-15 0-15 0-15,$\mathrm{W}_{em}$ 0-11,$\mathrm{W}_{em}$
Proportion 100% Fully-closed Fully-closed Fully-closed Fully-closed
20M $\mathrm{W}_{k}$ $\mathrm{W}_{q},\mathrm{W}_{k}$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v}$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{d}$
50M $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v}$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{o}$ $MLP$ 0
100M $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},MLP$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{o},MLP$ 0, $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v}$ 0-1
160M $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{o},MLP$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{o},MLP$ 0-1 0-2
200M 0 0 0-1, $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{d},MLP_{\text{f1}}$ 0-3
300M 0, $\mathrm{W}_{q},\mathrm{W}_{v},\mathrm{W}_{o},MLP_{\text{up}}$ 0, $\mathrm{W}_{q},\mathrm{W}_{v},\mathrm{W}_{o},MLP_{\text{up}}$ 0-3 0-5
Quantity 600M 0-2 0-2 0-7 0-11
We further computed $\Delta\text{ARR}$ by close-sourcing varying quantities and proportions of parameters under FT-all attacks on three additional models. As shown in Figure 11 and Figure 6(b), we observed the same pattern as with Llama2-7B, where resilience emerges once a sufficient number of parameters are closed-sourced. For example, on Mistral-7B, resilience occurs after closed-sourcing 100 million parameters, which is less than a single decoder layer. Closed-sourcing fewer parameters leads to a notable drop in resilience, with $\Delta\text{ARR}$ rising to around 40%. Beyond this threshold, resilience stabilizes near 0% $\Delta\text{ARR}$. This pattern holds across all models, highlighting a critical threshold for effective closed-source. Furthermore, different architectures require varying closed-sourcing quantities to achieve resilience, even with similar model sizes. For instance, Mistral-7B reaches resilience by closed-sourcing 100 million parameters, Llama2-7B requires 200 million, and Phi-1.5 needs a higher rate of 7%, compared to 3% for Llama2-7B.
Figure 11: $\Delta$ ARR for different closed parameter quantities and proportions.
B.10 Effectiveness of recovery difficulty
The complete Pearson and Spearman results are presented in Table 15, revealing a negative correlation between RS and the average recovery ratio. For example, in Llama2-7B, both Pearson and Spearman coefficients fall below -0.77. Similar trends are seen in models with varying architectures and sizes, confirming that RD is a reliable predictor of recovered model performance and demonstrating the effectiveness of SCARA. Additionally, Figure 12 shows scatter plots depicting the relationship between $\Delta$ARR and Recovery Difficulty($\uparrow$)s across four models, along with the corresponding Pearson and Spearman correlation coefficients. The Recovery Difficulty($\uparrow$)s were obtained from Section 5.3. As illustrated in Figure 12, we observe a clear trend: an increase in $\Delta$ARR corresponds to a decrease in model scores across all models analyzed. This inverse relationship is consistently supported by strong negative values for both Pearson and Spearman correlation coefficients, with the most significant negative correlation seen in Phi2-2.7B, indicating a substantial drop in model scores as $\Delta$ARR increases.
Table 15: Correlation coefficients (Spearman | Pearson) between recovery ratio and recovery difficult.
Model Rsn. Read. Knl. Code & Math Gen. Avg.
Llama2-7B -0.83 |||| -0.97 -0.77 |||| -0.96 -0.83 |||| -0.95 -0.85 |||| -0.90 -0.82 |||| -0.93 -0.80 |||| -0.98
Mistral-7B -0.83 |||| -0.89 -0.72 |||| -0.91 -0.82 |||| -0.94 -0.78 |||| -0.95 -0.55 |||| -0.87 -0.67 |||| -0.92
Phi-2 -0.93 |||| -0.96 -0.84 |||| -0.96 -0.74 |||| -0.87 -0.84 |||| -0.80 -0.84 |||| -0.84 -0.87 |||| -0.95
Phi-1.5 -0.86 |||| -0.97 -0.78 |||| -0.94 -0.83 |||| -0.94 -0.90 |||| -0.80 -0.84 |||| -0.89 -0.80 |||| -0.94
Figure 12: Correlation Analysis of $\Delta$ARR and Recovery Difficulty Across Different Models.
B.11 Adversarial Attack
In this section, we provide a detailed comparison of SCARA and SAP-DP in their effectiveness against three types of black-box adversarial attacks on the Llama2-7B model. The attacks considered include Membership Inference Attacks (MIA), Attribute Inference Attacks (AIA), and Prompt Injection Attacks (PIA).
Membership Inference Attack (MIA): This attack aims to determine whether a specific data point was included in the training dataset of the model. Attackers utilize model outputs to infer membership status, potentially exposing sensitive information about the training data (Fu et al., 2023; Chen & Pattabiraman, 2024). We conducted our experiment following SPV_MIA  [^10] https://github.com/wjfu99/MIA-LLMs, which provides a robust framework for assessing model vulnerabilities. We focus on the AUC scores for SPV-MIA against semi-open models across Ag News datasets (Zhang et al., 2016).
Attribute Inference Attack (AIA): In this scenario, the adversary attempts to infer specific attributes of training data based on the model’s outputs. This can lead to privacy breaches, particularly when sensitive attributes are involved (Staab et al., 2023; Li et al., 2024). We conducted our experiments following the methodology outlined in Staab et al. (2023)  [^11] https://github.com/eth-sri/llmprivacy and evaluated the top-3 accuracy on the PersonalReddit (PR) Dataset.
Prompt Injection Attack (PIA): This attack manipulates input prompts to coerce the model into producing desired outputs that may compromise the integrity or security of the system (Zhao et al., 2024; Xu et al., 2024). In our experiment, we follow AutoDAN  [^12] https://github.com/SheltonLiu-N/AutoDAN, which can automatically generate stealthy jailbreak prompts by the carefully designed hierarchical genetic algorithm.We evaluate the effectiveness of these prompts using the keyword-based attack success rate (ASR), which measures the presence of predefined keywords in responses generated LLMs. For gold standard, LED  [^13] https://github.com/ledllm/ledllm, significantly enhances the resilience of LLMs against prompt injection attacks (PIA), reducing the ASR to 0.
Appendix C Detailed Results
C.1 Comparison in two semi-open Llama2-70B
In this experiment, we examine two semi-open Llama2-70B models, where either the first two decoder layers are closed-source (referred to as Semi-Open-1) or the last two decoder layers are closed-source (referred to as Semi-Open-2). The objective is to compare their performance in terms of customization and their resilience under the recovery attack. The results are summarized in Table 16 and Table 17.
Table 16: Customization Performance of Llama2-70B under Different Closed-Sourced Layers
Math Code Medical Law Finance Alignment
Fully Closed-sourced 53.15 24.90 53.68 79.63 37.54 7.19
Semi-Open-1 62.40 43.99 62.73 93.85 87.51 7.46
Semi-Open-2 62.53 42.36 62.72 93.91 87.90 7.46
Table 17: Recovery Performance of Llama2-70B under Different Closed-Sourced Layers
Benchmarks Fully Closed-sourced Semi-Open-1 Semi-Open-2
Rsn. PIQA 50.82 50.49 79.05
winogrande 51.07 51.22 72.93
arc_easy 25.17 25.63 76.30
arc_challenge 23.55 20.48 50.17
Hellaswag 26.65 25.77 79.49
Read. lambada 0.00 0.01 57.25
BoolQ 43.30 37.92 84.95
SQuADv2_EM 0.00 0.00 1.54
SQuADv2_f1 0.23 1.01 35.59
OBQA 25.60 24.40 44.00
Knl. NQ 0.00 0.00 15.18
TriviaQA 0.00 0.00 52.67
Code mbpp 0.00 0.00 16.00
HumanEval 0.00 0.00 13.41
Math GSM8K 0.03 0.01 27.75
Gen. MMLU 23.01 23.22 63.61
BBH 0.00 0.00 49.45
Average Recovery Ratio($\downarrow$) 22.55 21.73 74.94
C.2 Evaluation Results under FT-all attack
In this section, we provide a comprehensive analysis of the evaluation results, comparing SCARA with two baseline methods: SAP-DP and a fully-closed approach. This comparison is conducted across 16 benchmarks under the FT-all attack scenario. The detailed results for Llama2-70B are presented in Table 18, while the results for Llama2-7B and Mistral-7B are shown in Table 19. Additionally, the outcomes for Phi-2 and Phi-1.5 are provided in Tables 20.
Table 18: Evaluation results of Llama2-70B under FT-all attack
Pre-train SCARA SAP-DP Fully-closed
Rsn. PIQA 80.69 50.49 48.26 50.82
Winogrande 74.74 51.22 50.59 51.07
ARC-easy 80.35 25.63 26.35 25.17
ARC-challenge 53.24 20.48 20.31 23.55
Hellaswag 82.15 25.77 25.76 26.65
Read. LAMBADA 75.07 0.01 0.00 0.00
BoolQ 86.70 37.92 37.83 43.30
SQuADv2_EM 51.23 0.00 0.00 0.00
SQuADv2_f1 67.43 1.01 1.13 0.23
OBQA 44.80 24.40 24.40 25.60
Knl. NaturalQuestions 32.38 0.00 0.00 0.00
TriviaQA 73.47 0.00 0.02 0.00
Code MBPP 24.80 0.00 0.00 0.00
HumanEval 25.00 0.00 0.00 0.00
Math GSM8K 53.15 0.01 0.00 0.03
Gen. MMLU 63.09 23.22 24.19 23.01
BBH 61.40 0.00 0.00 0.00
Average Recovery Ratio($\downarrow$) - 21.73 21.64 22.55
Table 19: Evaluation results of 7B models under FT-all attack
Llama2-7B Mistral-7B
SCARA SAP-DP Fully-closed SCARA SAP-DP Fully-closed
Rsn. PIQA 49.56 49.56 49.47 51.63 50.22 49.35
Winogrande 50.99 49.66 50.83 49.78 51.07 50.59
ARC-easy 27.04 26.43 25.98 26.12 28.03 25.83
ARC-challenge 21.07 20.56 22.47 19.94 21.42 22.35
Hellaswag 25.56 25.69 26.39 26.10 25.97 25.39
Read. LAMBADA 0.01 0.00 0.01 0.12 0.00 0.01
BoolQ 44.30 41.70 48.34 39.05 37.83 45.80
SQuADv2_EM 0.00 0.00 0.00 0.00 0.00 0.00
SQuADv2_f1 0.49 0.63 0.59 1.21 0.26 0.66
OBQA 25.13 23.00 25.93 25.60 25.20 25.00
Knl. NaturalQuestions 0.01 0.01 0.04 0.00 0.00 0.02
TriviaQA 0.00 0.00 0.02 0.00 0.00 0.01
Code MBPP 0.00 0.00 0.00 0.00 0.00 0.00
HumanEval 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 24.26 22.92 24.45 25.24 23.05 23.26
BBH 0.00 0.00 0.00 0.00 0.00 0.00
Average Recovery Ratio($\downarrow$) 25.03 24.16 25.62 22.41 22.28 22.68
Table 20: Evaluation results of small models under FT-all attack
Phi-2 Phi-1.5
SCARA SAP-DP Fully-closed SCARA SAP-DP Fully-closed
Rsn. PIQA 54.17 52.01 52.07 53.43 52.61 50.44
Winogrande 51.56 48.93 48.91 51.09 49.25 49.12
ARC_easy 34.57 28.20 27.03 30.81 28.79 27.50
ARC_challenge 19.45 19.37 18.66 20.56 19.80 21.22
Hellaswag 27.61 25.32 25.26 26.27 25.66 25.05
Read. LAMBADA 0.75 0.02 0.00 0.59 0.00 0.00
BoolQ 45.29 40.21 44.60 46.98 41.80 46.28
SQuADv2_EM 0.02 0.00 0.00 0.00 0.00 0.00
SQuADv2_f1 2.61 0.28 0.64 0.78 0.65 1.60
OBQA 24.80 26.60 25.80 26.60 28.60 26.87
Knl. NaturalQuestions 0.00 0.00 0.02 0.04 0.00 0.00
TriviaQA 0.01 0.00 0.01 0.01 0.00 0.00
Code MBPP 0.00 0.00 0.00 0.00 0.00 0.00
HumanEval 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 24.16 22.87 22.95 24.07 22.95 22.95
BBH 0.01 0.00 0.00 0.00 0.00 0.00
C.3 Customization Performance of Models
In this section, we present detailed evaluation results of the model customization performance across six downstream tasks used in our experiments. The detailed results for Llama2-70B are presented in Table 21, while the results for Llama2-7B and Mistral-7B are shown in Table 22 and Table 23. Additionally, the outcomes for Phi-2 and Phi-1.5 are provided in Tables 24 and Table 25.
Table 21: Detailed results of Llama2-70B closed by SCARA on six downstream tasks.
Math Code Medical Law Finance Alignment
Fully-Closed 53.15 24.90 53.68 79.63 55.63 7.19
SAP-DP 61.10 36.87 54.55 83.40 65.78 7.41
SCARA 62.40 43.99 62.73 93.85 87.51 7.46
Fully-Open 64.06 44.58 63.40 94.17 88.22 7.42
Table 22: Detailed results of Llama2-7B closed by SCARA on six downstream tasks.
Math Code Medical Law Finance Alignment
Fully-Closed 20.24 13.75 36.91 51.80 38.71 6.51
SAP-DP 20.24 13.75 36.91 51.80 38.71 6.52
SCARA 28.96 21.37 46.52 90.84 81.95 6.63
Fully-Open 29.34 21.265 47.60 90.49 84.09 6.63
Table 23: Detailed results of Mistral-7B closed by SCARA on six downstream tasks.
Math Code Medical Law Finance Alignment
Fully-Closed 38.21 33.83 61.50 50.47 37.39 3.20
SAP-DP 41.47 34.44 63.08 50.37 38.10 2.47
SCARA 46.10 43.16 66.78 84.94 86.19 3.87
Fully-Open 45.26 46.08 66.47 88.13 84.91 3.78
Table 24: Detailed results of Phi-2 closed by SCARA on six downstream tasks.
Math Code Medical Law Finance Alignment
Fully-Closed 57.77 47.59 43.13 56.46 54.07 5.22
SAP-DP 58.52 46.65 43.40 56.81 54.37 5.11
SCARA 59.59 47.79 45.85 57.11 56.26 5.26
Fully-Open 59.60 48.40 45.93 57.19 56.68 5.27
Table 25: Detailed results of Phi-1.5 closed by SCARA on six downstream tasks.
Math Code Medical Law Finance Alignment
Fully-Closed 30.33 35.09 30.78 52.18 34.60 3.24
SAP-DP 30.25 35.45 32.66 51.99 34.27 3.68
SCARA 33.66 37.10 33.14 52.26 39.60 3.87
Fully-Open 34.49 37.45 33.23 52.34 39.90 3.68
C.4 Comparison in closing baselines on llama2-70B
We compare the recovery resilience of SCARA with SAP-DP and Fully-closed as baselines under FT-closed and SEM attack strategies. The evaluation results on sixteen benchmarks are shown in Table 26.
Table 26: Evaluation results of Llama2-70B under FT-closed and SEM attack
FT-closed SEM
SCARA SAP-DP SCARA SAP-DP
Rsn. PIQA 49.78 49.40 48.62 49.00
Winogrande 51.30 49.01 50.99 51.13
ARC-easy 26.43 25.59 25.33 24.55
ARC-challenge 21.41 21.42 22.01 20.93
Hellaswag 26.07 26.10 25.90 25.22
Read. LAMBADA 0.00 0.00 0.00 0.00
BoolQ 45.09 37.83 44.95 39.80
SQuADv2_EM 0.00 0.00 0.00 0.00
SQuADv2_f1 0.98 1.01 0.59 1.00
OBQA 24.40 23.80 25.03 22.96
Knl. NaturalQuestions 0.00 0.00 0.00 0.00
TriviaQA 0.00 0.00 0.00 0.00
Code MBPP 0.00 0.00 0.00 0.00
HumanEval 0.00 0.00 0.00 0.00
Math GSM8K 0.00 0.00 0.00 0.00
Gen. MMLU 23.18 23.66 22.98 22.83
BBH 0.00 0.00 0.00 0.00
Average Recovery Ratio($\downarrow$) 22.60 21.80 22.40 22.30
C.5 Comparison in Recovery Attack Strategies
In this section, we present detailed evaluation results of the model recovery performance of SCARA under FT-closed and SEM attack strategies across six functionalities used in our experiments. The detailed results under the FT-closed recovery strategy are presented in Table 27. The results under SEM attack strategies are shown in Table 28.
Table 27: Recovery Performance of SCARA under FT-Closed attacks.
Llama2-7B Mistral-7B Phi-2 Phi-1.5
Rsn. PIQA 49.95 49.55 54.57 52.45
Winogrande 49.88 49.68 52.33 52.41
ARC-easy 27.65 25.88 33.33 31.06
ARC-challenge 20.81 22.69 19.03 18.77
Hellaswag 26.04 25.01 27.62 26.88
Read. LAMBADA 0.00 0.00 0.77 0.71
BoolQ 38.13 46.01 44.34 57.49
SQuADv2_EM 0.00 0.00 0.00 0.00
SQuADv2_f1 0.22 0.36 3.07 2.27
OBQA 25.70 25.12 24.40 25.20
Knl. NaturalQuestions 0.00 0.00 0.00 0.00
TriviaQA 0.00 0.00 0.01 0.00
Code MBPP 0.00 0.00 0.00 0.00
HumanEval 0.00 0.00 0.00 0.00
Math GSM8K 0.00 0.00 0.00 0.00
Gen. MMLU 24.23 23.56 23.03 24.10
BBH 0.00 0.00 0.00 0.00
Average Recovery Ratio($\downarrow$) 24.80 22.50 23.56 26.97
Table 28: Recovery Performance of SCARA under SEM attacks.
Llama2-7B Mistral-7B Phi-2 Phi-1.5
Rsn. PIQA 51.52 48.53 49.46 50.82
Winogrande 50.28 51.02 48.70 50.59
ARC-easy 24.83 25.83 25.93 24.62
ARC-challenge 24.99 22.35 20.65 21.08
Hellaswag 25.58 25.39 25.84 25.39
Read. LAMBADA 0.00 0.01 0.00 0.01
BoolQ 53.30 45.80 38.41 61.07
SQuADv2_EM 0.00 0.00 0.00 0.00
SQuADv2_f1 0.77 0.66 0.00 1.35
OBQA 25.00 25.00 27.80 30.40
Knl. NaturalQuestions 0.00 0.02 0.00 0.00
TriviaQA 0.00 0.01 0.01 0.00
Code MBPP 0.00 0.00 0.00 0.00
HumanEval 0.00 0.00 0.00 0.00
Math GSM8K 0.00 0.00 0.00 0.00
Gen. MMLU 25.39 23.26 22.95 23.11
BBH 0.00 0.00 0.00 0.00
Average Recovery Ratio($\downarrow$) 25.00 22.00 22.10 24.70
C.6 Comparison in Recovery datasets scales
To investigate the impact of attack dataset scales on the efficiency of SCARA, we conduct model recovery attack on the Llama2-7B model using four different attack datasets of varying sizes: 100k, 200k, 300k, and 500k. The evaluation performance under different attack set scales are in Table 29
Table 29: Evaluation Results of SCARA on Llama2-7B under Various Attack Set Scales.
51K 100K 200K 300K 500K
Rsn. PIQA 49.56 49.89 49.18 49.18 49.59
Winogrande 50.99 47.99 49.49 50.20 50.20
ARC-easy 27.04 27.06 27.06 27.02 27.01
ARC-challenge 21.07 21.33 20.90 21.16 21.48
Hellaswag 25.56 26.49 26.46 26.50 26.19
Read. LAMBADA 0.01 0.01 0.00 0.00 0.01
BoolQ 44.30 44.41 44.10 44.07 44.96
SQuADv2_EM 0.00 0.00 0.00 0.02 0.00
SQuADv2_f1 1.05 0.32 0.51 0.52 0.71
OBQA 25.13 25.00 23.80 25.20 25.60
Knl. NaturalQuestions 0.01 0.08 0.08 0.06 0.06
TriviaQA 0.00 0.02 0.01 0.03 0.01
Code MBPP 0.00 0.00 0.00 0.00 0.00
HumanEval 0.00 0.00 0.00 0.00 0.00
Math GSM8K 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 24.26 25.34 25.43 26.14 26.41
BBH 0.00 0.00 0.00 0.00 0.00
Average Recovery Ratio($\downarrow$) 25.07 25.03 24.89 25.26 25.48
C.7 Transition Layer Results.
Resilience Performance. We close same-sized layer sets with different start points, and attack them using FT-all. Specifically, the sets consist of one layer for Llama2-7B (Table 30, Table 31), and two layers for Phi-2 (Table 34, Table 35). We further computed the $\Delta\text{ARR}$ for each closed-source set within Mistral-7B-v0.1 and Phi-1.5 in Appendix B.8. The results for the Mistral-7B-v0.1 model are presented in Table 32 and Table 33. Additionally, the performance outcomes for the Phi-1.5 model can be found in Table 36.
In all the above tables, “Pretrain” represents the model’s original performance without any layers closed-sourced. These columns indicate the index of layers in the model that have been closed-sourced. “*” indicates fully-closed. All evaluation scores are averages from three different seed tests, corresponding to the values 20, 42, and 1234, following the details of the Sixteen Functionality Benchmarks in Appendix B.5.
Table 30: Evaluation Results of Llama2-7B under Different Closed Layers (Part1)
Pretrain 0 1 2 3 4 5 6 7
Rsn. PIQA 76.66 49.56 51.43 49.53 50.45 49.84 50.27 50.96 51.09
Hellaswag 75.45 25.56 25.75 25.88 26.16 25.91 27.20 29.39 28.89
Winogrande 66.38 50.99 50.86 50.15 49.75 49.96 50.91 51.64 51.36
ARC_easy 74.41 27.04 27.23 26.10 26.30 25.51 26.44 28.24 27.96
ARC_challenge 44.11 21.07 20.31 20.19 21.30 22.04 21.56 20.62 22.92
Read. OpenBookQA 68.49 0.01 0.11 0.02 0.02 0.01 0.00 0.05 0.04
LAMBADA 80.67 44.30 41.22 38.36 41.43 38.08 38.14 38.40 41.55
BoolQ 59.48 0.00 0.04 0.00 0.00 0.00 0.00 0.01 0.03
SQuADv2_em 71.88 1.05 1.31 0.63 1.07 0.45 0.44 1.13 1.10
SQuADv2_f1 43.80 25.13 24.60 23.60 24.93 25.67 24.47 25.07 26.00
Knl. NaturalQuestions 22.47 0.01 0.00 0.01 0.03 0.02 0.01 0.13 0.08
TriviaQA 57.23 0.00 0.01 0.00 0.02 0.01 0.01 0.07 0.10
Code HumanEval 10.90 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 16.60 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 20.24 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 45.83 24.26 25.37 23.98 24.26 24.75 24.01 25.23 27.45
BBH 39.86 0.00 0.00 0.00 0.00 0.00 0.00 0.50 0.38
Avg. Performance Score($\downarrow$) 51.44 15.82 15.78 15.20 15.63 15.43 15.50 15.97 16.41
Average Recovery Ratio($\downarrow$) - 30.76 30.67 29.55 30.39 29.99 30.13 31.04 31.90
Recovery Difficulty($\uparrow$) - 11.11 11.27 10.87 10.31 10.83 10.33 10.90 11.11
Table 31: Evaluation Results of Llama2-7B under Different Closed-sourced Layers (Part2). “*” indicates the fully closed-sourced model.
16 18 20 22 24 26 28 30 *
Rsn. PIQA 51.47 52.99 58.22 65.83 69.60 73.45 75.46 75.99 49.47
Hellaswag 31.38 36.55 45.61 56.60 62.70 67.88 71.37 72.94 26.39
Winogrande 53.09 55.98 58.96 64.12 64.80 65.25 65.46 66.53 50.83
ARC_easy 30.58 35.35 43.85 55.92 62.56 68.36 70.85 72.60 25.98
ARC_challenge 24.26 26.85 30.97 35.38 38.17 41.41 43.00 44.17 22.47
Read. OpenBookQA 0.28 1.58 6.79 30.88 44.58 56.23 62.33 63.11 0.01
LAMBADA 57.55 70.53 71.36 78.85 79.69 80.29 79.39 80.40 48.34
BoolQ 0.08 0.90 2.34 7.07 6.04 6.87 3.54 9.46 0.00
SQuADv2_em 2.21 13.48 21.47 35.72 36.96 39.32 37.08 42.08 0.59
SQuADv2_f1 27.33 28.20 30.47 32.13 34.93 39.27 39.93 41.53 25.93
Knl. NaturalQuestions 0.13 0.41 1.60 2.94 4.29 2.69 7.28 11.87 0.04
TriviaQA 0.25 1.79 4.93 11.02 15.73 17.95 33.19 42.26 0.02
Code HumanEval 0.00 0.00 0.00 0.00 0.00 3.25 8.34 10.98 0.00
MBPP 0.00 0.00 0.00 0.07 0.47 2.27 8.80 13.27 0.00
Math GSM8K 0.00 0.00 0.00 0.13 0.81 8.42 6.90 15.77 0.00
Gen. MMLU 43.17 48.20 49.38 49.58 49.72 50.03 50.75 50.61 24.45
BBH 0.76 11.44 19.79 28.87 31.16 35.98 38.24 40.54 0.00
Avg. Performance Score($\downarrow$) 18.97 22.60 26.22 32.65 35.42 38.76 41.29 44.36 16.15
Average Recovery Ratio($\downarrow$) 36.89 43.94 50.98 63.48 68.87 75.35 80.27 86.24 31.39
Recovery Difficulty($\uparrow$) 10.42 9.49 8.86 7.12 6.14 4.72 3.40 3.06 11.19
Table 32: Evaluation Results of Mistral-7B under Different Closed-sourced Layers (Part1)
Pretrain 0 1 2 3 4 5 6 7
Rsn. PIQA 81.99 51.63 53.20 53.63 53.47 51.56 52.61 50.71 55.15
Hellaswag 81.04 26.10 26.36 26.36 26.66 27.10 25.51 26.18 28.16
Winogrande 74.03 49.78 49.78 51.01 50.38 49.91 50.14 49.70 51.17
ARC_easy 80.77 33.03 31.96 30.71 29.66 30.25 30.35 26.44 32.38
ARC_challenge 50.26 19.94 21.27 20.45 19.60 20.05 21.36 21.25 20.73
Read. OpenBookQA 44.40 25.60 25.20 25.20 25.47 25.87 26.33 25.07 27.20
LAMBADA 73.29 0.12 0.44 1.91 2.08 0.80 0.30 0.17 1.95
BoolQ 83.67 39.05 53.12 45.95 38.61 47.35 38.06 46.44 47.66
SQuADv2_em 64.04 0.00 0.00 0.01 0.01 0.01 0.00 0.00 0.01
SQuADv2_f1 71.37 1.21 0.84 1.05 1.03 1.27 0.43 0.07 0.86
Knl. NaturalQuestions 28.98 0.00 0.01 0.00 0.04 0.01 0.00 0.02 0.07
TriviaQA 70.79 0.00 0.00 0.02 0.01 0.01 0.01 0.00 0.16
Code HumanEval 29.88 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 38.40 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 38.21 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 62.50 25.24 24.68 25.11 23.43 23.65 24.26 24.26 24.99
BBH 56.40 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.01
Avg. Performance Score($\downarrow$) 60.59 15.98 16.87 16.55 15.91 16.34 15.84 15.90 17.09
Average Recovery Ratio($\downarrow$) - 26.38 27.85 27.32 26.25 26.97 26.15 26.24 28.20
Average Recovery Ratio($\downarrow$) - 11.50 11.31 11.48 10.71 10.77 11.44 11.02 10.71
Table 33: Evaluation Results of Mistral-7B under Different Closed-sourced Layers (Part2)
16 18 20 22 24 26 28 30 *
Rsn. PIQA 54.50 52.32 52.72 57.13 62.82 64.67 67.23 75.61 49.35
Hellaswag 29.31 29.02 29.99 33.46 46.21 52.12 52.46 67.73 25.39
Winogrande 51.20 54.17 51.07 55.75 58.59 62.41 63.09 66.33 50.59
ARC_easy 32.84 29.35 30.80 38.04 47.24 51.99 54.74 69.95 25.83
ARC_challenge 21.19 23.04 23.78 26.34 30.86 33.22 35.04 40.53 22.35
Read. OpenBookQA 26.00 27.87 26.87 29.67 28.73 32.67 33.40 36.40 25.00
LAMBADA 2.61 0.18 1.28 4.17 21.89 29.93 24.49 48.32 0.01
BoolQ 53.98 53.60 58.79 55.76 64.10 74.72 68.48 81.30 45.80
SQuADv2_em 0.01 0.00 0.47 0.13 2.39 3.59 1.87 1.82 0.00
SQuADv2_f1 0.96 0.18 1.27 2.60 14.88 22.61 21.12 34.16 0.66
Knl. NaturalQuestions 0.01 0.10 0.19 0.58 1.84 3.15 3.53 8.87 0.02
TriviaQA 0.03 0.01 0.61 0.62 5.14 7.51 10.32 25.44 0.01
Code HumanEval 0.00 0.00 0.00 0.61 2.24 4.88 2.44 9.75 0.00
MBPP 0.00 0.00 0.00 2.00 4.33 8.33 0.93 13.07 0.00
Math GSM8K 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.25 0.00
Gen. MMLU 24.30 25.84 29.54 24.55 34.77 40.77 40.84 50.44 23.26
BBH 0.00 0.00 0.02 0.30 7.55 18.76 21.05 30.07 0.00
Avg. Performance Score($\downarrow$) 17.47 17.39 18.08 19.51 25.51 30.08 29.47 38.83 15.78
Average Recovery Ratio($\downarrow$) 28.83 28.71 29.84 32.20 42.09 49.64 48.64 64.08 26.05
Average Recovery Ratio($\downarrow$) 11.34 11.11 10.45 10.59 10.23 10.34 9.59 8.53 11.20
Table 34: Evaluation Results of Phi-2 under Different Closed-sourced Layers (Part 1)
Pretrain 0 2 4 6 8 10 12 14
Rsn. PIQA 79.27 54.17 72.85 73.76 75.03 76.75 78.00 78.91 77.84
Hellaswag 73.73 27.61 56.49 57.73 60.47 62.84 66.39 66.91 66.95
Winogrande 75.45 51.56 59.17 59.98 59.88 64.32 68.11 68.95 70.38
ARC_easy 79.92 34.57 72.94 73.40 73.97 76.51 78.33 78.66 78.63
ARC_challenge 52.90 19.45 41.75 39.82 44.11 45.65 47.92 49.74 48.78
Read. OpenBookQA 51.20 25.80 35.73 37.47 40.13 42.00 44.00 45.67 44.80
LAMBADA 56.28 3.25 28.55 30.42 34.64 40.05 45.41 45.52 46.66
BoolQ 83.36 47.29 65.20 62.64 66.39 71.39 73.42 72.95 75.83
SQuADv2_em 61.30 0.02 10.49 17.63 21.94 33.94 19.54 19.15 29.14
SQuADv2_f1 71.38 2.61 37.22 40.35 45.53 59.16 48.21 50.09 54.87
Knl. NaturalQuestions 9.58 0.00 3.60 4.97 6.13 7.55 7.95 8.10 9.25
TriviaQA 39.29 0.01 13.57 16.29 24.74 28.60 31.58 33.71 32.79
Code HumanEval 48.78 0.00 1.42 6.50 10.98 16.66 22.76 19.51 23.17
MBPP 46.80 0.00 5.07 6.87 9.47 19.60 25.67 23.47 25.73
Math GSM8K 57.77 0.00 7.25 8.64 4.42 9.63 14.18 11.35 17.31
Gen. MMLU 56.73 26.16 34.29 37.01 39.90 43.11 45.63 48.17 49.82
BBH 59.53 0.01 15.27 18.37 16.38 14.58 4.93 4.35 11.37
Avg. Performance Score($\downarrow$) 59.02 17.21 32.99 34.81 37.30 41.90 42.47 42.66 44.90
Average Recovery Ratio($\downarrow$) - 29.15 55.90 58.99 63.21 71.00 71.97 72.28 76.09
Recovery Difficulty($\uparrow$) - 10.07 7.07 4.95 4.09 3.63 3.31 3.31 3.11
Table 35: Evaluation Results of Phi-2 under Different Closed-sourced Layers (Part2). “*” indicates the fully closed-sourced model.
16 18 20 22 24 26 28 30 *
Rsn. PIQA 77.44 77.80 77.69 76.77 76.89 77.55 78.16 78.58 52.07
Hellaswag 67.20 66.90 67.13 68.00 68.86 70.01 71.44 71.18 25.26
Winogrande 70.82 71.40 73.11 74.46 75.79 75.72 75.93 74.77 48.91
ARC_easy 78.30 77.27 77.33 76.82 78.09 77.76 79.53 79.56 27.03
ARC_challenge 49.71 48.29 48.52 48.04 49.80 50.68 53.16 52.67 18.66
Read. OpenBookQA 46.53 46.47 45.87 45.27 46.33 45.53 46.53 48.27 20.80
LAMBADA 45.67 46.88 47.95 50.17 50.54 52.77 53.01 53.23 0.00
BoolQ 80.56 80.72 82.22 83.31 83.98 83.54 82.54 83.41 39.60
SQuADv2_em 7.88 1.30 1.69 1.31 0.15 0.23 3.54 10.03 0.56
SQuADv2_f1 40.84 34.51 34.25 35.94 35.64 36.68 39.57 44.87 0.90
Knl. NaturalQuestions 8.90 6.09 6.40 6.79 6.86 6.85 7.20 8.37 0.02
TriviaQA 31.48 27.03 25.08 24.54 22.89 22.99 24.24 26.93 0.01
Code HumanEval 22.56 21.34 25.41 32.52 38.01 46.14 46.54 43.90 0.00
MBPP 26.73 25.33 24.80 31.73 36.67 41.80 43.13 43.20 0.00
Math GSM8K 16.68 16.02 14.66 12.31 17.24 30.12 45.41 49.79 0.00
Gen. MMLU 52.69 53.45 55.68 56.61 56.93 56.59 56.86 56.47 22.95
BBH 3.42 17.36 8.33 18.24 30.09 48.12 52.28 56.36 0.00
Avg. Performance Score($\downarrow$) 42.79 42.25 42.12 43.70 45.57 48.42 50.53 51.86 15.10
Average Recovery Ratio($\downarrow$) 72.51 71.58 71.38 74.04 77.22 82.04 85.63 87.87 25.59
Recovery Difficulty($\uparrow$) 3.07 3.29 3.03 3.01 2.70 2.32 1.98 2.13 11.32
Table 36: Evaluation Results of Phi-1.5 under Different Closed-sourced Layers
Pretrain 0-1 2-3 4-5 6-7 8-9 10-11 12-13 14-15 16-17 18-19 20-21 22-23 *
Rsn. PIQA 75.68 53.43 69.52 71.53 73.50 74.76 75.08 74.94 74.64 73.90 74.63 74.54 74.81 50.44
Hellaswag 62.56 26.27 46.66 50.71 52.98 54.51 55.11 56.01 56.78 57.90 58.76 59.35 58.58 25.05
Winogrande 72.69 51.09 54.91 59.22 61.75 64.85 67.95 68.88 68.98 71.25 71.19 72.87 70.66 49.12
ARC_easy 76.14 30.81 61.70 65.70 70.10 71.38 70.01 71.72 71.93 72.34 73.39 74.16 73.74 27.50
ARC_challenge 44.62 20.56 32.85 34.10 38.08 40.05 40.30 39.48 40.87 41.52 42.84 42.58 45.42 21.22
Read. OpenBookQA 48.00 26.60 33.93 35.73 40.40 41.13 40.67 41.73 41.67 40.27 41.33 43.27 45.47 26.87
LAMBADA 44.10 0.59 17.96 26.45 29.37 33.83 33.85 36.46 37.06 37.96 39.98 41.10 40.49 0.00
BoolQ 75.05 46.98 59.12 52.42 57.41 65.68 68.52 63.47 65.12 66.52 73.91 75.17 77.0 46.28
SQuADv2_em 48.01 0.00 5.82 10.94 18.34 13.96 14.70 23.22 16.98 26.05 22.04 20.16 26.86 0.00
SQuADv2_f1 60.84 0.78 24.49 26.04 34.86 32.17 32.36 43.14 38.23 48.03 45.75 45.56 49.62 1.60
Knl. NaturalQuestions 5.46 0.04 1.68 2.73 3.41 3.06 3.21 4.25 4.03 4.06 4.54 4.17 4.45 0.01
TriviaQA 16.94 0.01 5.70 7.77 10.85 11.03 9.11 12.11 11.84 11.86 12.02 12.11 13.19 0.01
Code HumanEval 35.98 0.00 3.05 10.57 12.20 16.26 13.82 17.48 18.70 23.17 29.68 31.91 31.71 0.00
MBPP 35.40 0.00 2.80 7.80 10.93 17.40 16.53 16.13 16.67 22.27 27.33 28.27 28.53 0.00
Math GSM8K 30.33 0.00 0.05 0.73 0.15 0.23 0.75 0.50 2.17 4.98 9.73 17.77 23.45 0.00
Gen. MMLU 42.44 24.07 26.56 28.77 32.51 32.87 36.09 39.42 39.72 43.23 42.51 42.82 43.66 23.95
BBH 28.80 0.00 2.07 3.97 8.38 7.37 2.81 7.79 4.12 10.63 6.94 10.34 11.45 0.00
Avg. Performance Score($\downarrow$) 47.24 16.54 26.40 29.13 32.66 34.15 34.17 36.28 35.85 38.59 39.80 40.95 42.30 15.94
Average Recovery Ratio($\downarrow$) - 35.02 55.90 61.66 69.14 72.29 72.34 76.80 75.90 81.68 84.25 86.69 89.56 33.75
Average Recovery Ratio($\downarrow$) - 10.08 7.18 4.70 3.50 2.93 2.83 2.53 2.36 2.27 2.16 2.06 2.46 9.33
Customizability Performance. We close varying numbers of layers from the start and fine-tune the open set, and then we observe the customizability transition in models. Table 37 shows the detailed evaluation results of Llama2-7B and Phi-2 on GSM8k benchmark.
Table 37: Customization Performance under Different Closed Sets
Llama2-7B Phi-2
Closed Layers GSM8K($\uparrow$) Closed Layers GSM8K($\uparrow$)
Fully-open 29.34 Fully-open 59.60
0 28.96 0-1 59.59
0-4 21.76 0-5 58.60
0-8 21.46 0-9 58.45
0-12 20.85 0-13 55.19
0-16 20.11 0-17 56.25
0-20 21.46 0-21 54.59
0-24 21.44 0-25 55.34
0-28 18.73 0-29 54.59
Fully-Closed 20.32 Fully-Closed 57.77
C.8 Evaluation Results under Different Closed size
In this section, we present a comprehensive evaluation of the model’s performance across sixteen benchmarks utilized in our experiments. The evaluation results for LLaMA2-7B, categorized by varying quantities and proportions of closed-source parameters, are displayed in Table 38 and Table 39, respectively. For the Mistral-7B model, the results are summarized in Table 40 and Table 41. Furthermore, the evaluation outcomes for the Phi-2 model can be found in Tables 42 and Table 43. The performance results for Phi-1.5 are also included in Tables 44 and Table 45 for comparison. For further details regarding the closed-source settings employed in our experiments, please refer to Appendix C.8.
Table 38: Evaluation Results of Llama2-7B under Different Closed-source Proportion
0.25% 0.5% 1% 3% 7% 15% 30% 50% 100%
Rsn. PIQA 77.78 77.69 67.73 49.42 49.55 50.05 49.98 49.31 49.47
Hellaswag 71.40 71.54 52.39 25.74 26.03 26.25 25.67 25.48 26.39
Winogrande 64.64 65.64 54.12 50.38 50.43 49.65 49.59 49.62 50.83
ARC_easy 74.69 75.04 53.82 26.03 26.76 26.46 26.64 26.66 25.98
ARC_challenge 43.66 43.29 26.99 20.16 21.39 19.74 21.44 21.73 22.47
Read. OpenBookQA 63.15 63.62 33.20 0.01 0.00 0.02 0.01 0.01 0.01
LAMBADA 80.66 80.78 62.10 38.22 39.33 43.45 39.39 41.83 48.34
BoolQ 11.39 12.14 5.47 0.00 0.00 0.00 0.00 0.00 0.00
SQuADv2_em 40.24 40.74 32.65 0.78 0.20 0.24 2.09 2.13 0.59
SQuADv2_f1 40.73 40.67 30.47 22.93 23.40 25.53 24.07 23.07 25.93
Knl. NaturalQuestions 7.83 7.89 5.61 0.00 0.01 0.02 0.01 0.00 0.04
TriviaQA 44.29 45.95 18.78 0.00 0.01 0.00 0.00 0.00 0.02
Code HumanEval 11.39 12.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 15.20 15.33 1.00 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 13.22 13.29 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 45.04 45.03 30.90 24.06 24.04 25.01 23.19 23.11 24.45
BBH 37.45 37.51 17.36 0.00 0.00 0.00 0.00 0.00 0.00
Avg. Performance Score($\downarrow$) 43.69 44.01 28.98 15.16 15.36 15.67 15.42 15.47 16.15
Average Recovery Ratio($\downarrow$) 84.94 85.56 56.33 29.48 29.86 30.47 29.97 30.07 31.39
Recovery Difficulty($\uparrow$) 1.96 1.93 8.66 10.87 11.75 11.48 11.65 11.57 11.19
Table 39: Evaluation Results of Llama2-7B under Different Closed-source Quantity
20M 50M 100M 160M 200M 300M 600M
Rsn. PIQA 77.78 73.49 67.55 67.12 49.42 50.36 49.97
Hellaswag 71.40 63.47 51.67 51.27 25.74 25.70 25.78
Winogrande 64.64 57.54 53.07 52.04 50.38 49.28 50.49
ARC_easy 74.69 66.50 51.97 52.11 26.03 26.43 26.29
ARC_challenge 43.66 36.04 26.51 25.99 20.16 20.79 21.70
Read. OpenBookQA 63.15 45.34 30.22 28.75 0.01 0.05 0.01
LAMBADA 80.66 69.47 62.28 62.59 38.22 39.03 40.80
BoolQ 11.39 2.21 4.18 7.24 0.00 0.00 0.01
SQuADv2_em 40.24 33.98 28.98 31.05 0.78 0.74 0.37
SQuADv2_f1 40.73 33.93 29.13 30.00 22.93 23.80 23.53
Knl. NaturalQuestions 7.83 2.98 5.33 5.73 0.00 0.00 0.02
TriviaQA 44.29 15.28 13.71 17.25 0.00 0.00 0.01
Code HumanEval 11.39 0.41 0.00 0.00 0.00 0.00 0.00
MBPP 15.20 6.87 1.00 0.80 0.00 0.00 0.00
Math GSM8K 9.00 0.10 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 45.04 36.15 28.95 29.04 24.06 23.70 23.45
BBH 37.45 28.53 14.99 16.99 0.00 0.00 0.00
Avg. Performance Score($\downarrow$) 43.44 33.66 27.62 28.12 15.16 15.29 15.44
Average Recovery Ratio($\downarrow$) 84.46 65.44 53.69 54.66 29.48 29.72 30.01
Recovery Difficulty($\uparrow$) 1.96 5.48 8.95 9.25 10.87 10.93 10.81
Table 40: Evaluation Results of Mistral-7B under Different Closed-sourced Proportion
0.25% 1% 0.5% 3% 7% 15% 30% 50% 100%
Rsn. PIQA 77.79 74.36 52.16 53.34 52.07 52.19 50.04 50.60 49.35
Hellaswag 71.31 65.50 26.50 26.16 25.92 25.91 25.87 25.61 25.39
Winogrande 67.09 60.32 49.22 51.65 50.01 51.36 51.36 49.65 50.59
ARC_easy 74.52 69.51 29.95 30.82 29.73 30.44 28.20 27.45 25.83
ARC_challenge 42.32 38.40 20.76 20.71 21.10 20.25 22.61 22.47 22.35
Read. OpenBookQA 42.13 34.60 25.13 25.33 26.47 26.07 25.20 25.87 25.00
LAMBADA 55.99 44.36 0.73 1.66 0.96 0.31 0.03 0.02 0.01
BoolQ 78.35 74.06 43.18 42.01 42.09 40.02 38.53 39.91 45.80
SQuADv2_em 13.91 6.97 0.00 0.01 0.00 0.00 0.00 0.00 0.00
SQuADv2_f1 41.13 33.88 1.60 0.93 1.27 0.71 0.99 0.86 0.66
Knl. NaturalQuestions 8.46 5.82 0.03 0.00 0.02 0.03 0.00 0.00 0.02
TriviaQA 34.04 17.03 0.01 0.01 0.02 0.01 0.00 0.00 0.01
Code HumanEval 11.99 6.51 0.00 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 16.93 12.80 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 6.32 0.45 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 44.17 37.98 23.98 24.34 25.10 23.91 23.68 24.12 23.26
BBH 35.44 27.27 0.02 0.00 0.00 0.00 0.00 0.00 0.00
Avg. Performance Score($\downarrow$) 42.46 35.87 16.08 16.29 16.16 15.95 15.68 15.68 15.78
Average Recovery Ratio($\downarrow$) 70.08 59.20 26.53 26.89 26.67 26.33 25.87 25.88 26.05
Recovery Difficulty($\uparrow$) 2.22 5.48 10.92 11.29 11.35 11.19 11.17 11.20 11.20
Table 41: Evaluation Results of Mistral-7B under Different Closed-sourced Quantity
20M 50M 100M 160M 200M 300M 600M
Rsn. PIQA 77.79 73.74 51.36 52.86 53.34 50.98 51.62
Hellaswag 71.31 65.51 26.49 27.98 26.16 26.27 26.04
Winogrande 67.09 64.51 50.06 49.51 51.65 50.17 50.85
ARC_easy 74.52 68.29 27.84 30.95 30.82 27.36 28.30
ARC_challenge 42.32 37.97 20.85 21.67 20.71 21.28 20.17
Read. OpenBookQA 42.13 37.27 25.60 25.87 25.33 26.60 27.00
LAMBADA 55.99 47.63 1.16 4.74 1.66 0.43 0.53
BoolQ 78.35 75.00 40.17 47.05 42.01 42.05 39.03
SQuADv2_em 13.91 8.65 0.01 0.04 0.01 0.01 0.00
SQuADv2_f1 41.13 35.50 1.01 0.49 0.93 0.28 0.39
Knl. NaturalQuestions 8.46 7.82 0.02 0.05 0.00 0.01 0.02
TriviaQA 34.04 22.89 0.02 0.19 0.01 0.01 0.01
Code HumanEval 11.99 7.93 0.00 0.00 0.00 0.00 0.00
MBPP 16.93 11.87 0.00 0.00 0.00 0.00 0.00
Math GSM8K 6.32 2.48 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 44.17 41.28 24.22 24.44 24.34 23.78 23.33
BBH 35.44 33.43 0.00 0.40 0.00 0.00 0.00
Avg. Performance Score($\downarrow$) 42.46 37.75 15.81 16.84 16.29 15.84 15.72
Average Recovery Ratio($\downarrow$) 70.08 62.31 26.10 27.79 26.89 26.14 25.95
Recovery Difficulty($\uparrow$) 2.22 3.44 11.14 10.85 11.10 11.23 11.22
Table 42: Evaluation Results of Phi-2 under Different Closed-sourced Proportion
0.25% 0.5% 1% 3% 7% 15% 30% 50% 100%
Rsn. PIQA 70.40 70.71 74.64 54.43 54.17 54.75 54.37 52.39 52.07
Hellaswag 53.13 52.99 62.84 27.88 27.61 27.77 28.01 26.30 25.26
Winogrande 66.17 66.43 69.93 51.49 51.56 51.46 51.44 49.12 48.91
ARC_easy 64.62 65.33 72.55 33.39 34.57 32.00 32.18 29.97 27.03
ARC_challenge 43.26 43.86 40.67 20.82 19.45 20.00 20.56 19.88 18.66
Read. OpenBookQA 41.80 42.67 38.87 26.87 25.80 26.33 26.53 26.07 20.80
LAMBADA 32.51 32.25 40.24 10.58 3.25 3.87 6.06 0.66 0.00
BoolQ 65.77 65.27 76.84 48.13 47.29 45.62 46.15 40.50 39.60
SQuADv2_em 0.36 9.09 3.31 0.02 0.02 0.01 0.01 0.00 0.56
SQuADv2_f1 24.81 30.83 30.47 0.45 2.61 0.57 2.52 1.67 0.90
Knl. NaturalQuestions 5.70 5.06 1.14 0.03 0.00 0.01 0.07 0.03 0.02
TriviaQA 20.27 21.50 8.78 2.02 0.01 0.02 0.01 0.01 0.01
Code HumanEval 22.16 26.83 17.68 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 25.07 26.40 9.73 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 29.26 31.36 2.00 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 41.76 42.17 43.86 30.31 26.16 25.79 24.85 24.03 22.95
BBH 18.98 21.55 9.59 3.06 0.01 0.79 0.24 0.00 0.00
Avg. Performance Score($\downarrow$) 36.83 38.49 35.48 18.20 17.21 17.00 17.24 15.92 15.10
Average Recovery Ratio($\downarrow$) 62.40 65.22 60.12 30.95 29.15 28.81 29.21 26.97 25.59
Recovery Difficulty($\uparrow$) 6.70 6.65 2.00 9.14 10.07 10.13 10.14 9.82 11.32
Table 43: Evaluation Results of Phi-2 under Different Closed-sourced Quantity
20M 50M 100M 160M 200M 300M 600M
Rsn. PIQA 73.70 70.00 53.90 54.17 53.01 54.75 54.28
Hellaswag 59.75 55.64 28.26 27.61 26.90 27.77 28.61
Winogrande 66.61 67.17 51.96 51.56 52.28 51.46 50.88
ARC_easy 70.96 67.02 35.17 34.57 31.84 32.00 31.62
ARC_challenge 48.30 42.52 21.84 19.45 20.39 20.00 20.56
Read. OpenBookQA 45.33 41.27 26.13 25.80 25.60 26.33 26.53
LAMBADA 35.64 25.34 1.93 3.25 2.17 3.87 5.78
BoolQ 75.37 66.25 51.66 47.29 40.81 45.62 47.69
SQuADv2_em 10.62 0.10 0.14 0.02 0.02 0.01 0.00
SQuADv2_f1 38.28 22.83 1.33 2.61 1.36 0.57 1.13
Knl. NaturalQuestions 5.44 4.51 0.06 0.00 0.02 0.01 0.05
TriviaQA 12.34 12.77 0.05 0.01 0.01 0.02 0.01
Code HumanEval 20.94 10.98 0.00 0.00 0.00 0.00 0.00
MBPP 12.60 13.40 0.00 0.00 0.00 0.00 0.00
Math GSM8K 7.52 7.78 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 43.07 39.45 26.26 26.16 25.85 25.79 25.38
BBH 12.35 18.02 0.00 0.01 0.00 0.79 0.12
Avg. Performance Score($\downarrow$) 37.57 33.24 17.57 17.21 16.49 17.00 17.22
Average Recovery Ratio($\downarrow$) 63.67 56.32 29.77 29.15 27.93 28.81 29.17
Recovery Difficulty($\uparrow$) 2.07 7.96 9.25 9.96 10.08 10.13 10.22
Table 44: Evaluation Results of Phi-1.5 under Different Closed-sourced Proportion
0.25% 0.5% 1% 3% 7% 15% 30% 50% 100%
Rsn. PIQA 68.21 68.37 69.68 65.85 53.43 52.94 52.36 51.25 50.44
Hellaswag 49.05 49.18 49.30 30.72 26.27 26.74 27.02 26.10 25.05
Winogrande 63.83 64.91 61.20 58.04 51.09 51.38 50.25 50.22 49.12
ARC_easy 62.94 62.89 62.25 35.15 30.81 29.27 29.64 27.99 27.50
ARC_challenge 36.98 37.49 32.91 25.97 20.56 20.36 20.08 20.88 21.22
Read. OpenBookQA 39.07 40.20 35.00 33.87 26.60 27.67 27.73 26.47 26.87
LAMBADA 24.71 24.99 25.36 0.11 0.59 0.78 1.15 0.06 0.00
BoolQ 59.43 59.35 63.49 41.01 46.98 51.59 46.46 44.02 46.28
SQuADv2_em 15.65 16.00 3.13 0.50 0.00 0.01 0.03 0.00 0.00
SQuADv2_f1 32.62 32.62 14.88 0.56 0.78 1.24 2.29 1.58 1.60
Knl. NaturalQuestions 2.72 2.64 0.32 0.03 0.04 0.03 0.05 0.03 0.01
TriviaQA 8.17 7.96 5.69 0.01 0.01 0.01 0.01 0.01 0.01
Code HumanEval 14.43 13.41 2.03 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 17.20 18.67 6.47 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 4.88 4.90 0.25 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 30.12 29.88 28.98 27.78 24.07 24.22 24.66 24.28 22.95
BBH 4.34 3.19 0.98 0.50 0.00 0.00 0.00 0.00 0.00
Avg. Performance Score($\downarrow$) 31.43 31.57 27.17 19.41 16.54 16.84 16.57 16.05 15.94
Average Recovery Ratio($\downarrow$) 66.54 66.83 57.52 41.11 35.02 35.64 35.08 33.98 33.75
Recovery Difficulty($\uparrow$) 6.18 6.15 2.76 9.28 10.08 11.19 10.54 10.23 11.26
Table 45: Evaluation Results of Phi-1.5 under Different Closed-sourced Quantity
20M 50M 100M 160M 200M 300M 600M
Rsn. PIQA 69.80 65.85 53.43 52.52 52.94 53.06 53.81
Hellaswag 49.51 25.72 30.27 26.31 26.74 27.05 26.51
Winogrande 62.56 58.04 51.09 50.83 51.38 50.57 49.99
ARC_easy 62.41 30.15 30.81 29.14 29.27 29.62 29.67
ARC_challenge 32.51 25.97 20.56 19.97 20.36 20.48 20.79
Read. OpenBookQA 35.53 33.87 26.60 26.93 27.67 28.20 26.87
LAMBADA 28.14 0.11 0.59 0.45 0.78 1.30 0.61
BoolQ 64.77 41.01 46.98 47.33 51.59 46.09 45.59
SQuADv2_em 4.67 0.50 0.00 0.00 0.01 0.01 0.00
SQuADv2_f1 22.47 0.56 0.78 1.02 1.24 2.31 2.01
Knl. NaturalQuestions 1.64 0.03 0.04 0.05 0.03 0.06 0.03
TriviaQA 5.93 0.01 0.01 0.01 0.01 0.02 0.01
Code HumanEval 7.73 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 7.87 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 0.28 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 31.11 27.78 24.07 23.41 24.22 24.54 24.68
BBH 3.38 0.50 0.00 0.00 0.00 0.00 0.00
Avg. Performance Score($\downarrow$) 28.84 19.89 16.54 16.35 16.84 16.67 16.50
Average Recovery Ratio($\downarrow$) 61.06 41.11 35.02 34.61 35.64 35.28 34.94
Recovery Difficulty($\uparrow$) 2.81 9.28 10.26 11.65 11.19 10.87 10.49
C.9 Limitation on OPT-350M
[…]
Table 46: Evaluation Results of OPT-350M under Different Closed-sourced Layers. “*” indicates the fully closed-sourced model.
Pretrain 0-2 3-5 6-8 9-11 12-14 15-17 18-20 21-23 24-26 27-29 30-32 33-35 *
Rsn. PIQA 64.69 61.40 62.50 61.11 56.46 58.47 58.94 61.86 62.59 63.13 61.93 62.67 63.11 49.53
Hellaswag 36.68 34.03 34.27 33.69 31.79 32.24 32.78 33.27 33.68 33.25 33.94 33.63 33.07 25.77
Winogrande 52.09 51.62 52.96 51.57 50.83 52.83 51.06 51.52 51.85 52.06 52.04 51.99 50.94 49.85
ARC_easy 44.02 40.46 40.66 40.07 35.41 37.50 37.81 39.91 40.70 41.12 41.19 40.84 39.92 26.53
ARC_challenge 20.82 22.27 22.61 21.25 21.39 21.25 22.01 21.36 20.25 20.48 19.88 20.99 20.28 19.82
Read. OpenBookQA 28.00 27.60 27.47 27.40 27.40 27.27 26.20 26.47 27.67 26.80 28.67 27.67 27.13 27.47
LAMBADA 40.47 30.62 32.97 28.62 21.65 23.87 28.23 29.07 29.83 29.81 31.72 31.43 18.08 0.00
BoolQ 57.74 50.87 48.51 50.58 51.60 52.83 53.42 54.37 53.30 51.42 59.79 53.14 60.42 37.83
SQuADv2_em 11.34 6.87 7.88 4.74 4.19 0.27 0.87 2.22 3.79 3.05 4.11 4.69 2.35 0.00
SQuADv2_f1 19.35 16.27 17.00 12.00 11.72 9.04 6.92 10.11 10.90 10.08 8.88 11.47 7.30 0.01
Knl. NaturalQuestions 1.08 1.05 0.83 0.83 0.78 0.55 0.69 0.41 1.00 0.85 0.71 0.52 0.75 0.04
TriviaQA 4.48 2.24 2.66 2.01 2.16 1.41 1.06 2.39 2.38 2.29 1.57 1.90 1.76 0.02
Code HumanEval 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 1.59 0.15 0.25 0.00 0.08 0.00 0.05 0.00 0.03 0.18 0.00 0.00 0.00 0.00
Gen. MMLU 26.05 25.52 26.02 25.20 25.60 25.05 25.73 23.97 25.57 26.04 25.13 25.17 25.73 22.95
BBH 16.97 6.87 12.58 5.51 5.11 2.55 5.98 2.74 11.15 13.98 14.25 13.02 12.57 0.00
Avg. Performance Score($\downarrow$) 25.02 22.23 22.89 21.44 20.36 20.30 20.69 21.16 22.04 22.03 22.58 22.30 21.38 15.28
Average Recovery Ratio($\downarrow$) - 88.83 91.49 85.71 81.38 81.13 82.69 84.55 88.08 88.05 90.23 89.13 85.43 61.08
Recovery Difficulty($\uparrow$) - 5.92 9.32 9.04 8.60 8.83 8.73 7.17 5.82 5.18 4.65 4.92 4.15 10.89