### v2

# Position: On-Premises LLM Deployment Demands a Middle Path: Preserving Privacy Without Sacrificing Model Confidentiality

Abstract
Current LLM customization typically relies on two deployment strategies: closed-source APIs, which require users to upload private data to external servers, and open-weight models, which allow local fine-tuning but pose misuse risks. In this position paper, we argue that (1) deploying closed-source LLMs within user-controlled infrastructure (on-premises deployment) enhances data privacy and mitigates misuse risks, and (2) a well-designed on-premises deployment must ensure model confidentiality—by preventing model theft—and offer privacy-preserving customization. Prior research on small models has explored securing only the output layer within hardware-secured devices to balance confidentiality and customization efficiency. However, we show that this approach is insufficient for defending large-scale LLMs against distillation attacks. We therefore introduce a semi-open deployment framework that secures only a few, carefully chosen layers, achieving distillation resistance comparable to fully secured models while preserving fine-tuning flexibility. Through extensive experiments, we show that securing bottom layers significantly reduces functional extraction risks. Our findings demonstrate that privacy and confidentiality can coexist, paving the way for secure on-premises AI deployment that balances usability and protection.
Machine Learning, Secure on-premises LLM Deployment
1 Introduction
Vendors of Large Language Models (LLMs) have introduced advanced models with remarkable capabilities to address diverse user needs (Minaee et al., 2024; Zhao et al., 2023). To enable customization and drive industry progress, vendors typically adopt two approaches. Closed-source vendors, like OpenAI, provide fine-tuning APIs that allow users to upload data to customize proprietary models such as GPT-4. In contrast, companies like Meta offer open-weight models, such as Llama3 (Dubey et al., 2024), which users can adapt within their own infrastructure, ensuring greater flexibility and control.
However, both approaches present significant limitations for privacy-sensitive users, such as financial institutions, healthcare organizations, and government agencies, which prioritize data security. Fine-tuning APIs from closed-source vendors offer encryption for sensitive data and comply with strict privacy regulations (Dong et al., 2023; Pang et al., 2024). However, their security fundamentally depends on vendor trust, raising ethical concerns and exposing users to potential data breaches (Sun et al., 2023).
Alternatively, fine-tuning open-weight models within private infrastructure—commonly referred to as on-premises deployment, where all data and model adaptation occur locally—ensures data protection and customization (Nevo et al., 2024). However, full disclosure of model architectures and weights heightens the risk of misuse by malicious actors, enabling misinformation generation, automated cyberattacks, and security bypasses (Hendrycks et al., 2023). These risks discourage vendors from releasing state-of-the-art (SOTA) models as open-weight, as uncontrolled distribution could lead to widespread harm. Beyond security concerns, maintaining high-quality open-weight models imposes substantial computational and financial costs (Wolfe et al., 2024), further disincentivizing full disclosure.
In this position paper, we argue that deploying closed-source SOTA models on-premises with restricted access to authorized users offers a middle path to mitigate privacy leakage and model misuse risks. However, prior research has shown that hardware-based extraction attacks can recover model parameters and architectures directly from CPUs and memory (Hu et al., 2020) within local infrastructures, enabling unauthorized users to steal proprietary model information from the deployment server. This underscores the need for stronger security measures. Therefore, we further argue that a well-designed on-premises deployment strategy must fully prevent model theft while preserving data privacy, ensuring no reliance on untrusted servers.
Our work provides an affirmative step toward addressing the fundamental challenge of balancing security and customization in on-premises LLM deployment. Existing approaches have sought to protect proprietary models using Trusted Execution Environments (TEEs) (Narra et al., 2019), but fully enclosing models within TEEs incurs prohibitive computational costs, limiting their practicality (Nayan et al., 2024). Prior research has attempted to mitigate this trade-off by securing only critical layers, such as the output layer, while leaving the remaining layers exposed for fine-tuning (Zhang et al., 2024a; Mo et al., 2020). However, our findings align with prior work showing that securing only the output layer is insufficient against distillation attacks (Zanella-Beguelin et al., 2021). Extending these attacks to Llama2-70B, we confirm that this vulnerability persists at scale, enabling near-complete functionality distillation in six domains, as shown in Figure 1. These vulnerabilities raise skepticism about whether model confidentiality and customization can truly coexist in on-premises deployment, highlighting the need for a security paradigm beyond final-layer protection. Without a viable solution, vendors risk intellectual property theft, while users must compromise data privacy when relying on external servers.
Figure 1: Semi-open Deployment.
Our work suggests that this dilemma can be resolved. We propose a theoretically inspired semi-open deployment strategy that secures only a few, carefully chosen layers, balancing security and customization while mitigating distillation risks. This enables secure on-premises AI adoption in privacy-sensitive industries such as healthcare, finance, and government. More broadly, it highlights the need for hybrid security solutions that balance usability and protection, lowering barriers to responsible AI deployment.
By demonstrating that confidentiality and customization can coexist, our work shifts the narrative from skepticism to opportunity, paving the way for AI-driven innovation in regulated sectors. Future research should explore optimal layer selection, adaptive security mechanisms, and distillation-resistant architectures to further enhance model confidentiality. We hope this work inspires the AI community to develop scalable solutions that ensure AI remains both secure and widely accessible. We summarize our contribution as follows.
  • Semi-open on-premises LLM deployment. We propose a semi-open deployment framework that secures a single, carefully chosen layer, achieving distillation resistance comparable to full encryption while preserving customization flexibility comparable to full parameter fine-tuning.
  • Layer-wise security transitions. We provide the first theoretical result identifying a transition layer that resists distillation attacks, revealing fundamental differences in security across model layers.
  • Comprehensive empirical evaluations. We have conducted extensive empirical evaluations across five models (Llama2 (70B,7B), Mistral-B, Phi-2 (2.7B), and Phi-1.5 (1.3B) ) under three distillation strategies with a 51k attack set, comparing its performance with three baselines (DarkneTZ, SAP-DP, Fully-secured).
  • Discussion of alternative positions. We objectively examine alternative positions, and highlight their benefits and limitations, reinforcing the need for a new framework that balances security, privacy, and customization. Our discussion aims to stimulate further advancements and foster new solutions in this evolving field.
2 Related Works
2.1 Data Privacy Risks
Using LLM services for customization introduces significant data privacy risks, as user data is exposed during transmission, storage, and processing (Li et al., 2024b). To address this, prior research has explored Homomorphic Encryption (Lee et al., 2022), which allows computations on encrypted data, and Differential Privacy (Wei et al., 2021), which injects noise to prevent sensitive data memorization. However, these methods suffer from high computational overhead (Zhou et al., 2024) and limited protection coverage (Wei et al., 2020). A more effective approach is on-premises deployment, which ensures data remains under user control (Nevo et al., 2024). While this mitigates privacy concerns, it shifts the risk to LLM vendors, who lose control over model usage and face heightened risks of theft and misuse (Hendrycks et al., 2023).
2.2 Model Theft Risks
Deploying LLMs on user-controlled infrastructure increases exposure to model extraction threats (Atli et al., 2020). Attackers can recover model parameters through hardware-based techniques, including side-channel attacks targeting GPUs (Nayan et al., 2024) and bus monitoring attacks (Hu et al., 2020). In addition to direct parameter extraction, distillation attacks allow adversaries to replicate model functionality using only black-box access, a process known as functional distillation (Nevo et al., 2024; Xu et al., 2024; Ezzeddine et al., 2024). While distillation attacks have been extensively studied in smaller models, such as CNNs (Orekondy et al., 2018), BERT (Sanh et al., 2020; Zanella-Beguelin et al., 2021), and ReLU-based models (Canales-Martínez et al., 2024; Jagielski et al., 2020), their effectiveness against large-scale LLMs remains an open question. Our work extends these attacks to Llama2-70B and demonstrates that securing only the output layer remains insufficient to prevent near-complete functionality replication.
2.3 Traditional Approaches
Balancing data privacy and model security in on-premises deployment requires novel solutions. One direction is privacy-preserving federated deployment, where a model is split between user-controlled infrastructure and vendor-managed servers (Shen et al., 2023; Huang et al., 2024). Users fine-tune local layers while the vendor controls deeper layers, with privacy-preserving techniques like differential privacy (Zhou et al., 2024) protecting sensitive data. However, industries with strict data regulations, such as banking, healthcare, and government, often require full on-premises deployment (Schillaci, 2024; Guerra-Manzanares et al., 2023), ensuring models remain within their infrastructure while allowing customization.
To secure models in on-premises settings, research has explored hardware-based protections such as Arm TrustZone (Pinto & Santos, 2019) and secure execution environments (Zhang et al., 2024b; Li et al., 2024a). These approaches isolate computations to prevent unauthorized extraction but impose high resource demands and restrict customization (Mo et al., 2020). A more flexible alternative is layer-wise security, where only critical layers are protected while others remain exposed (Lin et al., 2024; Chen et al., 2024; Zhang et al., 2024a). Studies suggest different strategies, including securing shallow layers (Elgamal & Nahrstedt, 2020), intermediate layers (Shen et al., 2022), or the output layer (Huang et al., 2024). However, most research has focused on smaller models, leaving the effectiveness of different security placements in LLMs unclear.
Our work investigates the layer-wise security in LLMs, showing that securing bottom layers better resists distillation. This suggests that protecting a single, well-chosen layer can defend against attacks while maximizing fine-tuning flexibility, enabling more secure on-premises deployment.
3 Preliminaries
3.1 Security Threat: Model Distillation
Semi-open LLMs. Let $f:\mathbb{R}^{n\times d}\rightarrow\mathcal{Y}$ denote a victim model that processes the feature matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ to produce an output in set $\mathcal{Y}$. Modern LLMs typically adopt a multi-layer architecture to capture complex patterns in the input data. Specifically, $f(\mathbf{X};\bm{\theta})=\varphi_{L}\circ...\circ\varphi_{1}(\mathbf{X})$ is a composition of multiple decoder layers $\varphi_{1},...,\varphi_{L}$, each shares the same architecture but with distinct parameters. The parameters of all layers are denoted by the vector $\bm{\theta}$. We consider a semi-open setting, in line with (Zanella-Beguelin et al., 2021) and (Xu et al., 2021), where certain layers of the LLM are secured while others remain public.
Figure 2: Workflow of model distillation attack
Adversary’s Objective. The adversary aims to replicate the functionality of a semi-open victim LLM, partially secured in a protected environment, by building a replacement model. The agreement between the victim and replicated models is measured through accuracy and fidelity on a specified testing dataset.
Adversary’s Knowledge. It is assumed that the adversary knows the architectures of both secured and unsecured modules, as prior work (Gou et al., 2021; Boix-Adsera, 2024) has shown that using the same architecture as the secured module significantly improves the effectiveness of distillation attacks. However, the adversary knows only the parameters of the unsecured module, while those of the secured module remain unknown.
Adversary’s Capability. The adversary is capable of querying the semi-open model, obtaining both the semantic output produced by the complete model and the representation vector generated by the secured module. Utilizing this information, the adversary constructs a distillation attack dataset denoted as $\mathcal{D}$. Since the adversary knows the architecture of the secured module, the adversary next replaces the secured module with a randomly initialized module of the same architecture. Using the constructed set $\mathcal{D}$, the adversary employs three distinct supervised distillation strategies to replicate the functionality of the secured module: (1) FT-all: Fine-tunes both the replacement and unsecured modules using output of the entire model as training labels. (2) FT-closed: Fine-tunes only the replacement model using output of the entire model, keeping the unsecured module fixed. (3) SEM (Tamber et al., 2024): Fine-tunes the replacement model using outputs from the secured module without involving the unsecured component.
3.2 Problem Formulation
In this paper, we analyze the performance of a large language model under a defined distribution $\mathbb{P}_{\mathbf{X}\times Y}$, describing the relationship between input matrix $\mathbf{X}$ and label $Y$. We assume the victim LLM $f(\mathbf{X};\bm{\theta})$ performs well on this distribution, and the attack set $\mathcal{D}$ comprises samples drawn from $\mathbb{P}_{\mathbf{X}\times Y}$. To evaluate agreement between the distilled LLM and ground-truth labels, we use a scoring function $s:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$. Secured layers are indexed by $I\subseteq[L]=\{1,\dots,L\}$. Let $\bm{\theta}_{\text{dist}}(I,\mathcal{D})$ represent the parameter vector of the distilled replica of a victim model, where layers indexed by $I$ are secured, and adversaries utilize the attack set $\mathcal{D}$ to replicate its functionality. For each secured set $I$, we define the "Distillation Ratio" $R(I)$, which quantifies how well the distilled model $\bm{\theta}_{\text{dist}}(I,\mathcal{D})$ replicates the behavior of $f(\mathbf{X};\bm{\theta})$, expressed as
$R(I)=\frac{\mathbb{E}[s(f(\mathbf{X};\bm{\theta}_{\text{dist}}(I,\mathcal{D})),Y)]}{\mathbb{E}[s(f(\mathbf{X};\bm{\theta}),Y)]}.$ (1)
Here, $\mathbb{E}$ in the numerator reflects the expectation computed over random samples $(\mathbf{X},Y)$ drawn from $\mathbb{P}_{\mathbf{X}\times Y}$, the random attack set $\mathcal{D}$, and the random initialization of parameters within the secured layers during fine-tuning. Conversely, the term $\mathbb{E}$ in the denominator solely considers the expectation over random samples. With this definition, $R([L])$ represents the distillation ratio when the entire model is secured, reflecting the highest level of security. This leads to the question:
What is the smallest secured set $\text{What is the smallest secured set }I\text{ such that }$
$\quad\quad R(I)\text{ closely approximates }R([L])?$
This question aims to identify the minimal secured set $I$ such that securing the layers indexed by $I$ achieves a level of security comparable to securing the entire model.
4 Impact of Secured Layers Placement
Figure 3: Security and adaptability comparison in Llama2-70B. Lower scores indicate better security in Fig. (a) and weaker adaptability in Fig. (b). Details can be found in Appendix C.1
In this section, we investigate the impact of securing specific layers on security and customization against distillation attacks. We begin with an experiment with two semi-open deployments of Llama2-70B: one securing the bottom two decoder layers (Bottom2-Secured) and the other securing the top two decoder layers (Top2-Secured). As shown in Figure 3, both deployments achieve similar customization performance in six downstream tasks. However, securing the bottom layers provides significantly stronger security. Additionally, comparing Bottom2-Secured to fully-secured deployment reveals comparable security with improved customizability. This suggests that securing a certain number of bottom layers can effectively balance strong security against distillation attacks and high customization performance.
4.1 Security Transition in Deep Transformers
Model Overview. In this subsection, we consider a deep transformer $f$ with $L$ layers, expressed as $f(\mathbf{X};\bm{\theta})=\varphi_{L}\circ\dots\circ\varphi_{1}(\mathbf{X})$. The input feature matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ consists of $n$ rows, each representing a $d$-dimensional token vector. Each layer $\varphi_{i}$ is a transformer that incorporates a normalized residual self-attention mechanism, defined as:
$\varphi_{i}(\mathbf{X};K_{i},Q_{i})=\mathbf{X}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\mathbf{X}.$
Here, $Q_{i}\in\mathbb{R}^{d\times d_{Q}}$ and $K_{i}\in\mathbb{R}^{d\times d_{Q}}$ are projection matrices for the query and key components, respectively. The terms $\sqrt{d_{Q}}$ and $\|\mathbf{X}\|$ serve as normalization factors, ensuring stable computations within the attention mechanism. We consider the semi-open deployment of securing the $\alpha L$-th layer with $\alpha\in[0,1]$ and $\alpha L\in\mathbb{N}$ while keeping other layers unsecured. After the distillation attack, we assume the parameters of the distilled model in the unsecured layers are identical to the victim model, while those in the secured layer deviate. Let $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ denote the distilled weight matrix of the proprietary layer, i.e., ${\bm{\theta}}_{\text{dist}}(\{\alpha L\})=\{(K_{1},Q_{1}),...,(\hat{K}_{\alpha L},\hat{Q}_{\alpha L}),...,({K}_{L},{Q}_{L})\}$. Let $\hat{\varphi}_{\alpha L}$ denote the function of the distilled proprietary layer, i.e., the $\alpha L-$th layer, in the distilled model. In this subsection, we consider the normalized output of an infinitely deep model whose $\alpha L$-th layer is hidden and subjected to the attack. The output of the distilled model is
$\hat{f}_{\infty}(\mathbf{X})=\lim_{L\rightarrow\infty}\frac{f(\mathbf{X};{\bm{\theta}}_{\text{dist}}(\{\alpha L\}))}{\|f(\mathbf{X};{\bm{\theta}}_{\text{dist}}(\{\alpha L\}))\|_{F}},$
where $\|\cdot\|_{F}$ denotes the Frobenius norm. We consider an infinitely deep network as the ideal model, reflecting the sufficient depth of most large-scale models in practice. The following theorem establishes the existence of a critical value $\alpha^{*}$ such that if $\alpha<\alpha^{*}$, the output matrix of the distilled LLM has rank one. Conversely, if $\alpha>\alpha^{*}$, the output matrix has rank strictly greater than one.
Theorem 1.
Assume that $\mathbb{P}_{\mathbf{X}\times Y}$ is defined on a countable domain $\mathcal{X}\times\mathcal{Y}$ with $\mathbf{0}_{n\times d}\notin\mathcal{X}$. Assume that parameter matrices $\{K_{i},Q_{i}\}_{i\geq 1}$ in the victim model $f$ have uniform bounded norms, i.e., $\|K_{i}\|\leq D$ and $\|Q_{i}\|\leq D$ for some $D>0$. There exists an $\alpha^{*}\in(0,1)$ depending on $D$ such that the following two statements are true.
(1) If $\alpha<\alpha^{*}$ and $\{K_{i},Q_{i}\}_{i\geq 1}$ are parameter matrices of the victim model, with $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ as distilled parameters drawn from a continuous distribution on $\mathbb{R}^{n\times d}$, the matrix $\hat{f}_{\infty}(\mathbf{X})$ exists and almost surely has rank one for all $\mathbf{X}$.
(2) If $\alpha>\alpha^{*}$, there exists a victim model with parameter sequence $\{K_{i},Q_{i}\}_{i\geq 1}$ such that for any distilled parameters $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$, the matrix $\hat{f}_{\infty}(\mathbf{X})$ exists but has rank greater than one for some $\mathbf{X}$.
Remark 1: The proof is provided in Appendix A. This theorem demonstrates that if the distilled parameters of the bottom layers (i.e., $\alpha<\alpha^{*}$) are obtained through a randomized algorithm, such as stochastic gradient descent, with a continuous distribution supported on $\mathbb{R}^{n\times d}$, the distillation will certainly fail, as the feature matrix degenerate. In contrast, keeping the later layers secured (i.e., $\alpha>\alpha^{*}$) does not maintain this property, indicating that it is more effective to secure the bottom layers before the transition layer, rather than the later ones.
Remark 2: The existence of $\hat{f}_{\infty}(\mathbf{X})$ is a non-trivial result. While the mapping $\varphi_{i}$ admits a fixed point at $\mathbf{X}=\mathbf{0}_{n\times d}$, the convergence of the iterative process governed by $\varphi_{i}$ cannot be guaranteed using the contraction mapping theorem, as $\varphi_{i}$ does not satisfy the contraction property for any pair $(Q_{i},K_{i})$. This complexity becomes particularly evident in the special case where $n=1$ and $\mathbf{X}$ is a column vector. Here, the output of $\varphi_{i}$ satisfies the relation $\langle\mathbf{1}_{d},\varphi_{i}(\mathbf{X};K_{i},Q_{i})\rangle=2\langle\mathbf{1}_{d},\mathbf{X}\rangle$, implying that the iteration diverges unless $\mathbf{X}$ is orthogonal to $\mathbf{1}_{d}$. However, the divergence is not arbitrary; rather, the theorem reveals that it occurs in a fixed, well-defined direction. This insight ensures the existence of a normalized output, which remains stable and meaningful despite the lack of strict convergence.
Remark 3: The existence of $\alpha^{*}\in(0,1)$ is also a non-trivial statement, as $\alpha^{*}$ could potentially be zero, which would imply the absence of a critical layer such that securing layers prior to it guarantees the failure of the recovered model’s functionality. The primary challenge lies in demonstrating that perturbations to the earlier layers result in rank-one outputs, a property that does not universally hold for arbitrary perturbations. To address this, we establish an alternative result: given an input matrix $\mathbf{X}$, rank-one outputs can be guaranteed if the perturbation matrices $K_{i}$ and $Q_{i}$ are chosen to avoid specific zero-measure sets, denoted as $\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}(\mathbf{X})$, respectively. Assuming a countable domain $\mathcal{X}\times\mathcal{Y}$, which is typical for structured inputs such as sentences or images, it follows that the perturbation matrices to be avoided belong to the countable union of these sets, defined as $\mathcal{K}=\bigcup_{\mathbf{X}\in\mathcal{X}}\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}=\bigcup_{\mathbf{X}\in\mathcal{X}}\mathcal{Q}(\mathbf{X})$. Since this union remains a zero-measure set, avoiding these specific sets ensures that the conditions of the theorem are satisfied for any input matrix $\mathbf{X}$.
Theorem 1 shows that protecting bottom layers enhances resilience against model distillation attacks, indicating that securing layers starting from the first may be effective. In the next subsection, we propose SOLID, a preliminary LLM deployment framework solution that balance model protection and customization.
4.2 SOLID: Semi-Open Local Infrastructure Deployment Framework
We propose a method to approximately find the smallest bottom layer index set $I$ that satisfies $R(I)\leq(1+\varepsilon)R([L])$ for any small $\varepsilon>0$. A simple approach is to start with $I_{l}=\{1,\ldots,l\}$ for each $l$ beginning from 1, then evaluate the distillation ratio $R(I_{l})$ after the attack, and identify the smallest $l$ that meets the inequality. This extensive fine-tuning process is time-consuming, prompting the critical question: Can we create a fine-tuning-free metric that predicts LLM performance under model distillation attacks? Hence, our goal is to establish a metric directly correlated with the distillation ratio.
In the distillation ratio $R(I)$, each $I$ has the same denominator, so our focus is on a metric related to the numerator, specifically $\mathbb{E}[s(f(\mathbf{X};\bm{\theta}_{\text{FT}}(I,\mathcal{D})),Y)]$, which measures the average performance score of the distilled model. This average performance score generally inversely correlates with the average testing loss with the expression $L(\bm{\theta})\triangleq\mathbb{E}_{\mathbf{X}\times Y}[\ell(f(\mathbf{X};\bm{\theta}),Y)]$, where $\ell$ denotes the cross-entropy loss employed by LLM. Hence, we aim at finding the smallest $I$ such that
$L({\bm{\theta}}_{\text{dist}}(I,\mathcal{D}))\geq(1-\varepsilon)L({\bm{\theta}}_{\text{dist}}([L],\mathcal{D})).$
However, calculating both sides of this inequality requires knowing the distilled parameters from the fine-tuning process. To bypass this, we aim for an approximate solution. The distilled parameters are generated through gradient descent, starting from the initial parameters $\bm{\theta}_{0}(I)$, with the hidden layers being randomly initialized. Using the Taylor Expansion, we find
$L({\bm{\theta}}_{\text{dist}}(I,\mathcal{D}))$ $=L({\bm{\theta}}_{\text{0}}(I,\mathcal{D}))$
$\quad+\mathcal{O}(\mathbb{E}\|\bm{\theta}_{\text{dist}}(I,\mathcal{D})-\bm{\theta}_{0}(I)\|_{2}).$
Previous research (Choi et al., 2024; Bailly et al., 2022) indicates that the difference $\|\bm{\theta}_{\text{dist}}(I,\mathcal{D})-\bm{\theta}_{0}(I)\|_{2}$ is minimal in large networks compared to the dataset size $|\mathcal{D}|$. In models such as single-layer ReLU networks (Anthony et al., 1999; Zou et al., 2020), this difference scales as $\mathcal{O}\left(\frac{|\mathcal{D}|}{\sqrt{N}}\right)$ (Jacot et al., 2018; Wei et al., 2019), where $N$, the number of model parameters, far exceeds the dataset size in large language models (LLMs) (Dubey et al., 2024; Liu et al., 2024). The first term, independent of fine-tuning, dominates and effectively predicts the distillation ratio. We refer to this term as the Distillation Difficulty (DD$(I)$), defined as
$\text{DD}(I)=\mathbb{E}[L(\bm{\theta}_{0}(I))].$
This score, which can be estimated using a sample average, represents the distilled model performance of the model when specific layers $I$ are secured. A higher DD$(I)$ suggests better security performance, indicating a lower distillation ratio $R(I)$. Therefore, our SOLID operates in the following way. SOLID begins by sampling evaluation data targeting general capabilities from the underlying distribution, and then computes DD$(I_{l})$ for each set of secured layers $I_{l}=\{1,...,l\}$ for $l=1,...,L$. SOLID stops at the smallest $l^{*}$ that satisfies DD$(I_{l^{*}})\geq(1-\varepsilon)\text{DD}([L])$.
5 Experiments
In this section, we conduct experiments to validate the correctness of our positions. Specifically, we answer the following research questions:
  • RQ1. Can query-based distillation attacks distill the functionality of the entire model under the baseline deployment that secures the top layer?
  • RQ2. How do secured layer location and amount affect the security-customization trade-off?
  • RQ3. Does securing bottom layers (SOLID) offer a better balance between model theft risk and customization performance compared to baseline deployments?
5.1 Experimental Settings
We begin by introducing our experimental setups. Details can be found in Appendix B.
Models. We consider five open-source, decoder-only structured LLMs with various architectures. Specifically, we select Llama2-70B-chat, Llama2-7B-chat (Touvron et al., 2023), Mistral-7B-v0.1 (Jiang et al., 2023), Phi-2 (Abdin et al., 2024), and Phi-1.5 (Li et al., 2023). We designate these pre-trained models as the base models for adaptation and victims in model distillation attacks.
Attack Methods. We distill models produced by different protection approaches using three attack methods: FT-all, FT-closed and SEM. Following (He et al., 2021), a diverse attack set is required for full distillation. Therefore, we merge data evenly from two general datasets, MMLU (Hendrycks et al., 2021) and Alpaca 52k (Wang et al., 2022), resulting in a 51k combined set.
Baselines. We evaluate the effectiveness of SOLID by comparing it with three baselines: DarkneTZ (Mo et al., 2020), SAP-DP, and the fully-secured approach (Eiras et al., 2024). DarkneTZ protects only the final decoder layer while the SAP (Shen et al., 2023) framework exposes the first six decoder layers and secures the rest. SAP-DP extends SAP by adding Laplace noise to model outputs to enhance protection (Lee et al., 2018). The fully-secured approach represents the extreme, securing all layers for maximal security.
Implementation Details of SOLID. We apply the SOLID algorithm to identify the smallest secure set $I$ such that $R(I)\leq(1+\varepsilon)R([L])$. To calculate distillation difficulty (DD), we use cross-entropy loss and approximate the expectation over samples distributed on the general domain and randomly initialized secured parameters. This is done using a 1,500-sample evaluation set randomly sampled from the MMLU and Alpaca 52k, with secured parameters initialized via Xavier initialization and averaged over three random seeds (20, 42, 1234). In our experiments, we find that $\varepsilon=0.05$ yields optimal performance.
Table 1: Distillation ratios on 6 functionalities under FT-all (SOLID||||SAP-DP||||Fully-secured||||DarkneTZ). “H.E.” in Code domain presents the benchmark “HumanEval”. More results and details are available in Appendix C.2.
Benchmark Llama2-70B Llama2-7B Mistral-7B Phi-2
Rsn. PIQA 62.6||||59.8||||63.0||||99.3 64.7||||64.7||||64.6||||99.1 63.0||||61.2||||60.2||||92.2 68.3||||65.6||||65.7||||99.1
Winogrande 68.5||||67.7||||68.3||||98.3 76.8||||74.8||||76.6||||100. 67.2||||69.0||||68.3||||89.5 68.3||||64.9||||64.8||||99.1
ARC-easy 31.9||||32.8||||31.3||||98.5 36.3||||35.5||||34.9||||97.6 32.3||||34.7||||32.0||||86.6 43.2||||35.3||||33.9||||99.5
ARC-challenge 38.5||||38.1||||44.2||||99.2 47.8||||46.6||||50.9||||100. 39.7||||42.6||||44.5||||81.4 36.8||||36.6||||35.3||||99.5
Hellaswag 31.4||||31.4||||32.4||||98.1 33.9||||34.0||||35.0||||96.6 32.2||||32.0||||31.3||||84.6 37.4||||37.3||||34.3||||96.5
Read. LAMBADA 0.01||||0.00||||0.00||||88.6 0.02||||0.00||||0.01||||92.2 0.16||||0.00||||0.01||||67.9 1.34||||0.04||||0.00||||94.6
BoolQ 47.2||||47.1||||53.9||||100. 59.5||||56.0||||65.0||||99.6 48.3||||46.8||||56.7||||97.3 56.7||||50.3||||55.8||||100.
SQuADv2 1.50||||1.68||||0.34||||55.3 0.68||||0.88||||0.82||||59.5 1.69||||0.36||||0.93||||50.7 3.65||||0.39||||0.90||||62.9
OBQA 54.5||||54.5||||57.1||||99.6 57.4||||52.5||||59.2||||94.8 57.7||||56.8||||56.3||||84.0 0.00||||0.00||||0.02||||94.3
Knl. NaturalQuestions 0.00||||0.02||||0.00||||40.1 0.01||||0.01||||0.08||||53.6 0.00||||0.00||||0.02||||31.8 0.01||||0.00||||0.06||||87.4
TriviaQA 0.00||||0.02||||0.00||||72.3 0.00||||0.00||||0.03||||73.8 0.00||||0.00||||0.01||||38.7 0.01||||0.00||||0.01||||68.9
Code MBPP&H.E. 0.00||||0.00||||0.00||||58.6 0.00||||0.00||||0.00||||90.9 0.00||||0.00||||0.00||||40.2 0.00||||0.00||||0.00||||91.1
Math GSM8K 0.02||||0.00||||0.06||||79.6 0.00||||0.00||||0.00||||78.6 0.00||||0.00||||0.00||||31.1 0.00||||0.00||||0.00||||86.2
Gen. MMLU 36.8||||38.3||||36.5||||96.7 52.9||||50.0||||53.3||||110. 40.4||||36.9||||37.2||||81.7 42.6||||40.3||||40.5||||99.5
BBH 0.00||||0.00||||0.00||||93.3 0.00||||0.00||||0.00||||101. 0.00||||0.00||||0.00||||63.3 0.01||||0.00||||0.00||||94.8
Average Distillation Ratio($\downarrow$) 21.9||||21.8||||22.8||||77.9 25.3||||24.4||||25.9||||86.5 22.5||||22.4||||22.8||||73.7 23.9||||22.3||||22.4||||88.9
Secured Ratio($\downarrow$) 2.50||||92.5||||100.||||1.25 3.16||||81.3||||100.||||3.16 3.16||||81.3||||100.||||3.16 6.25||||81.3||||100.||||3.16
Evaluation Benchmarks We assess adaptability on six downstream tasks: Code (Zheng et al., 2024b), Math (Yue et al., 2023), Medical (Zhang et al., 2023), Finance (Wang et al., 2023b), Law (Guha et al., 2024), and Alignment (Meng et al., 2024). To fully evaluate recovered functionalities, we focus on six capabilities domains following Llama2 report (Touvron et al., 2023). Specifically, we assess the recovered model across sixteen benchmarks grouped into (1) Commonsense Reasoning (Rsn.); (2) Reading Comprehension (Read.); (3) World Knowledge (Knl.); (4) Code; (5) Math; and (6) General Ability (Gen.).
Metrics. We measure customization through model’s performance on benchmarks. For security, we calculate the “Average Distillation Ratio” (ADR) by averaging the distillation ratios across benchmarks. A lower ADR indicates higher security offered by the secure set.
5.2 Failure in Defense (RQ1)
We evaluate the security of DarkneTZ using three distillation strategies. Based on the results shown in Tables 1 and 2, we have following observations.
Observation 1: DarkneTZ, which secures only the last decoder layer, fails to protect the model against all three attacks (Table 1). DarkneTZ achieves ADRs generally exceeding 73% across four models, significantly higher than the average 24% ADR obtained by the fully-secured approach that offers maximum security. Specifically, on Llama2-70B and Phi-2, the distillation ratio for the Reasoning (Rsn.) task consistently exceeds 98%, indicating near-complete replication of model functionality. On Llama2-7B, the attack achieves a distillation ratio over 100% on the MMLU and BBH, suggesting the distilled model even outperforms the victim on these tasks. Similarly, Table 2 highlights consistent failure patterns against FT-closed and SEM attacks, with ADRs consistently above 75%. This confirms that securing only the top layers is insufficient to prevent the replication of model functionality.
Figure 4: (a) shows the trade-off between security and customization for Llama2-7B and Phi-2 with different placements of same-sized secured sets. (b) shows the trade-off as the secured set size increases from the first decoder layer. Smaller ADR indicates higher security and higher ACC reflects better customizability.
5.3 Security-Customization Trade-off (RQ2)
We conduct two experiments to analyze the impact of secured layer placement and quantity on the trade-off between security and customization. First, we secure one layer in Llama2-7B and two in Phi-2, varying their placement. Second, we incrementally secure both models by adding protected layers, starting from the smallest module (k_project) of the first decoder layer. These models are evaluated under the FT-all distillation attack and customized for the math domain. The results, as shown in Figure 4, lead to the following observations. More results and details can be found in Appendix B.7 and B.8.
Observation 2: Secured layer placement significantly impacts security, consistent with Theorem 1, but has limited effect on customization performance (Figure 4(a)). For Llama2-7B, security transitions at the fourteenth layer, with ADR consistently near 20% for earlier sets, indicating stronger security than protecting later layers. Meanwhile, customization accuracy remains stable across placements, highlighting the advantage of securing pre-transition layers. In contrast, Phi-2 transitions earlier at the first layer set, where only the first set balances security and customization, with later sets reducing security. These results suggest that securing layers before the transition layer optimizes the security-customization trade-off.
Observation 3: Increasing the number of secured layers enhances security but reduces customization (Figure 4(b)). The ADR of Llama2-7B decreases from 85% to 22% after securing an entire decoder layer, indicating improved security. However, customization accuracy drops from 29% to 21% as the number of secured layers increases from one to five, reflecting reduced customization flexibility. A similar trend is observed in Phi-2, suggesting that while increasing the number of secured layers enhances security (lower ADR), it negatively impacts customization flexibility (lower ACC) in both models.
5.4 Effectiveness of Securing Bottom Layers (RQ3)
We compare the security of SOLID with baseline deployments across three distillation strategies. The results lead to the following observations.
Table 2: Distillation ratios of Llama2-70B under FT-closed (FT-c.) and SEM attacks.
Strat. Method Rsn. Read. Knl. C.&M. Gen. ADR
FT-c. SOLID 47.1 21.6 0.00 0.03 18.7 22.6
SAP-DP 46.2 19.5 0.00 0.00 19.0 21.8
F-Secured 47.8 21.2 0.00 0.08 18.5 22.8
DarkneTZ 98.7 69.3 58.3 65.9 95.0 78.1
SEM SOLID 48.2 21.9 0.00 0.00 18.5 22.4
SAP-DP 47.1 21.1 0.00 0.00 18.3 22.3
F-Secured 47.8 21.2 0.00 0.08 18.5 22.8
DarkneTZ 98.8 71.2 54.2 66.3 94.1 77.4
Observation 4: SOLID achieves security against model distillation comparable to fully secured protection while securing significantly fewer parameters (Table 1). SOLID attains a similar security level (ADR) to SAP-DP and the fully secured approach across four architectures and multiple domains, securing at most 6.25% of parameters—compared to at least 80% for the others. For instance, on Llama2-70B, SOLID secures only 1.25% of parameters yet achieves an ADR of 21.9%, comparable to SAP-DP (21.8%) and the fully secured approach (22.8%). Additionally, SOLID maintains this security under FT-closed and SEM attacks, with ADR differences from SAP-DP and the fully secured approach remaining below 2.1% across six domains (Table 2). Details are in Appendix C.4 and C.5.
Figure 5: Customization performance comparison of secured models on six downstream tasks.
Observation 5: SOLID approaches the customization flexibility of full parameter fine-tuning, consistently outperforming baseline deployments (Figure 5). We observe that the performance of SOLID consistently matches full parameter fine-tuning across four architectures, with differences within 4%. Additionally, in the Law domain, SOLID improves performance by 10% over SAP-DP and the fully-secured approach on Llama2-70B, with a 35% improvement on the 7B models. Similar gains are evident on Phi-2, though the improvement in the Law domain reduces to 1%. These results demonstrate that securing a small portion of parameters minimally impacts customization, while providing strong protection against distillation attacks. Further results can be found in Appendix B.6 and C.3.
6 Alternative Views
Acknowledging diverse views and potential curiosities regarding LLM deployment design, we objectively examine several alternate viewpoints.
  • Alternative: Enhanced Regulation. Enhanced government data regulations can partially mitigate ethical privacy concerns and strengthen user confidence in data uploads and fine-tuning APIs, thereby promoting industry development. However, they do not eliminate the risk of data breaches and may struggle to address adaptive threats, rendering them insufficient for organizations with stringent privacy requirements, such as government agencies.
  • Alternative: Hybrid Privacy-Aware Deployment. Federated hybrid deployment enables users to deploy early model layers and process sensitive data locally while hosting later layers in the cloud for enhanced protection (Shen et al., 2023). It employs privacy-preserving techniques (e.g., differential privacy) to sanitize inputs before generating intermediate representations, ensuring secure data transmission. This approach preserves customization performance while improving both data privacy and model security. However, supervised fine-tuning still requires data uploads, and intermediate exchanges remain vulnerable to breaches (Zhou et al., 2022), thereby compromising the overall security assurance.
  • Alternative: Advanced Open-Weight Models. The success of DeepSeek-R1 (Guo et al., 2025) has showcased the power of open-weight models, raising expectations for more robust open-source alternatives. Developing open-weight models with strong security alignment could offer a viable solution to the ongoing dilemma between data privacy and model confidentiality. These models, capable of ensuring both usability and robust misuse detection, could address ethical concerns and serve as an alternative to the privacy-preserving deployment methods. However, their adoption may be constrained by limited profitability.
7 Conclusion
In this position paper, we advocate for a balanced approach to on-premises LLM deployment, addressing the dilemma between data privacy and model confidentiality. Building on our theorem, we propose SOLID, a simple yet effective semi-open solution that efficiently secures a few bottom layers of LLMs. We believe that our work shifts the view from skepticism to opportunity, paving a viable middle path for secure AI deployment that protects user privacy, safeguards vendor intellectual property, and preserves model customization flexibility. While our position may invite both support and critique, we hope it provides meaningful insights and fosters ongoing discourse. If this paper contributes to constructive engagement within the community, it will have fulfilled its purpose.
Impact Statement
The primary objective of this position paper is to advocate for a balanced approach to on-premises LLM deployment, ensuring both data privacy and model security. We introduce an efficient theoretically inspired semi-open deployment framework that mitigates functional extraction risks while enabling customization, providing a preliminary yet effective solution for privacy-sensitive industries. While our work primarily contributes to academic discourse, it also has broader implications for secure AI deployment in sectors where data confidentiality is critical. Ethically, we highlight the potential to simultaneously prevent data leakage, safeguard model security, and mitigate misuse risks. While we foresee no immediate societal consequences requiring specific emphasis, we acknowledge the importance of ongoing ethical considerations and the potential for future societal impacts as this interdisciplinary field evolves.
Appendix A Proof of Theorem 1
In this section, we prove Theorem 1. We first revisit the our model, present several important lemmas and finally present the proof.
A.1 Model Overview
The distilled model $f(\mathbf{X};\bm{\theta})$ is structured as a sequence of $L$ transformer layers,
$f(\mathbf{X})=\varphi_{L}\circ\varphi_{L-1}\circ...\circ\varphi_{\alpha L+1}\circ\hat{\varphi}_{\alpha L}\circ_{\alpha L-1}\circ...\circ\varphi_{1}(\mathbf{X}),$ (2)
where $\mathbf{X}\in\mathbb{R}^{n\times d}$ represents the input, interpreted as an assembly of $n$ tokens, each possessing $d$ hidden dimensions. Each transformer layer, indexed by $1\leq i\leq L$, is represented by $\varphi_{i}$, which maps $\mathbb{R}^{n\times d}$ to $\mathbb{R}^{n\times d}$ and can be defined as follows,
$\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)=\left[\mathbf{I}_{n}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\right]\mathbf{X},$ (3)
where $Q_{i}\in\mathbb{R}^{d\times d_{Q}}$, $K_{i}\in\mathbb{R}^{d\times d_{Q}}$ represent projection parameter matrices. Here, the $\alpha L$-th layer is the distilled layer and the others are the public layers. For simplicity, we use the function $\hat{\varphi}_{\alpha L}$ to denote mapping of the distilled layer, i.e., $\hat{\varphi}_{\alpha L}(\mathbf{X})=\varphi_{\alpha L}(\mathbf{X};\hat{K}_{\alpha L},\hat{Q}_{\alpha L})$.
A.2 Bounds on Different Orthogonal Components
Lemma 1.
For any $1\leq l\leq L$, $1\leq p\leq d$, any $\mathbf{X}\in\mathbb{R}^{n\times d}$, we have
$\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\varphi_{l}\left(\mathbf{X};K_{l},Q_{l}\right)[p]\right|\leq(1+\beta_{D})\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\mathbf{X}[p]\right|,$ (4)
where $\mathbb{I}_{n}$ is a column vector with dimensions $n\times 1$ and each element is 1, $\mathbf{X}[p]$ is the $p$-th column of the input $\mathbf{X}$, $\varphi_{l}\left(\mathbf{X};K_{l},Q_{l}\right)[p]$ is the $p$-th column of the $l$-th self-attention output, the coefficient $\beta_{D}$ satisfies $10<\beta_{D}<$ 1 and it is related to the upper bound of the L2-norm of matrices $K_{l},Q_{l}$.
Proof.
Let $\bm{u}=\left\{\bm{u}_{l,1}=\frac{\mathbb{I}_{n}}{\sqrt{n}},\bm{u}_{l,2},\ldots,\bm{u}_{l,n}\right\}$ denote the eigenvectors of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. Assume $\sigma_{l,1},\sigma_{l,2},\ldots,\sigma_{l,n}$ denote the eigenvalues of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$ and $-1<\sigma_{l,n}<\beta_{D}$ for any $l,n$. Thus we have
$\bm{v}^{\top}\varphi_{l}\left(\mathbf{X};K_{l},Q_{l}\right)[p]$ $=\bm{v}^{\top}\left[\mathbf{I}_{n}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\right]\mathbf{X}[p]$ (5a)
$=\bm{v}^{\top}\left[\mathbf{I}_{n}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\right]\sum_{k=1}^{n}\alpha_{pk}\bm{u}_{l,k}$ (5b)
$=\bm{v}^{\top}\sum_{k=1}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{u}_{l,k}$ (5c)
$\leq\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{v}^{\top}{u}_{l,k}\right|$ (5d)
$=\left\|\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{u}_{l,k}\right\|_{2}$ (5e)
$=\left[\sum_{k=2}^{n}\alpha^{2}_{pk}(1+\sigma_{l,k})^{2}\right]^{1/2}$ (5f)
$\leq(1+\beta_{D})\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\mathbf{X}[p]\right|,$ (5g)
where
$\beta_{D}=\max_{\left\|K_{l}\right\|_{2}\leq D,\left\|Q_{l}\right\|_{2}\leq D}\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}{\left\|\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\bm{v}\right\|_{2}}<1.$
The equation (5c) is due to $\bm{u}_{l,k}$ are the eigenvectors of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. The inequality (5e) is because when $\bm{v}=\frac{\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{u}_{l,k}}{\left\|\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k}){u}_{l,k}\right\|_{2}}$, we have the maximum value.
Lemma 2.
For any $K_{l},Q_{l}\in\mathbb{R}^{d\times s}$ and any $\mathbf{X}\in\mathbb{R}^{n\times d}$, the following equation always holds:
$\left|\mathbb{I}_{n}^{\top}\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]\right|=2\left|\mathbb{I}_{n}^{\top}\mathbf{X}[p]\right|,$ (6)
where $\mathbf{X}[p]$ is the $p$-th column of the input $\mathbf{X}$, $\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]$ is the $p$-th column of the $l$-th self-attention output.
Proof.
Assume that a set of orthogonal basis for $\mathbb{R}^{n}$ is $\left\{{\bm{u_{1}},\bm{u_{2}},\dots,\bm{u_{n}}}\right\}$, where $\bm{u_{1}}=\frac{\mathbb{I}_{n}}{\sqrt{n}}$. Then we can rewrite $\mathbf{X}[p]$ as $\mathbf{X}[p]=\sum_{j=1}^{n}\alpha_{pj}\bm{u_{j}}$, where $\alpha_{pj}$$(1\leq p\leq d)$ are the corresponding coefficients for the $p$-th column of $\mathbf{X}$ under the orthogonal basis. Next, we calculate $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|$ and $\left|\mathbb{I}_{n}^{\top}\mathbf{X}[p]\right|$, respectively. Note that $\mathbb{I}_{n}^{\top}\bm{u_{j}}=0$ for all $j\neq 1$. Therefore, we can obtain that,
$\mathbb{I}_{n}^{\top}\mathbf{X}[p]=\sqrt{n}\alpha_{p1}.$ (7)
Then we can get
$\left|\mathbb{I}_{n}^{\top}\mathbf{X}[p]\right|=|\sqrt{n}\alpha_{p1}|.$ (8)
Let $\sigma_{i1},\sigma_{i2},\ldots,\sigma_{in}$ denote the eigenvalues of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. Applying the Perron–Frobenius theorem for Markov matrices (Lemmens & Nussbaum, 2012), we deduce that for the matrix $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$, there exists only one eigenvalue equal to 1, while all other eigenvalues in absolute value are strictly less than 1. Without loss of generality, we assume $\sigma_{i1}=1$, implying $\left|\sigma_{ij}\right|<1$ for $j\neq 1$. Recalling the definition of $\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)$ and considering the linear operation, we can rewrite it as follows:
$\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]=\sum_{j=1}^{n}\alpha_{pj}\left(1+\sigma_{ij}\right)\bm{u_{j}}.$ (9)
Then we calculate the term $\left|\mathbb{I}_{n}^{\top}\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]\right|$ as follows,
$\left|\mathbb{I}_{n}^{\top}\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]\right|$ $=\left|\mathbb{I}_{n}^{\top}(\sum_{j=1}^{n}\alpha_{pj}\left(1+\sigma_{ij}\right)\bm{u_{j}}\right|$ (10a)
$=\left|\sqrt{n}\left(\alpha_{p1}(1+\sigma_{i1})\right)\right|$ (10b)
$=2|\sqrt{n}\alpha_{p1}|,$ (10c)
where (10a) is induced by substituting the equation (9) into $\left|\mathbb{I}_{n}^{\top}\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]\right|$, (10b) is due to $\mathbb{I}_{n}^{\top}\bm{u_{j}}=0$ for all $j\neq 1$, (10c) follows the fact that $\sigma_{i1}=1$
A.3 Proof of Theorem 1
We first prove the following result. For simplicity of notations, we use $f(\mathbf{X})\left[p\right]$ to denote the $p$-th ($1\leq p\leq d$) column of the the distilled model $f(\mathbf{X})$, where the parameters in the $\alpha L$-th layer is replaced with the matrices $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$. We use the function $\hat{\varphi}_{\alpha L}(\mathbf{X})=\varphi_{\alpha L}(\mathbf{X};\hat{K}_{\alpha L},\hat{Q}_{\alpha L})$ to denote the mapping of the $(\alpha L)$-th layer. Then we are going to show that there exists $\alpha^{\star}=\log_{2}{\frac{2}{1+\beta_{D}}}$ and $10<\beta_{D}<$ 1 makes the following equations hold.
(1) Assume $\alpha<\alpha^{\star}$. For any $\mathbf{X}$, $\left\|K_{i}\right\|_{2}\leq D,\left\|Q_{i}\right\|_{2}\leq D$, there exists a zero measure set $\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}(\mathbf{X})$ such that
$\lim_{L\to\infty}\left\|\frac{f(\mathbf{X})\left[p\right]}{\left\|f(\mathbf{X})\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}=0.$ (11)
(2) For any $\alpha>\alpha^{\star}$, there exists a sequence of matrix $\{K_{i},Q_{i}\}_{i\geq 1}$ such that for any distilled matrix $K_{\alpha L}$ and $Q_{\alpha L}$, we have $\left\|K_{i}\right\|_{2}\leq D,\left\|Q_{i}\right\|_{2}\leq D$, we have,
$\lim_{L\to\infty}\left\|\frac{f(\mathbf{X})\left[p\right]}{\left\|f(\mathbf{X})\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}=\sqrt{2}.$ (12)
Proof.
Based on Lemma (1), we obtain that
$\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}f\left(\mathbf{X}\right)[p]\right|\leq(1+\beta)^{L}\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\mathbf{X}[p]\right|.$ (13)
Based on Lemma (2), we know that
$\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=2^{(1-\alpha)L-1}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|.$ (14)
We firstly prove the equation (11). When
$\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=2^{(1-\alpha)L-1}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|\neq 0,$ (15)
then we have
$\left\|\frac{f(\mathbf{X})\left[p\right]}{\left\|f(\mathbf{X})\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}$ $=\left[2-\frac{2\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]}{\sqrt{n}\sqrt{\frac{(\mathbb{I}_{n}^{\top}f(\mathbf{X})[p])^{2}}{n}+(\bm{v}^{\top}f(\mathbf{X})[p])^{2})}}\right]^{1/2}$ (16a)
$=\sqrt{2}\left[1-\frac{1}{\sqrt{1+\frac{n(\bm{v}^{\top}f(\mathbf{X})[p])^{2}}{(\mathbb{I}_{n}^{\top}f(\mathbf{X})[p])^{2}}}}\right]^{1/2}$ (16b)
$\leq\sqrt{2}\left[1-\frac{1}{\sqrt{1+\frac{n(1+\beta)^{2L}|\bm{v}^{\top}\mathbf{X}[p]|^{2}}{2^{2[(1-\alpha)L-1]}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|^{2}}}}\right]^{1/2}$ (16c)
$\leq 2\sqrt{2n}\left(\frac{1+\beta}{2^{1-\alpha}}\right)^{L}\frac{\left|\bm{v}^{\top}\mathbf{X}[p]\right|}{\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|},$ (16d)
where the inequality (16c) is based on the inequality (13) and (14). The inequality (16d) is based on Lemma (3). Therefore, if $\alpha<\log_{2}{\frac{2}{1+\beta_{D}}}$ and $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|\neq 0$, then we have $\lim_{L\to\infty}\left(\frac{1+\beta_{D}}{2^{1-\alpha}}\right)^{L}=0$. Now we can consider when $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=0.$ In fact, it is easy to show that this can only happens when $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ belong to certain sets making $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=0,$ which corresponds to zero measure set $\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}(\mathbf{X})$ depending on the input $\mathbf{X}$. Since the input space is countable, therefore, the union $\cup_{\mathbf{X}\in\mathcal{X}}\mathcal{K}(\mathbf{X})$ and $\cup_{\mathbf{X}\in\mathcal{X}}\mathcal{Q}(\mathbf{X})$ are also zero-measure sets.
To prove equation (12), let $K^{\star}$, $Q^{\star}$ with $\left\|K^{\star}\right\|_{2}\leq D,\left\|Q^{\star}\right\|_{2}\leq D$ satisfy the following condition,
$\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}{\left\|\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\bm{v}\right\|_{2}}=\beta_{D}.$ (17)
Let $\bm{v}^{\star}$ be the solver of the above optimization problem (17) and consider the $K_{l}=K^{\star}$, $Q_{l}=Q^{\star}$ and $\mathbf{X}^{\star}=[\bm{v}^{\star},\bm{v}^{\star},\cdots,\bm{v}^{\star}]$. Clearly, $\bm{v}^{\star}\perp\mathbb{I}_{n}.$ Assume there exists $\bm{u}:\left\|\bm{u}^{\star}\right\|_{2}=1$ satisfying $\bm{u}^{\star}\perp\mathbb{I}_{n}$, $\bm{u}^{\star}\perp\bm{v}^{\star}$, therefore we can rewrite $f(\mathbf{X}^{\star})\left[p\right]$ as follows,
$f(\mathbf{X}^{\star})\left[p\right]=\frac{\mathbb{I}_{n}^{\top}}{\sqrt{n}}f(\mathbf{X}^{\star})\frac{\mathbb{I}_{n}}{\sqrt{n}}+\bm{v}^{\star\top}f(\mathbf{X}^{\star})\bm{v}^{\star}+\bm{u}^{\star\top}f(\mathbf{X}^{\star})\bm{u}^{\star}.$ (18)
For any $1\leq l\leq L$, based on Lemma (1), we know that
$\left|\bm{v}^{*\top}f\left(\mathbf{X}^{\star}\right)[p]\right|=(1+\beta_{D})^{L}\left|\bm{v}^{*\top}\mathbf{X}^{\star}[p]\right|.$ (19)
Since
$\left|\mathbb{I}_{n}^{\top}f\left(\mathbf{X}^{\star}\right)[p]\right|=2^{L}\left|\mathbb{I}_{n}^{\top}\mathbf{X}^{\star}[p]\right|=|\mathbb{I}_{n}^{\top}\bm{v}^{\star}|=0$ (20)
and
$\left|\bm{v}^{*\top}f\left(\mathbf{X}^{\star}\right)[p]\right|=(1+\beta_{D})^{L}\left|\bm{v}^{*\top}\mathbf{X}^{\star}[p]\right|\neq 0,$ (21)
then we have
$\left\|\frac{f(\mathbf{X}^{\star})\left[p\right]}{\left\|f(\mathbf{X}^{\star})\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}$ $=\left[2-\frac{2\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p]}{\sqrt{n}\left\|f(\mathbf{X}^{\star})\left[p\right]\right\|_{2}}\right]^{1/2}$ (22a)
$=\left[2-\frac{2\mathbb{I}_{n}^{\top}}{\sqrt{n}}\frac{f(\mathbf{X}^{\star})[p]}{\sqrt{\frac{1}{n}(\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p])^{2}+(\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p])^{2}+(\bm{u}^{\star\top}f(\mathbf{X}^{\star})[p])^{2}}}\right]^{1/2}$ (22b)
$\geq\left[2-\frac{2\mathbb{I}_{n}^{\top}}{\sqrt{n}}\frac{f(\mathbf{X}^{\star})[p]}{\sqrt{\frac{1}{n}(\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p])^{2}+(\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p])^{2}}}\right]^{1/2}$ (22c)
$=\left[2-2\frac{\frac{\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p]}{\sqrt{n}\left|\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p]\right|}}{\sqrt{1+\frac{\left|\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p]\right|^{2}}{n\left|\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p])\right|^{2}}}}\right]^{1/2}$ (22d)
$=\left[2-2\frac{\frac{2^{(1-\alpha)L-1}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X}^{\star})[p]\right|}{\sqrt{n}(1+\beta_{D})^{L}\left|\bm{v}^{\star\top}\mathbf{X}^{\star}[p]\right|}}{\sqrt{1+\frac{2^{2[(1-\alpha)L-1]}}{n(1+\beta_{D})^{2L}}\frac{\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X}^{\star})[p]\right|^{2}}{\left|\bm{v}^{\star\top}\mathbf{X}^{\star}[p]\right|^{2}}}}\right]^{1/2},$ (22e)
where equation (22b) is based on (18), equation (22e) is based on (21) and (14). When $\alpha>\log_{2}{\frac{2}{1+\beta_{D}}}$, we have $\lim_{L\to\infty}\left(\frac{2^{1-\alpha}}{1+\beta_{D}}\right)^{L}=0$. Thus we have $\lim_{L\to\infty}\left\|\frac{f(\mathbf{X}^{\star})\left[p\right]}{\left\|f(\mathbf{X}^{\star}\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}=\sqrt{2}$. This indicates that the $p$-th column of the output matrix $f(\mathbf{X}^{\star})$ is not parallel to $\mathbf{I}_{n}$ for any $p$. This further indicates that the output matrix does not have the identical vector in each row. ∎
A.4 Technical Lemma
Lemma 3.
For any $x\in(0,1)$, it always holds $\left[1-\frac{1}{\sqrt{1+x^{2}}}\right]^{1/2}\leq x$.
Proof.
To establish the inequality $\left[1-\frac{1}{\sqrt{1+x^{2}}}\right]^{1/2}\leq x$, we begin by proving,
$1-\frac{1}{\sqrt{1+x^{2}}}\leq x^{2}.$ (23)
To demonstrate (23), we equivalently show
$1-x^{2}\leq\frac{1}{\sqrt{1+x^{2}}}.$ (24)
Subsequently, it suffices to verify
$(1-x^{2})(\sqrt{1+x^{2}})\leq 1.$ (25)
This is equivalent to proving
$(1-x^{2})^{2}(1+x^{2})\leq 1.$ (26)
Thus, our focus shifts to demonstrating
$(1-x^{2})(1-x^{4})\leq 1.$ (27)
Clearly, (27) holds true for any $x\in(0,1)$. ∎
Appendix B Experiment Details
To more intuitively compare the security differences between the SOLID method and a fully-secured approach, we define $\Delta\textbf{ADR}(I)=\text{ADR}(I)-\text{ADR}([L])$ to assess the resilience of the secured set $I$ relative to the fully-secured approach. A smaller value of $\Delta\textbf{ADR}$ indicates resilience similar to that of the fully-secured model.
B.1 Model Details.
The foundation models we use in our experiments are selected from open-source repositories, and Table 3 shows the basic information of the models and their sources. Specifically, we employ Llama2-70B-chat [^1] https://huggingface.co/meta-llama/Llama-2-70b-chat-hf, Llama2-7B-chat [^2] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, and Mistral-7B-v0.1333https://huggingface.co/mistralai/Mistral-7B-v0.1. For smaller models, we select Phi-2444https://huggingface.co/microsoft/phi-2 and Phi-1.5555https://huggingface.co/microsoft/phi-1_5. We also consider OPT model [^6] https://huggingface.co/facebook/opt-350m, which has only 350 million parameters and 24 decoder layers.
Table 3: Model Info
Model Size Decoder Layers
Llama2-70B-chat (Touvron et al., 2023) 70B 80
Llama2-7B-chat (Touvron et al., 2023) 7B 32
Mistral-7B-v0.1 (Jiang et al., 2023) 7B 32
Phi-2 (Abdin et al., 2024) 2.7B 32
Phi-1.5 (Li et al., 2023) 1.3B 24
OPT (Zhang et al., 2022) 350M 24
B.2 Distillation Attacks.
Attack implementation details. In performing FT-all and FT-secure model distillation attacks, we adhere to the training hyper-parameters outlined in the Llama2 report (Touvron et al., 2023), employing the AdamW optimizer with a cosine learning rate scheduler. The initial learning rate is set to $2\times 10^{-5}$, with a weight decay of 0.1, a batch size of 128, and bfloat16 precision for input sequences of 512 tokens. The LLaMA2-70B model is trained for 3 epochs with a random seed of 42, while other models are trained for 5 epochs across three seeds: 42, 1234, and 20. Despite limiting training to 3 epochs for the 70B model, the training loss stabilized effectively. Our implementation builds upon the llama-recipes repository provided by META.
For SEM attacks, distinct configurations were employed for SOLID and SAP-DP. In the case of SOLID, hidden representations from the secure-source components were collected and paired with the input data to train a substitute model. In contrast, for SAP-DP, representations from the sixth decoder layer and the model’s final logits were utilized to construct the training dataset. In accordance with (Tamber et al., 2024), we applied a learning rate of 1.5e-4, a weight decay of 0.01, and a linear learning rate scheduler with 500 warmup steps. Both training and validation batch sizes were set to 32, with MSE as the loss function. SOLID was trained for 30 epochs due to its smaller model size, whereas SAP-DP was trained for 5 epochs.
All distillation experiments were conducted on Nvidia 4090 24G, 6000 Ada 48G, and A100 80G GPUs, utilizing PyTorch 2.2.0 and CUDA 11.8 on Ubuntu 20.04.6 LTS.
Base 51k Distillation Dataset. We ensure dataset coverage and reliability by using a 1:1 ratio of the MMLU auxiliary training set  [^7] https://github.com/hendrycks/test and Alpaca dataset  [^8] https://github.com/tatsu-lab/stanford_alpaca/blob/main/alpaca_data.json, extracting 25.5k samples from each. From the MMLU auxiliary training data (Hendrycks et al., 2021), we sample 50%, and from Alpaca (Taori et al., 2023), we use a step size of 2 to enhance diversity. The datasets are then formatted for model training, applying Alpaca and MMLU prompts from Table 4.
Table 4: Prompts for Alpaca and MMLU auxiliary training data
Dataset Prompt Type Description
Alpaca with input Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request.
w/o input Below is an instruction that describes a task. Write a response that appropriately completes the request.
MMLU Question Answering Below is a question with no choices. Write the correct answer that appropriately solves the question.
Multiple Choice The following is a multiple choice question, paired with choices. Answer the question in the format: “Choice:content”.
Validation Datasets. Table 5 outlines the composition of the validation dataset. We extracted 50% from each of the 57 MMLU validation sub-datasets, totaling 1.5K instances, paired with Alpaca data selected using a step size of 751.
Table 5: Composition of validation datasets of different sizes
Raw Data Set Validation Set Evaluation Set
Alpaca 765 765
MMLU auxiliary training set 751 751
Total Length 1516 1516
B.3 Baselines.
In this section, we provide further details on the baselines used in our comparisons: SAP-DP and fully-secured. These schemes represent different strategies, each with distinct trade-offs in terms of customizability and security against model distillation attacks.
SAP. The Split-and-Privatize (SAP) framework (Shen et al., 2023) offers an approach to balance between protecting model privacy and data privacy while maintaining competitive performance. Specifically, the SAP framework keeps the bottom six encoder layers open, allowing user access and fine-tuning while securing the deeper layers on the vendor.
SAP-DP. To further strengthen protection while maintaining competitive performance, we extend SAP by incorporating differential privacy techniques by adding Laplace noise to perturb the logits during the fine-tuning process (Lee et al., 2018). The Laplace Distribution with mean $\mu$ and scale $b$ is the distribution with probability density function:
$\text{Laplace}(x|\mu,b)=\frac{1}{2b}\exp\left(-\frac{|x-\mu|}{b}\right)$
Specifically, in SAP-DP, the noise $n$ is sampled: $n\sim\text{Laplace}(0,0.5)$ and added to the output logits of the model to balance privacy protection and model performance.
Fully-secured. Following (Eiras et al., 2024), we use the fully-secured approach as a baseline. This assumes the adversary has no access to internal model parameters, treating the model as a black-box, where only output data can be collected. We slightly broaden this setup by assuming the adversary knows the model’s architecture but no other details. Thus, distilling the fully-secured model involves using the collected data to retrain a model with the same architecture to restore its general functionality.
DarkneTZ. Based on the work of (Mo et al., 2020), we use DarkneTZ as a baseline to test whether protecting only the output layers is sufficient to defend against distillation attacks. In this setup, we assume the adversary has no access to the model parameters of the output layers, specifically the last decoder layer. Similar to the SAP framework, this approach allows the adversary to access and fine-tune all layers except the final decoder layer.
B.4 Implementation Details of SOLID.
Evaluation Datasets. We created a 1.5K Evaluation Set to assess model security under various secure-sourcing strategies. This set includes 50% of entries from each of the 57 MMLU validation sub-datasets (Hendrycks et al., 2021), distinct from Validation Set outlined in Table 5. Additionally, we selected an equal number of Alpaca dataset (Taori et al., 2023), using a step size of 751, ensuring no overlap with the Validation Set.
Hyper-parameter Sensitivity. As shown in Figure 6, we evaluate SOLID’s sensitivity to tolerance magnitude $\varepsilon$, adjusting it from 0.05 to 1 in 0.05 increments while calculating the $\Delta\text{ADR}$ for six distilled models. The results indicate that SOLID is minimally sensitive to changes in $\varepsilon$, with $\Delta\text{ADR}$ values stabilizing as $\varepsilon$ increases. This stability arises from the need for a smaller secured layer at higher $\varepsilon$, allowing the condition $R(I)\leq(1+\varepsilon)R([L])$ to be met with fewer layers. Additionally, the increase in $\Delta\text{ADR}$ is smaller for larger models, suggesting that privatizing more parameters beyond a certain point offers diminishing returns in security.
Figure 6: Sensitivity on $\varepsilon$.
B.5 Evaluation Benchmarks
Most of our evaluations are conducted using the lm-evaluation suite (Gao et al., 2023), the bigcode-evaluation-harness platform (Ben Allal et al., 2022), and MT-Bench (Zheng et al., 2023). For specific domains, such as finance and law, we utilize the official benchmark testing codes provided by their respective communities, as detailed below.
Evaluation on Customizabilities. We assess the customizability of models across six domains, as detailed in Table 6. Each domain includes specific benchmarks and metrics designed to evaluate different aspects of the model’s performance in relation to customizability. In particular, for evaluating medical capabilities, we select two subcategories from the MMLU benchmark that are related to the medical domain: mmlu_anatomy and mmlu_professional_medicine. For assessing legal reasoning, we select 10 multiple-choice and judgment-based subcategories from Legalbench. The performance of the model in these legal tasks is measured using perplexity, following the prompt structure provided by Legalbench. Specifically, the selected subcategories include: cuad_audit_rights, canada_tax_court_outcomes, definition_classification, cuad_affiliate_license-licensee, learned_hands_business, contract_nli_survival_of_obligations, contract_nli_explicit_identification, contract_nli_confidentiality_of_agreement, hearsay, and contract_qa.
Table 6: Details of the Six Customizability Benchmarks
Domain Benchmark Metric n-shot Reference
Code HumanEval Pass@1 0 (Chen et al., 2021)
MBPP Pass@1 1 (Austin et al., 2021)
Math GSM8K Exact Match 8 (Cobbe et al., 2021)
Medical MMLU_Medical Accuracy 5 (Hendrycks et al., 2021)
Finance FPB F1 0 (Wang et al., 2023a)
Law LegalBench Accuracy 0 (Guha et al., 2023)
Alignment MT-Bench Score (GPT-4) (Zheng et al., 2023)
Evaluation on Security. We follow the Llama-2 report (Touvron et al., 2023) to evaluate the distilled model, including 16 benchmarks, which are categorized into 6 groups. Table 7 summarizes the functionality benchmarks used in our experiments, along with their test methods and performance metrics. Our model ranks choices in multiple-choice tasks and generates answers for open-ended generation tasks.
Table 7: Details of the Sixteen Functionality Benchmarks
Domain Benchmark Metric n-shot Reference
Commonsense Reasoning PIQA Accuracy 0 (Bisk et al., 2020)
Hellaswag Accuracy 0 (Zellers et al., 2019)
Winogrande Accuracy 0 (Sakaguchi et al., 2019)
ARC_easy Accuracy 0 (Clark et al., 2018)
ARC_challenge Accuracy 0 (Clark et al., 2018)
Reading Comprehension OpenBookQ Accuracy 0 (Mihaylov et al., 2018)
LAMBADA Accuracy 0 (Paperno et al., 2016)
BoolQ Accuracy 0 (Clark et al., 2019)
SQuADv2 HasAns_EM 2 (Rajpurkar et al., 2018)
SQuADv2 HasAns_F1 2 (Rajpurkar et al., 2018)
World Knowledge NaturalQuestions Exact Match 5 (Kwiatkowski et al., 2019)
TriviaQA Exact Match 5 (Joshi et al., 2017)
Code HumanEval Pass@1 0 (Chen et al., 2021)
MBPP Pass@1 1 (Austin et al., 2021)
Math GSM8K Exact Match 8 (Cobbe et al., 2021)
General Ability MMLU Accuracy 5 (Hendrycks et al., 2021)
BBH Accuracy 3 (Suzgun et al., 2022)
B.6 Model Customization
Datasets. To fine-tune the models for domain-specific tasks, we utilized several datasets tailored to different sectors, including Code (Zheng et al., 2024b), Math (Yue et al., 2023), Medical (Zhang et al., 2023), Finance (Wang et al., 2023b), Law (Guha et al., 2024), and Alignment (Meng et al., 2024). Table 8 lists the customization training datasets used in the experiments. For the code domain, we combine the datasets from CodeFeedback and CodeAlpaca. For law and finance, we merge all training datasets from Legalbench and FinGPT respectively. These datasets are then prepared for model training using the Alpaca prompts outlined in Table 4. Additionally, we randomly select 3,000 samples to serve as the validation dataset.
Table 8: Customization Training Datasets Composition
Domain Dataset Name Size Reference
Code CodeFeedback 156k  (Zheng et al., 2024a)
CodeAlpaca 20k (Chaudhary, 2023)
Math MathInstruction 262K (Yue et al., 2023)
Medical MedMCQA 183k (Zhang et al., 2023)
Law Legalbench 90k (Guha et al., 2023)
Finance FinGPT 204k (Wang et al., 2023a)
Alignment Ultrafeedback 62k (Cui et al., 2024)
Customization Training Hyperparameters. In model customization, we use different hyperparameters depending on the model size. For LLaMA2-70B, we apply QLoRA with the settings outlined in Table 9, while for 7B models, we use LoRA. For smaller models like Phi2 and Phi-1.5, we fine-tune all model parameters. For LLaMA2-70B, we fine-tune it as a quantized 4-bit model over 1 epoch, starting with a learning rate of $1.5\times 10^{-6}$. For the 7B models, we train for 3 epochs, with a seed value of 42. The training setup includes a weight decay of 0.1, a batch size of 128, a warmup ratio of 0.03, and input sequences of 512 tokens, following standard experimental practices (Hu et al., 2021). For Phi2 and Phi-1.5, we use the training hyperparameters from the LLaMA2 report. We employ the AdamW optimizer with a cosine learning rate scheduler, starting with a learning rate of $2\times 10^{-5}$, a weight decay of 0.1, a batch size of 128, and use bfloat16 precision for 512-token input sequences. Specifically, for alignment, we follow SimPO (Meng et al., 2024) and set the preference parameters $\beta=2$ and $\gamma=1$. The learning rate is $1\times 10^{-6}$ for LLaMA2-70B and $5\times 10^{-7}$ for the 7B and smaller models. All experiments are conducted using the LLaMA-Factory on Nvidia 4090 24G, 6000 Ada 48G, and A100 80G GPUs, with PyTorch 2.2.0 and CUDA 11.8 on Ubuntu 20.04.6 LTS.
Table 9: The Hyperparameters for Customization Training.
Model Method Rank $r$ Lora $\alpha$ Dropout Learning Rate Epochs Warmup R.
Llama2-70B QLoRA 96 16 0.05 1.50E-04 1 0.03
Llama2-7B LoRA 32 64 0.05 2.00E-05 3 0.03
Mistral-7B LoRA 32 64 0.05 1.00E-06 3 0.03
B.7 Security and Customization Transitions
For the LLaMA2-7B model, the smallest secure-source layer set identified by SOLID consists of a single decoder layer, whereas for Phi-2, it includes two decoder layers. Consequently, for LLaMA2-7B, we opted to secure-source each even-indexed layer, while for Phi-2, we chose to secure-source non-overlapping pairs of layers (e.g., layers 0-1, 2-3). For each selected layer set, we first secure-source them, then subjected the semi-open model to FT-all attacks, and subsequently calculated the $\Delta\text{ADR}$ of the layer set to assess its security.
When verifying the customization transition, due to computational constraints, we validated only every other layer set for both models (e.g., secure-source layers 0, 0-4, 0-8 …). Specifically, we applied LoRA-based customization on LLaMA2-7B in the math domain, while for Phi-2, we utilized the full finetuning approach. The experimental hyperparameters remain consistent with those outlined in the Appendix B.6.
We further computed the $\Delta\text{ADR}$ for each secure-source set within Mistral-7B-v0.1 and Phi-1.5. In these models, the smallest secure-source set identified by SOLID consists of one decoder layer and two decoder layers, respectively. Following the same experimental configuration as LLaMA2-7B and Phi-2, we secured each even-indexed layer for Mistral-7B, and non-overlapping pairs of layers for Phi-1.5. The complete results demonstrating the transition layers within the Mistral-7B and Phi-1.5 model that secure two non-overlapping consecutive layers are depicted in Figure 7. Once again, we observed a distinct presence of transition layers. Specifically, in Mistral-7B, the transition layer appears at the 24th layer, while in Phi-1.5, it is located within the first layer set. Further results for can be found in Appendix C.7.
Figure 7: Security changes in Miatral-7B and Phi-1.5.
B.8 Security Across Secure Sizes
To examine the influence of Secure layer size on model security, we conduct experiments on Secure-sourcing different amounts and proportions of parameters in the model’s decoder layer. We give instructions on the detailed setting of secured models in Table 10. The module names are all derived from the overall implementation functions of each model in the Transformers open-source repositories in Table 3. We utilize abbreviated module names to denote specific settings.
Table 10: Secured Sizes Setting. “*” indicates an entire decoder layer.
Llama-7B Mistral-7B Phi2-2.7B Phi1.5-1.3B
0.25% $\mathrm{W}_{k}$ $\mathrm{W}_{q},\mathrm{W}_{k}$ $\mathrm{W}_{k}$ $\mathrm{W}_{k}$
0.50% $\mathrm{W}_{q},\mathrm{W}_{k}$ $\mathrm{W}_{o},MLP_{\text{up}}$ $\mathrm{W}_{q},\mathrm{W}_{k}$ $\mathrm{W}_{q},\mathrm{W}_{k}$
1% $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{o}$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{o}$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{d}$ $\mathrm{W}_{q},\mathrm{W}_{k},,\mathrm{W}_{v}$
3% 0 0 0 0
7% 0-1 0-1 0-1 0-1
15% 0-4 0-4 0-3 0-3
30% 0-9 0-9 0-9 0-6
50% 0-15 0-15 0-15,$\mathrm{W}_{em}$ 0-11,$\mathrm{W}_{em}$
Proportion 100% Fully-secured Fully-secured Fully-secured Fully-secured
20M $\mathrm{W}_{k}$ $\mathrm{W}_{q},\mathrm{W}_{k}$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v}$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{d}$
50M $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v}$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{o}$ $MLP$ 0
100M $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},MLP$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{o},MLP$ 0, $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v}$ 0-1
160M $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{o},MLP$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{o},MLP$ 0-1 0-2
200M 0 0 0-1, $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{d},MLP_{\text{f1}}$ 0-3
300M 0, $\mathrm{W}_{q},\mathrm{W}_{v},\mathrm{W}_{o},MLP_{\text{up}}$ 0, $\mathrm{W}_{q},\mathrm{W}_{v},\mathrm{W}_{o},MLP_{\text{up}}$ 0-3 0-5
Quantity 600M 0-2 0-2 0-7 0-11
We further computed $\Delta\text{ADR}$ by close-sourcing varying quantities and proportions of parameters under FT-all attacks on three additional models. As shown in Figure 8 and Figure LABEL:fig:appendix:appendix-specific-domain(a), we observed the same pattern as with Llama2-7B, where security emerges once a sufficient number of parameters are secured. For example, on Mistral-7B, security occurs after secure-sourcing 100 million parameters, which is less than a single decoder layer. Secure-sourcing fewer parameters leads to a notable drop in security, with $\Delta\text{ADR}$ rising to around 40%. Beyond this threshold, security stabilizes near 0% $\Delta\text{ADR}$. This pattern holds across all models, highlighting a critical threshold for effective secure-source. Furthermore, different architectures require varying secure-sourcing quantities to achieve security, even with similar model sizes. For instance, Mistral-7B reaches security by secure-sourcing 100 million parameters, Llama2-7B requires 200 million, and Phi-1.5 needs a higher rate of 7%, compared to 3% for Llama2-7B.
Figure 8: $\Delta$ ADR for different secure parameter quantities and proportions.
We explore how secured parameter ratio impacts the model security in Llama2-7B, as shown in Figure LABEL:fig:appendix:appendix-specific-domain(b). For instance, technical skills such as Math show earlier transitions, with security emerging at 1% parameters secured, whereas domains such as Commonsense Reasoning require hiding 3%. In summary, secure-sourcing a small portion of parameters can provide sufficient security against model distillation, meanwhile, technical capabilities tend to be more challenging to distill than other domains.
B.9 Effectiveness of distillation difficulty
DD is effective in larger models, with a clear negative correlation between DD and average distillation ratios. As shown in Figure 9 (a), the Pearson coefficient for Llama2-7B consistently remains below -0.80, reaching as low as -0.98. Similar trends are observed in other models with varying architectures and sizes, confirming DD as a reliable predictor of distilled model performance and demonstrating the effectiveness of SOLID. Furthermore, the complete Pearson and Spearman results, presented in Table 11, reveal a negative correlation between RS and the average distillation ratio. For instance, in Llama2-7B, both the Pearson and Spearman coefficients fall below -0.77. These trends hold across models of different architectures and sizes. Additionally, Figure 10 displays scatter plots illustrating the relationship between $\Delta$ADR and Distillation Difficulty($\uparrow$)s for four models, with corresponding Pearson and Spearman correlation coefficients. The Distillation Difficulty($\uparrow$)s, obtained from Section 5.3, show a clear inverse relationship: an increase in $\Delta$ADR leads to a decrease in model scores across all models analyzed. This trend is consistently supported by strong negative correlations, with the most significant drop in model scores seen in Phi2-2.7B, where $\Delta$ADR shows the most substantial negative correlation.
DD is ineffective in smaller OPT model, with notable inconsistencies with ADRs. As shown in Figure 9 (b), the negative correlation between DD and ADR in OPT-350M is weak (coefficients above -0.33), showing its unsuitability for predicting distillation performance. Additionally, optimal security is achieved by protecting the middle layers rather than the initial or output layers, making SOLID unable to identify the smallest secured set.
Figure 9: (a) presents the Pearson coefficient between distillation ratio (DR) and distillation difficulty (DD) across four models and six domains. (b) depicts the link between ADR and DD for Llama2-7B and OPT-350M.
Table 11: Correlation coefficients (Spearman | Pearson) between distillation ratio and distillation difficult.
Model Rsn. Read. Knl. Code & Math Gen. Avg.
Llama2-7B -0.83 |||| -0.97 -0.77 |||| -0.96 -0.83 |||| -0.95 -0.85 |||| -0.90 -0.82 |||| -0.93 -0.80 |||| -0.98
Mistral-7B -0.83 |||| -0.89 -0.82 |||| -0.91 -0.82 |||| -0.94 -0.78 |||| -0.95 -0.76 |||| -0.87 -0.87 |||| -0.92
Phi-2 -0.93 |||| -0.96 -0.84 |||| -0.96 -0.84 |||| -0.87 -0.84 |||| -0.80 -0.84 |||| -0.84 -0.87 |||| -0.95
Phi-1.5 -0.86 |||| -0.97 -0.78 |||| -0.94 -0.83 |||| -0.94 -0.90 |||| -0.80 -0.84 |||| -0.89 -0.80 |||| -0.94
Figure 10: Correlation Analysis of $\Delta$ADR and Distillation Difficulty Across Different Models.
Appendix C Detailed Results
C.1 Comparison in two semi-open Llama2-70B
In this experiment, we examine two semi-open Llama2-70B models, where either the first two decoder layers are secure-source (referred to as Bottom2-Secured) or the last two decoder layers are secure-source (referred to as Top2-Secured). The objective is to compare their performance in terms of customization and their security under the distillation attack. The results are summarized in Table 12 and Table 13.
Table 12: Customization Performance of Llama2-70B under Different Secured Layers
Math Code Medical Law Finance Alignment
Fully Secured 53.15 24.90 53.68 79.63 37.54 7.19
Bottom2-Secured 62.40 43.99 62.73 93.85 87.51 7.46
Top2-Secured 62.53 42.36 62.72 93.91 87.90 7.46
Table 13: Distillation Performance of Llama2-70B under Different Secured Layers
Benchmarks Fully Secured Semi-Open-1 Semi-Open-2
Rsn. PIQA 50.82 50.49 79.05
winogrande 51.07 51.22 72.93
arc_easy 25.17 25.63 76.30
arc_challenge 23.55 20.48 50.17
Hellaswag 26.65 25.77 79.49
Read. lambada 0.00 0.01 57.25
BoolQ 43.30 37.92 84.95
SQuADv2_EM 0.00 0.00 1.54
SQuADv2_f1 0.23 1.01 35.59
OBQA 25.60 24.40 44.00
Knl. NQ 0.00 0.00 15.18
TriviaQA 0.00 0.00 52.67
Code mbpp 0.00 0.00 16.00
HumanEval 0.00 0.00 13.41
Math GSM8K 0.03 0.01 27.75
Gen. MMLU 23.01 23.22 63.61
BBH 0.00 0.00 49.45
Average Distillation Ratio($\downarrow$) 22.55 21.73 74.94
C.2 Evaluation Results under FT-all attack
In this section, we provide a comprehensive analysis of the evaluation results, comparing SOLID with two baseline methods: SAP-DP and a fully-secured approach. This comparison is conducted across 16 benchmarks under the FT-all attack scenario. The detailed results for Llama2-70B are presented in Table 14, while the results for Llama2-7B and Mistral-7B are shown in Table 15. Additionally, the outcomes for Phi-2 and Phi-1.5 are provided in Tables 16.
Table 14: Evaluation results of Llama2-70B under FT-all attack
Pre-train SOLID SAP-DP Fully-secured
Rsn. PIQA 80.69 50.49 48.26 50.82
Winogrande 74.74 51.22 50.59 51.07
ARC-easy 80.35 25.63 26.35 25.17
ARC-challenge 53.24 20.48 20.31 23.55
Hellaswag 82.15 25.77 25.76 26.65
Read. LAMBADA 75.07 0.01 0.00 0.00
BoolQ 86.70 37.92 37.83 43.30
SQuADv2_EM 51.23 0.00 0.00 0.00
SQuADv2_f1 67.43 1.01 1.13 0.23
OBQA 44.80 24.40 24.40 25.60
Knl. NaturalQuestions 32.38 0.00 0.00 0.00
TriviaQA 73.47 0.00 0.02 0.00
Code MBPP 24.80 0.00 0.00 0.00
HumanEval 25.00 0.00 0.00 0.00
Math GSM8K 53.15 0.01 0.00 0.03
Gen. MMLU 63.09 23.22 24.19 23.01
BBH 61.40 0.00 0.00 0.00
Average Distillation Ratio($\downarrow$) - 21.73 21.64 22.55
Table 15: Evaluation results of 7B models under FT-all attack
Llama2-7B Mistral-7B
SOLID SAP-DP Fully-secured SOLID SAP-DP Fully-secured
Rsn. PIQA 49.56 49.56 49.47 51.63 50.22 49.35
Winogrande 50.99 49.66 50.83 49.78 51.07 50.59
ARC-easy 27.04 26.43 25.98 26.12 28.03 25.83
ARC-challenge 21.07 20.56 22.47 19.94 21.42 22.35
Hellaswag 25.56 25.69 26.39 26.10 25.97 25.39
Read. LAMBADA 0.01 0.00 0.01 0.12 0.00 0.01
BoolQ 44.30 41.70 48.34 39.05 37.83 45.80
SQuADv2_EM 0.00 0.00 0.00 0.00 0.00 0.00
SQuADv2_f1 0.49 0.63 0.59 1.21 0.26 0.66
OBQA 25.13 23.00 25.93 25.60 25.20 25.00
Knl. NaturalQuestions 0.01 0.01 0.04 0.00 0.00 0.02
TriviaQA 0.00 0.00 0.02 0.00 0.00 0.01
Code MBPP 0.00 0.00 0.00 0.00 0.00 0.00
HumanEval 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 24.26 22.92 24.45 25.24 23.05 23.26
BBH 0.00 0.00 0.00 0.00 0.00 0.00
Average Distillation Ratio($\downarrow$) 25.03 24.16 25.62 22.41 22.28 22.68
Table 16: Evaluation results of small models under FT-all attack
Phi-2 Phi-1.5
SOLID SAP-DP Fully-secured SOLID SAP-DP Fully-secured
Rsn. PIQA 54.17 52.01 52.07 53.43 52.61 50.44
Winogrande 51.56 48.93 48.91 51.09 49.25 49.12
ARC_easy 34.57 28.20 27.03 30.81 28.79 27.50
ARC_challenge 19.45 19.37 18.66 20.56 19.80 21.22
Hellaswag 27.61 25.32 25.26 26.27 25.66 25.05
Read. LAMBADA 0.75 0.02 0.00 0.59 0.00 0.00
BoolQ 45.29 40.21 44.60 46.98 41.80 46.28
SQuADv2_EM 0.02 0.00 0.00 0.00 0.00 0.00
SQuADv2_f1 2.61 0.28 0.64 0.78 0.65 1.60
OBQA 24.80 26.60 25.80 26.60 28.60 26.87
Knl. NaturalQuestions 0.00 0.00 0.02 0.04 0.00 0.00
TriviaQA 0.01 0.00 0.01 0.01 0.00 0.00
Code MBPP 0.00 0.00 0.00 0.00 0.00 0.00
HumanEval 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 24.16 22.87 22.95 24.07 22.95 22.95
BBH 0.01 0.00 0.00 0.00 0.00 0.00
C.3 Customization Performance of Models
In this section, we present detailed evaluation results of the model customization performance across six downstream tasks used in our experiments. The detailed results for Llama2-70B are presented in Table 17, while the results for Llama2-7B and Mistral-7B are shown in Table 18 and Table 19. Additionally, the outcomes for Phi-2 and Phi-1.5 are provided in Tables 20 and Table 21.
Table 17: Detailed results of Llama2-70B secure by SOLID on six downstream tasks.
Math Code Medical Law Finance Alignment
Fully-Secure 53.15 24.90 53.68 79.63 55.63 7.19
SAP-DP 61.10 36.87 54.55 83.40 65.78 7.41
SOLID 62.40 43.99 62.73 93.85 87.51 7.46
Fully-Open 64.06 44.58 63.40 94.17 88.22 7.42
Table 18: Detailed results of Llama2-7B secure by SOLID on six downstream tasks.
Math Code Medical Law Finance Alignment
Fully-Secure 20.24 13.75 36.91 51.80 38.71 6.51
SAP-DP 20.24 13.75 36.91 51.80 38.71 6.52
SOLID 28.96 21.37 46.52 90.84 81.95 6.63
Fully-Open 29.34 21.265 47.60 90.49 84.09 6.63
Table 19: Detailed results of Mistral-7B secure by SOLID on six downstream tasks.
Math Code Medical Law Finance Alignment
Fully-Secure 38.21 33.83 61.50 50.47 37.39 3.20
SAP-DP 41.47 34.44 63.08 50.37 38.10 2.47
SOLID 46.10 43.16 66.78 84.94 86.19 3.87
Fully-Open 45.26 46.08 66.47 88.13 84.91 3.78
Table 20: Detailed results of Phi-2 secure by SOLID on six downstream tasks.
Math Code Medical Law Finance Alignment
Fully-Secure 57.77 47.59 43.13 56.46 54.07 5.22
SAP-DP 58.52 46.65 43.40 56.81 54.37 5.11
SOLID 59.59 47.79 45.85 57.11 56.26 5.26
Fully-Open 59.60 48.40 45.93 57.19 56.68 5.27
Table 21: Detailed results of Phi-1.5 secure by SOLID on six downstream tasks.
Math Code Medical Law Finance Alignment
Fully-Secure 30.33 35.09 30.78 52.18 34.60 3.24
SAP-DP 30.25 35.45 32.66 51.99 34.27 3.68
SOLID 33.66 37.10 33.14 52.26 39.60 3.87
Fully-Open 34.49 37.45 33.23 52.34 39.90 3.68
C.4 Comparison in deployment baselines on llama2-70B
We compare the distillation security of SOLID with SAP-DP and Fully-secured as baselines under FT-secure and SEM attack strategies. The evaluation results on sixteen benchmarks are shown in Table 22.
Table 22: Evaluation results of Llama2-70B under FT-secure and SEM attack
FT-secure SEM
SOLID SAP-DP SOLID SAP-DP
Rsn. PIQA 49.78 49.40 48.62 49.00
Winogrande 51.30 49.01 50.99 51.13
ARC-easy 26.43 25.59 25.33 24.55
ARC-challenge 21.41 21.42 22.01 20.93
Hellaswag 26.07 26.10 25.90 25.22
Read. LAMBADA 0.00 0.00 0.00 0.00
BoolQ 45.09 37.83 44.95 39.80
SQuADv2_EM 0.00 0.00 0.00 0.00
SQuADv2_f1 0.98 1.01 0.59 1.00
OBQA 24.40 23.80 25.03 22.96
Knl. NaturalQuestions 0.00 0.00 0.00 0.00
TriviaQA 0.00 0.00 0.00 0.00
Code MBPP 0.00 0.00 0.00 0.00
HumanEval 0.00 0.00 0.00 0.00
Math GSM8K 0.00 0.00 0.00 0.00
Gen. MMLU 23.18 23.66 22.98 22.83
BBH 0.00 0.00 0.00 0.00
Average Distillation Ratio($\downarrow$) 22.60 21.80 22.40 22.30
C.5 Comparison in Distillation Attack Strategies
In this section, we present detailed evaluation results of the model distillation performance of SOLID under FT-secure and SEM attack strategies across six functionalities used in our experiments. The detailed results under the FT-secure distillation strategy are presented in Table 23. The results under SEM attack strategies are shown in Table 24.
Table 23: Distillation Performance of SOLID under FT-Secure attacks.
Llama2-7B Mistral-7B Phi-2 Phi-1.5
Rsn. PIQA 49.95 49.55 54.57 52.45
Winogrande 49.88 49.68 52.33 52.41
ARC-easy 27.65 25.88 33.33 31.06
ARC-challenge 20.81 22.69 19.03 18.77
Hellaswag 26.04 25.01 27.62 26.88
Read. LAMBADA 0.00 0.00 0.77 0.71
BoolQ 38.13 46.01 44.34 57.49
SQuADv2_EM 0.00 0.00 0.00 0.00
SQuADv2_f1 0.22 0.36 3.07 2.27
OBQA 25.70 25.12 24.40 25.20
Knl. NaturalQuestions 0.00 0.00 0.00 0.00
TriviaQA 0.00 0.00 0.01 0.00
Code MBPP 0.00 0.00 0.00 0.00
HumanEval 0.00 0.00 0.00 0.00
Math GSM8K 0.00 0.00 0.00 0.00
Gen. MMLU 24.23 23.56 23.03 24.10
BBH 0.00 0.00 0.00 0.00
Average Distillation Ratio($\downarrow$) 24.80 22.50 23.56 26.97
Table 24: Distillation Performance of SOLID under SEM attacks.
Llama2-7B Mistral-7B Phi-2 Phi-1.5
Rsn. PIQA 51.52 48.53 49.46 50.82
Winogrande 50.28 51.02 48.70 50.59
ARC-easy 24.83 25.83 25.93 24.62
ARC-challenge 24.99 22.35 20.65 21.08
Hellaswag 25.58 25.39 25.84 25.39
Read. LAMBADA 0.00 0.01 0.00 0.01
BoolQ 53.30 45.80 38.41 61.07
SQuADv2_EM 0.00 0.00 0.00 0.00
SQuADv2_f1 0.77 0.66 0.00 1.35
OBQA 25.00 25.00 27.80 30.40
Knl. NaturalQuestions 0.00 0.02 0.00 0.00
TriviaQA 0.00 0.01 0.01 0.00
Code MBPP 0.00 0.00 0.00 0.00
HumanEval 0.00 0.00 0.00 0.00
Math GSM8K 0.00 0.00 0.00 0.00
Gen. MMLU 25.39 23.26 22.95 23.11
BBH 0.00 0.00 0.00 0.00
Average Distillation Ratio($\downarrow$) 25.00 22.00 22.10 24.70
C.6 Comparison in Distillation datasets scales
To investigate the impact of attack dataset scales on the efficiency of SOLID, we conduct model distillation attack on the Llama2-7B model using four different attack datasets of varying sizes: 100k, 200k, 300k, and 500k. The evaluation performance under different attack set scales are in Table 25
Table 25: Evaluation Results of SOLID on Llama2-7B under Various Attack Set Scales.
51K 100K 200K 300K 500K
Rsn. PIQA 49.56 49.89 49.18 49.18 49.59
Winogrande 50.99 47.99 49.49 50.20 50.20
ARC-easy 27.04 27.06 27.06 27.02 27.01
ARC-challenge 21.07 21.33 20.90 21.16 21.48
Hellaswag 25.56 26.49 26.46 26.50 26.19
Read. LAMBADA 0.01 0.01 0.00 0.00 0.01
BoolQ 44.30 44.41 44.10 44.07 44.96
SQuADv2_EM 0.00 0.00 0.00 0.02 0.00
SQuADv2_f1 1.05 0.32 0.51 0.52 0.71
OBQA 25.13 25.00 23.80 25.20 25.60
Knl. NaturalQuestions 0.01 0.08 0.08 0.06 0.06
TriviaQA 0.00 0.02 0.01 0.03 0.01
Code MBPP 0.00 0.00 0.00 0.00 0.00
HumanEval 0.00 0.00 0.00 0.00 0.00
Math GSM8K 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 24.26 25.34 25.43 26.14 26.41
BBH 0.00 0.00 0.00 0.00 0.00
Average Distillation Ratio($\downarrow$) 25.07 25.03 24.89 25.26 25.48
C.7 Transition Layer Results.
Security Performance. We close same-sized layer sets with different start points, and attack them using FT-all. Specifically, the sets consist of one layer for Llama2-7B (Table 26, Table 27), and two layers for Phi-2 (Table 30, Table 31). We further computed the $\Delta\text{ADR}$ for each secure-source set within Mistral-7B-v0.1 and Phi-1.5 in Appendix B.7. The results for the Mistral-7B-v0.1 model are presented in Table 28 and Table 29. Additionally, the performance outcomes for the Phi-1.5 model can be found in Table 32.
In all the above tables, “Pretrain” represents the model’s original performance without any layers secured. These columns indicate the index of layers in the model that have been secured. “*” indicates fully-secured. All evaluation scores are averages from three different seed tests, corresponding to the values 20, 42, and 1234, following the details of the Sixteen Functionality Benchmarks in Appendix B.5.
Table 26: Evaluation Results of Llama2-7B under Different Secure Layers (Part1)
Pretrain 0 1 2 3 4 5 6 7
Rsn. PIQA 76.66 49.56 51.43 49.53 50.45 49.84 50.27 50.96 51.09
Hellaswag 75.45 25.56 25.75 25.88 26.16 25.91 27.20 29.39 28.89
Winogrande 66.38 50.99 50.86 50.15 49.75 49.96 50.91 51.64 51.36
ARC_easy 74.41 27.04 27.23 26.10 26.30 25.51 26.44 28.24 27.96
ARC_challenge 44.11 21.07 20.31 20.19 21.30 22.04 21.56 20.62 22.92
Read. OpenBookQA 68.49 0.01 0.11 0.02 0.02 0.01 0.00 0.05 0.04
LAMBADA 80.67 44.30 41.22 38.36 41.43 38.08 38.14 38.40 41.55
BoolQ 59.48 0.00 0.04 0.00 0.00 0.00 0.00 0.01 0.03
SQuADv2_em 71.88 1.05 1.31 0.63 1.07 0.45 0.44 1.13 1.10
SQuADv2_f1 43.80 25.13 24.60 23.60 24.93 25.67 24.47 25.07 26.00
Knl. NaturalQuestions 22.47 0.01 0.00 0.01 0.03 0.02 0.01 0.13 0.08
TriviaQA 57.23 0.00 0.01 0.00 0.02 0.01 0.01 0.07 0.10
Code HumanEval 10.90 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 16.60 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 20.24 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 45.83 24.26 25.37 23.98 24.26 24.75 24.01 25.23 27.45
BBH 39.86 0.00 0.00 0.00 0.00 0.00 0.00 0.50 0.38
Avg. Performance Score($\downarrow$) 51.44 15.82 15.78 15.20 15.63 15.43 15.50 15.97 16.41
Average Distillation Ratio($\downarrow$) - 30.76 30.67 29.55 30.39 29.99 30.13 31.04 31.90
Distillation Difficulty($\uparrow$) - 11.11 11.27 10.87 10.31 10.83 10.33 10.90 11.11
Table 27: Evaluation Results of Llama2-7B under Different Secured Layers (Part2). * indicates the fully secured model.
16 18 20 22 24 26 28 30 *
Rsn. PIQA 51.47 52.99 58.22 65.83 69.60 73.45 75.46 75.99 49.47
Hellaswag 31.38 36.55 45.61 56.60 62.70 67.88 71.37 72.94 26.39
Winogrande 53.09 55.98 58.96 64.12 64.80 65.25 65.46 66.53 50.83
ARC_easy 30.58 35.35 43.85 55.92 62.56 68.36 70.85 72.60 25.98
ARC_challenge 24.26 26.85 30.97 35.38 38.17 41.41 43.00 44.17 22.47
Read. OpenBookQA 0.28 1.58 6.79 30.88 44.58 56.23 62.33 63.11 0.01
LAMBADA 57.55 70.53 71.36 78.85 79.69 80.29 79.39 80.40 48.34
BoolQ 0.08 0.90 2.34 7.07 6.04 6.87 3.54 9.46 0.00
SQuADv2_em 2.21 13.48 21.47 35.72 36.96 39.32 37.08 42.08 0.59
SQuADv2_f1 27.33 28.20 30.47 32.13 34.93 39.27 39.93 41.53 25.93
Knl. NaturalQuestions 0.13 0.41 1.60 2.94 4.29 2.69 7.28 11.87 0.04
TriviaQA 0.25 1.79 4.93 11.02 15.73 17.95 33.19 42.26 0.02
Code HumanEval 0.00 0.00 0.00 0.00 0.00 3.25 8.34 10.98 0.00
MBPP 0.00 0.00 0.00 0.07 0.47 2.27 8.80 13.27 0.00
Math GSM8K 0.00 0.00 0.00 0.13 0.81 8.42 6.90 15.77 0.00
Gen. MMLU 43.17 48.20 49.38 49.58 49.72 50.03 50.75 50.61 24.45
BBH 0.76 11.44 19.79 28.87 31.16 35.98 38.24 40.54 0.00
Avg. Performance Score($\downarrow$) 18.97 22.60 26.22 32.65 35.42 38.76 41.29 44.36 16.15
Average Distillation Ratio($\downarrow$) 36.89 43.94 50.98 63.48 68.87 75.35 80.27 86.24 31.39
Distillation Difficulty($\uparrow$) 10.42 9.49 8.86 7.12 6.14 4.72 3.40 3.06 11.19
Table 28: Evaluation Results of Mistral-7B under Different Secured Layers (Part1)
Pretrain 0 1 2 3 4 5 6 7
Rsn. PIQA 81.99 51.63 53.20 53.63 53.47 51.56 52.61 50.71 55.15
Hellaswag 81.04 26.10 26.36 26.36 26.66 27.10 25.51 26.18 28.16
Winogrande 74.03 49.78 49.78 51.01 50.38 49.91 50.14 49.70 51.17
ARC_easy 80.77 33.03 31.96 30.71 29.66 30.25 30.35 26.44 32.38
ARC_challenge 50.26 19.94 21.27 20.45 19.60 20.05 21.36 21.25 20.73
Read. OpenBookQA 44.40 25.60 25.20 25.20 25.47 25.87 26.33 25.07 27.20
LAMBADA 73.29 0.12 0.44 1.91 2.08 0.80 0.30 0.17 1.95
BoolQ 83.67 39.05 53.12 45.95 38.61 47.35 38.06 46.44 47.66
SQuADv2_em 64.04 0.00 0.00 0.01 0.01 0.01 0.00 0.00 0.01
SQuADv2_f1 71.37 1.21 0.84 1.05 1.03 1.27 0.43 0.07 0.86
Knl. NaturalQuestions 28.98 0.00 0.01 0.00 0.04 0.01 0.00 0.02 0.07
TriviaQA 70.79 0.00 0.00 0.02 0.01 0.01 0.01 0.00 0.16
Code HumanEval 29.88 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 38.40 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 38.21 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 62.50 25.24 24.68 25.11 23.43 23.65 24.26 24.26 24.99
BBH 56.40 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.01
Avg. Performance Score($\downarrow$) 60.59 15.98 16.87 16.55 15.91 16.34 15.84 15.90 17.09
Average Distillation Ratio($\downarrow$) - 26.38 27.85 27.32 26.25 26.97 26.15 26.24 28.20
Average Distillation Ratio($\downarrow$) - 11.50 11.31 11.48 10.71 10.77 11.44 11.02 10.71
Table 29: Evaluation Results of Mistral-7B under Different Secured Layers (Part2)
16 18 20 22 24 26 28 30 *
Rsn. PIQA 54.50 52.32 52.72 57.13 62.82 64.67 67.23 75.61 49.35
Hellaswag 29.31 29.02 29.99 33.46 46.21 52.12 52.46 67.73 25.39
Winogrande 51.20 54.17 51.07 55.75 58.59 62.41 63.09 66.33 50.59
ARC_easy 32.84 29.35 30.80 38.04 47.24 51.99 54.74 69.95 25.83
ARC_challenge 21.19 23.04 23.78 26.34 30.86 33.22 35.04 40.53 22.35
Read. OpenBookQA 26.00 27.87 26.87 29.67 28.73 32.67 33.40 36.40 25.00
LAMBADA 2.61 0.18 1.28 4.17 21.89 29.93 24.49 48.32 0.01
BoolQ 53.98 53.60 58.79 55.76 64.10 74.72 68.48 81.30 45.80
SQuADv2_em 0.01 0.00 0.47 0.13 2.39 3.59 1.87 1.82 0.00
SQuADv2_f1 0.96 0.18 1.27 2.60 14.88 22.61 21.12 34.16 0.66
Knl. NaturalQuestions 0.01 0.10 0.19 0.58 1.84 3.15 3.53 8.87 0.02
TriviaQA 0.03 0.01 0.61 0.62 5.14 7.51 10.32 25.44 0.01
Code HumanEval 0.00 0.00 0.00 0.61 2.24 4.88 2.44 9.75 0.00
MBPP 0.00 0.00 0.00 2.00 4.33 8.33 0.93 13.07 0.00
Math GSM8K 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.25 0.00
Gen. MMLU 24.30 25.84 29.54 24.55 34.77 40.77 40.84 50.44 23.26
BBH 0.00 0.00 0.02 0.30 7.55 18.76 21.05 30.07 0.00
Avg. Performance Score($\downarrow$) 17.47 17.39 18.08 19.51 25.51 30.08 29.47 38.83 15.78
Average Distillation Ratio($\downarrow$) 28.83 28.71 29.84 32.20 42.09 49.64 48.64 64.08 26.05
Average Distillation Ratio($\downarrow$) 11.34 11.11 10.45 10.59 10.23 10.34 9.59 8.53 11.20
Table 30: Evaluation Results of Phi-2 under Different Secured Layers (Part 1)
Pretrain 0 2 4 6 8 10 12 14
Rsn. PIQA 79.27 54.17 72.85 73.76 75.03 76.75 78.00 78.91 77.84
Hellaswag 73.73 27.61 56.49 57.73 60.47 62.84 66.39 66.91 66.95
Winogrande 75.45 51.56 59.17 59.98 59.88 64.32 68.11 68.95 70.38
ARC_easy 79.92 34.57 72.94 73.40 73.97 76.51 78.33 78.66 78.63
ARC_challenge 52.90 19.45 41.75 39.82 44.11 45.65 47.92 49.74 48.78
Read. OpenBookQA 51.20 25.80 35.73 37.47 40.13 42.00 44.00 45.67 44.80
LAMBADA 56.28 3.25 28.55 30.42 34.64 40.05 45.41 45.52 46.66
BoolQ 83.36 47.29 65.20 62.64 66.39 71.39 73.42 72.95 75.83
SQuADv2_em 61.30 0.02 10.49 17.63 21.94 33.94 19.54 19.15 29.14
SQuADv2_f1 71.38 2.61 37.22 40.35 45.53 59.16 48.21 50.09 54.87
Knl. NaturalQuestions 9.58 0.00 3.60 4.97 6.13 7.55 7.95 8.10 9.25
TriviaQA 39.29 0.01 13.57 16.29 24.74 28.60 31.58 33.71 32.79
Code HumanEval 48.78 0.00 1.42 6.50 10.98 16.66 22.76 19.51 23.17
MBPP 46.80 0.00 5.07 6.87 9.47 19.60 25.67 23.47 25.73
Math GSM8K 57.77 0.00 7.25 8.64 4.42 9.63 14.18 11.35 17.31
Gen. MMLU 56.73 26.16 34.29 37.01 39.90 43.11 45.63 48.17 49.82
BBH 59.53 0.01 15.27 18.37 16.38 14.58 4.93 4.35 11.37
Avg. Performance Score($\downarrow$) 59.02 17.21 32.99 34.81 37.30 41.90 42.47 42.66 44.90
Average Distillation Ratio($\downarrow$) - 29.15 55.90 58.99 63.21 71.00 71.97 72.28 76.09
Distillation Difficulty($\uparrow$) - 10.07 7.07 4.95 4.09 3.63 3.31 3.31 3.11
Table 31: Evaluation Results of Phi-2 under Different Secured Layers (Part2). “*” indicates the fully secured model.
16 18 20 22 24 26 28 30 *
Rsn. PIQA 77.44 77.80 77.69 76.77 76.89 77.55 78.16 78.58 52.07
Hellaswag 67.20 66.90 67.13 68.00 68.86 70.01 71.44 71.18 25.26
Winogrande 70.82 71.40 73.11 74.46 75.79 75.72 75.93 74.77 48.91
ARC_easy 78.30 77.27 77.33 76.82 78.09 77.76 79.53 79.56 27.03
ARC_challenge 49.71 48.29 48.52 48.04 49.80 50.68 53.16 52.67 18.66
Read. OpenBookQA 46.53 46.47 45.87 45.27 46.33 45.53 46.53 48.27 20.80
LAMBADA 45.67 46.88 47.95 50.17 50.54 52.77 53.01 53.23 0.00
BoolQ 80.56 80.72 82.22 83.31 83.98 83.54 82.54 83.41 39.60
SQuADv2_em 7.88 1.30 1.69 1.31 0.15 0.23 3.54 10.03 0.56
SQuADv2_f1 40.84 34.51 34.25 35.94 35.64 36.68 39.57 44.87 0.90
Knl. NaturalQuestions 8.90 6.09 6.40 6.79 6.86 6.85 7.20 8.37 0.02
TriviaQA 31.48 27.03 25.08 24.54 22.89 22.99 24.24 26.93 0.01
Code HumanEval 22.56 21.34 25.41 32.52 38.01 46.14 46.54 43.90 0.00
MBPP 26.73 25.33 24.80 31.73 36.67 41.80 43.13 43.20 0.00
Math GSM8K 16.68 16.02 14.66 12.31 17.24 30.12 45.41 49.79 0.00
Gen. MMLU 52.69 53.45 55.68 56.61 56.93 56.59 56.86 56.47 22.95
BBH 3.42 17.36 8.33 18.24 30.09 48.12 52.28 56.36 0.00
Avg. Performance Score($\downarrow$) 42.79 42.25 42.12 43.70 45.57 48.42 50.53 51.86 15.10
Average Distillation Ratio($\downarrow$) 72.51 71.58 71.38 74.04 77.22 82.04 85.63 87.87 25.59
Distillation Difficulty($\uparrow$) 3.07 3.29 3.03 3.01 2.70 2.32 1.98 2.13 11.32
Table 32: Evaluation Results of Phi-1.5 under Different Secured Layers
Pretrain 0-1 2-3 4-5 6-7 8-9 10-11 12-13 14-15 16-17 18-19 20-21 22-23 *
Rsn. PIQA 75.68 53.43 69.52 71.53 73.50 74.76 75.08 74.94 74.64 73.90 74.63 74.54 74.81 50.44
Hellaswag 62.56 26.27 46.66 50.71 52.98 54.51 55.11 56.01 56.78 57.90 58.76 59.35 58.58 25.05
Winogrande 72.69 51.09 54.91 59.22 61.75 64.85 67.95 68.88 68.98 71.25 71.19 72.87 70.66 49.12
ARC_easy 76.14 30.81 61.70 65.70 70.10 71.38 70.01 71.72 71.93 72.34 73.39 74.16 73.74 27.50
ARC_challenge 44.62 20.56 32.85 34.10 38.08 40.05 40.30 39.48 40.87 41.52 42.84 42.58 45.42 21.22
Read. OpenBookQA 48.00 26.60 33.93 35.73 40.40 41.13 40.67 41.73 41.67 40.27 41.33 43.27 45.47 26.87
LAMBADA 44.10 0.59 17.96 26.45 29.37 33.83 33.85 36.46 37.06 37.96 39.98 41.10 40.49 0.00
BoolQ 75.05 46.98 59.12 52.42 57.41 65.68 68.52 63.47 65.12 66.52 73.91 75.17 77.0 46.28
SQuADv2_em 48.01 0.00 5.82 10.94 18.34 13.96 14.70 23.22 16.98 26.05 22.04 20.16 26.86 0.00
SQuADv2_f1 60.84 0.78 24.49 26.04 34.86 32.17 32.36 43.14 38.23 48.03 45.75 45.56 49.62 1.60
Knl. NaturalQuestions 5.46 0.04 1.68 2.73 3.41 3.06 3.21 4.25 4.03 4.06 4.54 4.17 4.45 0.01
TriviaQA 16.94 0.01 5.70 7.77 10.85 11.03 9.11 12.11 11.84 11.86 12.02 12.11 13.19 0.01
Code HumanEval 35.98 0.00 3.05 10.57 12.20 16.26 13.82 17.48 18.70 23.17 29.68 31.91 31.71 0.00
MBPP 35.40 0.00 2.80 7.80 10.93 17.40 16.53 16.13 16.67 22.27 27.33 28.27 28.53 0.00
Math GSM8K 30.33 0.00 0.05 0.73 0.15 0.23 0.75 0.50 2.17 4.98 9.73 17.77 23.45 0.00
Gen. MMLU 42.44 24.07 26.56 28.77 32.51 32.87 36.09 39.42 39.72 43.23 42.51 42.82 43.66 23.95
BBH 28.80 0.00 2.07 3.97 8.38 7.37 2.81 7.79 4.12 10.63 6.94 10.34 11.45 0.00
Avg. Performance Score($\downarrow$) 47.24 16.54 26.40 29.13 32.66 34.15 34.17 36.28 35.85 38.59 39.80 40.95 42.30 15.94
Average Distillation Ratio($\downarrow$) - 35.02 55.90 61.66 69.14 72.29 72.34 76.80 75.90 81.68 84.25 86.69 89.56 33.75
Average Distillation Ratio($\downarrow$) - 10.08 7.18 4.70 3.50 2.93 2.83 2.53 2.36 2.27 2.16 2.06 2.46 9.33
Customizability Performance. We close varying numbers of layers from the start and fine-tune the open set, and then we observe the customizability transition in models. Table 33 shows the detailed evaluation results of Llama2-7B and Phi-2 on GSM8k benchmark.
Table 33: Customization Performance under Different Secure Sets
Llama2-7B Phi-2
Secure Layers GSM8K($\uparrow$) Secure Layers GSM8K($\uparrow$)
Fully-open 29.34 Fully-open 59.60
0 28.96 0-1 59.59
0-4 21.76 0-5 58.60
0-8 21.46 0-9 58.45
0-12 20.85 0-13 55.19
0-16 20.11 0-17 56.25
0-20 21.46 0-21 54.59
0-24 21.44 0-25 55.34
0-28 18.73 0-29 54.59
Fully-Secure 20.32 Fully-Secure 57.77
C.8 Evaluation Results under Different Secure size
In this section, we present a comprehensive evaluation of the model’s performance across sixteen benchmarks utilized in our experiments. The evaluation results for LLaMA2-7B, categorized by varying quantities and proportions of secure-source parameters, are displayed in Table 34 and Table 35, respectively. For the Mistral-7B model, the results are summarized in Table 36 and Table 37. Furthermore, the evaluation outcomes for the Phi-2 model can be found in Tables 38 and Table 39. The performance results for Phi-1.5 are also included in Tables 40 and Table 41 for comparison. For further details regarding the secure-source settings employed in our experiments, please refer to Appendix C.8.
Table 34: Evaluation Results of Llama2-7B under Different Secure-source Proportion
0.25% 0.5% 1% 3% 7% 15% 30% 50% 100%
Rsn. PIQA 77.78 77.69 67.73 49.42 49.55 50.05 49.98 49.31 49.47
Hellaswag 71.40 71.54 52.39 25.74 26.03 26.25 25.67 25.48 26.39
Winogrande 64.64 65.64 54.12 50.38 50.43 49.65 49.59 49.62 50.83
ARC_easy 74.69 75.04 53.82 26.03 26.76 26.46 26.64 26.66 25.98
ARC_challenge 43.66 43.29 26.99 20.16 21.39 19.74 21.44 21.73 22.47
Read. OpenBookQA 63.15 63.62 33.20 0.01 0.00 0.02 0.01 0.01 0.01
LAMBADA 80.66 80.78 62.10 38.22 39.33 43.45 39.39 41.83 48.34
BoolQ 11.39 12.14 5.47 0.00 0.00 0.00 0.00 0.00 0.00
SQuADv2_em 40.24 40.74 32.65 0.78 0.20 0.24 2.09 2.13 0.59
SQuADv2_f1 40.73 40.67 30.47 22.93 23.40 25.53 24.07 23.07 25.93
Knl. NaturalQuestions 7.83 7.89 5.61 0.00 0.01 0.02 0.01 0.00 0.04
TriviaQA 44.29 45.95 18.78 0.00 0.01 0.00 0.00 0.00 0.02
Code HumanEval 11.39 12.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 15.20 15.33 1.00 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 13.22 13.29 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 45.04 45.03 30.90 24.06 24.04 25.01 23.19 23.11 24.45
BBH 37.45 37.51 17.36 0.00 0.00 0.00 0.00 0.00 0.00
Avg. Performance Score($\downarrow$) 43.69 44.01 28.98 15.16 15.36 15.67 15.42 15.47 16.15
Average Distillation Ratio($\downarrow$) 84.94 85.56 56.33 29.48 29.86 30.47 29.97 30.07 31.39
Distillation Difficulty($\uparrow$) 1.96 1.93 8.66 10.87 11.75 11.48 11.65 11.57 11.19
Table 35: Evaluation Results of Llama2-7B under Different Secure-source Quantity
20M 50M 100M 160M 200M 300M 600M
Rsn. PIQA 77.78 73.49 67.55 67.12 49.42 50.36 49.97
Hellaswag 71.40 63.47 51.67 51.27 25.74 25.70 25.78
Winogrande 64.64 57.54 53.07 52.04 50.38 49.28 50.49
ARC_easy 74.69 66.50 51.97 52.11 26.03 26.43 26.29
ARC_challenge 43.66 36.04 26.51 25.99 20.16 20.79 21.70
Read. OpenBookQA 63.15 45.34 30.22 28.75 0.01 0.05 0.01
LAMBADA 80.66 69.47 62.28 62.59 38.22 39.03 40.80
BoolQ 11.39 2.21 4.18 7.24 0.00 0.00 0.01
SQuADv2_em 40.24 33.98 28.98 31.05 0.78 0.74 0.37
SQuADv2_f1 40.73 33.93 29.13 30.00 22.93 23.80 23.53
Knl. NaturalQuestions 7.83 2.98 5.33 5.73 0.00 0.00 0.02
TriviaQA 44.29 15.28 13.71 17.25 0.00 0.00 0.01
Code HumanEval 11.39 0.41 0.00 0.00 0.00 0.00 0.00
MBPP 15.20 6.87 1.00 0.80 0.00 0.00 0.00
Math GSM8K 9.00 0.10 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 45.04 36.15 28.95 29.04 24.06 23.70 23.45
BBH 37.45 28.53 14.99 16.99 0.00 0.00 0.00
Avg. Performance Score($\downarrow$) 43.44 33.66 27.62 28.12 15.16 15.29 15.44
Average Distillation Ratio($\downarrow$) 84.46 65.44 53.69 54.66 29.48 29.72 30.01
Distillation Difficulty($\uparrow$) 1.96 5.48 8.95 9.25 10.87 10.93 10.81
Table 36: Evaluation Results of Mistral-7B under Different Secured Proportion
0.25% 1% 0.5% 3% 7% 15% 30% 50% 100%
Rsn. PIQA 77.79 74.36 52.16 53.34 52.07 52.19 50.04 50.60 49.35
Hellaswag 71.31 65.50 26.50 26.16 25.92 25.91 25.87 25.61 25.39
Winogrande 67.09 60.32 49.22 51.65 50.01 51.36 51.36 49.65 50.59
ARC_easy 74.52 69.51 29.95 30.82 29.73 30.44 28.20 27.45 25.83
ARC_challenge 42.32 38.40 20.76 20.71 21.10 20.25 22.61 22.47 22.35
Read. OpenBookQA 42.13 34.60 25.13 25.33 26.47 26.07 25.20 25.87 25.00
LAMBADA 55.99 44.36 0.73 1.66 0.96 0.31 0.03 0.02 0.01
BoolQ 78.35 74.06 43.18 42.01 42.09 40.02 38.53 39.91 45.80
SQuADv2_em 13.91 6.97 0.00 0.01 0.00 0.00 0.00 0.00 0.00
SQuADv2_f1 41.13 33.88 1.60 0.93 1.27 0.71 0.99 0.86 0.66
Knl. NaturalQuestions 8.46 5.82 0.03 0.00 0.02 0.03 0.00 0.00 0.02
TriviaQA 34.04 17.03 0.01 0.01 0.02 0.01 0.00 0.00 0.01
Code HumanEval 11.99 6.51 0.00 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 16.93 12.80 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 6.32 0.45 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 44.17 37.98 23.98 24.34 25.10 23.91 23.68 24.12 23.26
BBH 35.44 27.27 0.02 0.00 0.00 0.00 0.00 0.00 0.00
Avg. Performance Score($\downarrow$) 42.46 35.87 16.08 16.29 16.16 15.95 15.68 15.68 15.78
Average Distillation Ratio($\downarrow$) 70.08 59.20 26.53 26.89 26.67 26.33 25.87 25.88 26.05
Distillation Difficulty($\uparrow$) 2.22 5.48 10.92 11.29 11.35 11.19 11.17 11.20 11.20
Table 37: Evaluation Results of Mistral-7B under Different Secured Quantity
20M 50M 100M 160M 200M 300M 600M
Rsn. PIQA 77.79 73.74 51.36 52.86 53.34 50.98 51.62
Hellaswag 71.31 65.51 26.49 27.98 26.16 26.27 26.04
Winogrande 67.09 64.51 50.06 49.51 51.65 50.17 50.85
ARC_easy 74.52 68.29 27.84 30.95 30.82 27.36 28.30
ARC_challenge 42.32 37.97 20.85 21.67 20.71 21.28 20.17
Read. OpenBookQA 42.13 37.27 25.60 25.87 25.33 26.60 27.00
LAMBADA 55.99 47.63 1.16 4.74 1.66 0.43 0.53
BoolQ 78.35 75.00 40.17 47.05 42.01 42.05 39.03
SQuADv2_em 13.91 8.65 0.01 0.04 0.01 0.01 0.00
SQuADv2_f1 41.13 35.50 1.01 0.49 0.93 0.28 0.39
Knl. NaturalQuestions 8.46 7.82 0.02 0.05 0.00 0.01 0.02
TriviaQA 34.04 22.89 0.02 0.19 0.01 0.01 0.01
Code HumanEval 11.99 7.93 0.00 0.00 0.00 0.00 0.00
MBPP 16.93 11.87 0.00 0.00 0.00 0.00 0.00
Math GSM8K 6.32 2.48 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 44.17 41.28 24.22 24.44 24.34 23.78 23.33
BBH 35.44 33.43 0.00 0.40 0.00 0.00 0.00
Avg. Performance Score($\downarrow$) 42.46 37.75 15.81 16.84 16.29 15.84 15.72
Average Distillation Ratio($\downarrow$) 70.08 62.31 26.10 27.79 26.89 26.14 25.95
Distillation Difficulty($\uparrow$) 2.22 3.44 11.14 10.85 11.10 11.23 11.22
Table 38: Evaluation Results of Phi-2 under Different Secured Proportion
0.25% 0.5% 1% 3% 7% 15% 30% 50% 100%
Rsn. PIQA 70.40 70.71 74.64 54.43 54.17 54.75 54.37 52.39 52.07
Hellaswag 53.13 52.99 62.84 27.88 27.61 27.77 28.01 26.30 25.26
Winogrande 66.17 66.43 69.93 51.49 51.56 51.46 51.44 49.12 48.91
ARC_easy 64.62 65.33 72.55 33.39 34.57 32.00 32.18 29.97 27.03
ARC_challenge 43.26 43.86 40.67 20.82 19.45 20.00 20.56 19.88 18.66
Read. OpenBookQA 41.80 42.67 38.87 26.87 25.80 26.33 26.53 26.07 20.80
LAMBADA 32.51 32.25 40.24 10.58 3.25 3.87 6.06 0.66 0.00
BoolQ 65.77 65.27 76.84 48.13 47.29 45.62 46.15 40.50 39.60
SQuADv2_em 0.36 9.09 3.31 0.02 0.02 0.01 0.01 0.00 0.56
SQuADv2_f1 24.81 30.83 30.47 0.45 2.61 0.57 2.52 1.67 0.90
Knl. NaturalQuestions 5.70 5.06 1.14 0.03 0.00 0.01 0.07 0.03 0.02
TriviaQA 20.27 21.50 8.78 2.02 0.01 0.02 0.01 0.01 0.01
Code HumanEval 22.16 26.83 17.68 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 25.07 26.40 9.73 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 29.26 31.36 2.00 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 41.76 42.17 43.86 30.31 26.16 25.79 24.85 24.03 22.95
BBH 18.98 21.55 9.59 3.06 0.01 0.79 0.24 0.00 0.00
Avg. Performance Score($\downarrow$) 36.83 38.49 35.48 18.20 17.21 17.00 17.24 15.92 15.10
Average Distillation Ratio($\downarrow$) 62.40 65.22 60.12 30.95 29.15 28.81 29.21 26.97 25.59
Distillation Difficulty($\uparrow$) 6.70 6.65 2.00 9.14 10.07 10.13 10.14 9.82 11.32
Table 39: Evaluation Results of Phi-2 under Different Secured Quantity
20M 50M 100M 160M 200M 300M 600M
Rsn. PIQA 73.70 70.00 53.90 54.17 53.01 54.75 54.28
Hellaswag 59.75 55.64 28.26 27.61 26.90 27.77 28.61
Winogrande 66.61 67.17 51.96 51.56 52.28 51.46 50.88
ARC_easy 70.96 67.02 35.17 34.57 31.84 32.00 31.62
ARC_challenge 48.30 42.52 21.84 19.45 20.39 20.00 20.56
Read. OpenBookQA 45.33 41.27 26.13 25.80 25.60 26.33 26.53
LAMBADA 35.64 25.34 1.93 3.25 2.17 3.87 5.78
BoolQ 75.37 66.25 51.66 47.29 40.81 45.62 47.69
SQuADv2_em 10.62 0.10 0.14 0.02 0.02 0.01 0.00
SQuADv2_f1 38.28 22.83 1.33 2.61 1.36 0.57 1.13
Knl. NaturalQuestions 5.44 4.51 0.06 0.00 0.02 0.01 0.05
TriviaQA 12.34 12.77 0.05 0.01 0.01 0.02 0.01
Code HumanEval 20.94 10.98 0.00 0.00 0.00 0.00 0.00
MBPP 12.60 13.40 0.00 0.00 0.00 0.00 0.00
Math GSM8K 7.52 7.78 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 43.07 39.45 26.26 26.16 25.85 25.79 25.38
BBH 12.35 18.02 0.00 0.01 0.00 0.79 0.12
Avg. Performance Score($\downarrow$) 37.57 33.24 17.57 17.21 16.49 17.00 17.22
Average Distillation Ratio($\downarrow$) 63.67 56.32 29.77 29.15 27.93 28.81 29.17
Distillation Difficulty($\uparrow$) 2.07 7.96 9.25 9.96 10.08 10.13 10.22
Table 40: Evaluation Results of Phi-1.5 under Different Secured Proportion
0.25% 0.5% 1% 3% 7% 15% 30% 50% 100%
Rsn. PIQA 68.21 68.37 69.68 65.85 53.43 52.94 52.36 51.25 50.44
Hellaswag 49.05 49.18 49.30 30.72 26.27 26.74 27.02 26.10 25.05
Winogrande 63.83 64.91 61.20 58.04 51.09 51.38 50.25 50.22 49.12
ARC_easy 62.94 62.89 62.25 35.15 30.81 29.27 29.64 27.99 27.50
ARC_challenge 36.98 37.49 32.91 25.97 20.56 20.36 20.08 20.88 21.22
Read. OpenBookQA 39.07 40.20 35.00 33.87 26.60 27.67 27.73 26.47 26.87
LAMBADA 24.71 24.99 25.36 0.11 0.59 0.78 1.15 0.06 0.00
BoolQ 59.43 59.35 63.49 41.01 46.98 51.59 46.46 44.02 46.28
SQuADv2_em 15.65 16.00 3.13 0.50 0.00 0.01 0.03 0.00 0.00
SQuADv2_f1 32.62 32.62 14.88 0.56 0.78 1.24 2.29 1.58 1.60
Knl. NaturalQuestions 2.72 2.64 0.32 0.03 0.04 0.03 0.05 0.03 0.01
TriviaQA 8.17 7.96 5.69 0.01 0.01 0.01 0.01 0.01 0.01
Code HumanEval 14.43 13.41 2.03 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 17.20 18.67 6.47 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 4.88 4.90 0.25 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 30.12 29.88 28.98 27.78 24.07 24.22 24.66 24.28 22.95
BBH 4.34 3.19 0.98 0.50 0.00 0.00 0.00 0.00 0.00
Avg. Performance Score($\downarrow$) 31.43 31.57 27.17 19.41 16.54 16.84 16.57 16.05 15.94
Average Distillation Ratio($\downarrow$) 66.54 66.83 57.52 41.11 35.02 35.64 35.08 33.98 33.75
Distillation Difficulty($\uparrow$) 6.18 6.15 2.76 9.28 10.08 11.19 10.54 10.23 11.26
Table 41: Evaluation Results of Phi-1.5 under Different Secured Quantity
20M 50M 100M 160M 200M 300M 600M
Rsn. PIQA 69.80 65.85 53.43 52.52 52.94 53.06 53.81
Hellaswag 49.51 25.72 30.27 26.31 26.74 27.05 26.51
Winogrande 62.56 58.04 51.09 50.83 51.38 50.57 49.99
ARC_easy 62.41 30.15 30.81 29.14 29.27 29.62 29.67
ARC_challenge 32.51 25.97 20.56 19.97 20.36 20.48 20.79
Read. OpenBookQA 35.53 33.87 26.60 26.93 27.67 28.20 26.87
LAMBADA 28.14 0.11 0.59 0.45 0.78 1.30 0.61
BoolQ 64.77 41.01 46.98 47.33 51.59 46.09 45.59
SQuADv2_em 4.67 0.50 0.00 0.00 0.01 0.01 0.00
SQuADv2_f1 22.47 0.56 0.78 1.02 1.24 2.31 2.01
Knl. NaturalQuestions 1.64 0.03 0.04 0.05 0.03 0.06 0.03
TriviaQA 5.93 0.01 0.01 0.01 0.01 0.02 0.01
Code HumanEval 7.73 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 7.87 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 0.28 0.00 0.00 0.00 0.00 0.00 0.00
Gen. MMLU 31.11 27.78 24.07 23.41 24.22 24.54 24.68
BBH 3.38 0.50 0.00 0.00 0.00 0.00 0.00
Avg. Performance Score($\downarrow$) 28.84 19.89 16.54 16.35 16.84 16.67 16.50
Average Distillation Ratio($\downarrow$) 61.06 41.11 35.02 34.61 35.64 35.28 34.94
Distillation Difficulty($\uparrow$) 2.81 9.28 10.26 11.65 11.19 10.87 10.49
C.9 Limitation on OPT-350M
[…]
Table 42: Evaluation Results of OPT-350M under Different Secured Layers. “*” indicates the fully secured model.
Pretrain 0-2 3-5 6-8 9-11 12-14 15-17 18-20 21-23 24-26 27-29 30-32 33-35 *
Rsn. PIQA 64.69 61.40 62.50 61.11 56.46 58.47 58.94 61.86 62.59 63.13 61.93 62.67 63.11 49.53
Hellaswag 36.68 34.03 34.27 33.69 31.79 32.24 32.78 33.27 33.68 33.25 33.94 33.63 33.07 25.77
Winogrande 52.09 51.62 52.96 51.57 50.83 52.83 51.06 51.52 51.85 52.06 52.04 51.99 50.94 49.85
ARC_easy 44.02 40.46 40.66 40.07 35.41 37.50 37.81 39.91 40.70 41.12 41.19 40.84 39.92 26.53
ARC_challenge 20.82 22.27 22.61 21.25 21.39 21.25 22.01 21.36 20.25 20.48 19.88 20.99 20.28 19.82
Read. OpenBookQA 28.00 27.60 27.47 27.40 27.40 27.27 26.20 26.47 27.67 26.80 28.67 27.67 27.13 27.47
LAMBADA 40.47 30.62 32.97 28.62 21.65 23.87 28.23 29.07 29.83 29.81 31.72 31.43 18.08 0.00
BoolQ 57.74 50.87 48.51 50.58 51.60 52.83 53.42 54.37 53.30 51.42 59.79 53.14 60.42 37.83
SQuADv2_em 11.34 6.87 7.88 4.74 4.19 0.27 0.87 2.22 3.79 3.05 4.11 4.69 2.35 0.00
SQuADv2_f1 19.35 16.27 17.00 12.00 11.72 9.04 6.92 10.11 10.90 10.08 8.88 11.47 7.30 0.01
Knl. NaturalQuestions 1.08 1.05 0.83 0.83 0.78 0.55 0.69 0.41 1.00 0.85 0.71 0.52 0.75 0.04
TriviaQA 4.48 2.24 2.66 2.01 2.16 1.41 1.06 2.39 2.38 2.29 1.57 1.90 1.76 0.02
Code HumanEval 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
MBPP 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Math GSM8K 1.59 0.15 0.25 0.00 0.08 0.00 0.05 0.00 0.03 0.18 0.00 0.00 0.00 0.00
Gen. MMLU 26.05 25.52 26.02 25.20 25.60 25.05 25.73 23.97 25.57 26.04 25.13 25.17 25.73 22.95
BBH 16.97 6.87 12.58 5.51 5.11 2.55 5.98 2.74 11.15 13.98 14.25 13.02 12.57 0.00
Avg. Performance Score($\downarrow$) 25.02 22.23 22.89 21.44 20.36 20.30 20.69 21.16 22.04 22.03 22.58 22.30 21.38 15.28
Average Distillation Ratio($\downarrow$) - 88.83 91.49 85.71 81.38 81.13 82.69 84.55 88.08 88.05 90.23 89.13 85.43 61.08
Distillation Difficulty($\uparrow$) - 5.92 9.32 9.04 8.60 8.83 8.73 7.17 5.82 5.18 4.65 4.92 4.15 10.89